### v6

# Boundary of the Moduli Space of Stable Cubic Fivefolds

Abstract
We study the boundary of the GIT quotient for cubic fivefolds from three complementary viewpoints: irreducible components, normal forms of closed orbits, and singularities. We give explicit normal forms representing closed (polystable) $\mathrm{SL}(7)$-orbits that appear on the boundary, and describe the geometry of the singular locus of each normal form (including the appearance of low-degree curves).
A key feature of the strictly semistable locus is that isolated singularities occur on the GIT boundary: we show that there are exactly six analytic types of isolated singularities appearing on strictly semistable cubic fivefolds, each given by a quasi-homogeneous normal-form family (some with moduli). For each of these six types we compute the local minimal exponent and prove that it is always equal to $\frac{7}{3}$. This value is precisely the critical threshold $\frac{n+1}{d}$ for degree $d=3$ hypersurfaces in $\mathbb{P}^{n}$ with $n=6$. In particular, these six isolated singularities realize the equality (critical) case in Park’s criterion for minimal exponents. Motivated by this, we call these six types extremal cubic fivefold singularities.
We further obtain strong computational evidence for the non-inclusion among the $21$ boundary families. Assuming Conjecture 3.5 (supported by these computations), it follows that the boundary decomposes into $21$ irreducible components. Finally, using wall-crossing in Kirwan’s stratification, we determine the adjacency relations among the boundary families; in particular, we show that the only nontrivial adjacencies occur in eight pairs.
Introduction
This paper studies the Geometric Invariant Theory (GIT) compactification of the moduli space of cubic fivefolds, i.e. cubic hypersurfaces $X\subset\mathbb{P}^{6}$, via the natural action of $\mathrm{SL}(7)$ on
$\mathbb{P}\bigl(\mathrm{Sym}^{3}\mathbb{C}^{7}\bigr)\quad\text{and the quotient}\quad\mathbb{P}\bigl(\mathrm{Sym}^{3}\mathbb{C}^{7}\bigr)//\mathrm{SL}(7).$
Our goal is to describe not only the stable locus but also the boundary contributed by strictly semistable cubics, i.e. the image of the strictly semistable locus in the GIT quotient. While the case of cubic fourfolds has been intensively studied from both classical and modern perspectives, a comparably explicit description of the GIT boundary for cubic fivefolds has not been fully developed: in particular, one would like concrete normal forms for closed orbits, an explicit description of the boundary families, and a systematic analysis of the singularities that govern the geometry of degenerations.
From the GIT viewpoint, the boundary arises from the strictly semistable locus. Understanding it naturally breaks into three tasks: (i) identify polystable points (closed orbits) and construct explicit normal forms representing them; (ii) understand the boundary families they generate, together with the corresponding parameter spaces and dimensions; and (iii) extract geometrically meaningful information by analyzing the singular loci $\mathrm{Sing}(X)$ of these boundary representatives. The main contribution of this paper is to carry out these steps as explicitly as possible for cubic fivefolds, and to provide strong theoretical and computational evidence for the remaining structural assertions.
Irreducible components and normal forms. We first construct explicit polystable (closed-orbit) normal forms and the corresponding boundary families $\Phi_{k}$ ($1\leq k\leq 21$), together with their parameter spaces and dimensions. Assuming Conjecture 3.5 (which is strongly supported by extensive numerical experiments; see Remark 3.6), these families give exactly the $21$ irreducible components of the strictly semistable locus; see Theorem A. In any case, the construction provides a concrete coordinate-level model for the boundary representatives and their degenerations.
Singular loci on the boundary. Next, we determine the singular locus of each closed-orbit representative $X_{k}$. We show that $\mathrm{Sing}(X_{k})$ consists of a one-dimensional part (when present) together with finitely many isolated singular points. Moreover, the positive-dimensional part is a union of low-degree curves—namely lines, smooth conics, complete intersections of two quadrics, or an elliptic quartic—whereas isolated singularities occur only in finitely many families. Crucially, all isolated boundary singularities are quasi-homogeneous and fall into precisely six analytic types. These statements are summarized in Theorem C.
The critical minimal exponent threshold and extremal singularities. A second theme of the paper is that these six isolated singularity types are not accidental by-products of the classification: they capture the exact boundary of stability in a sharp sense coming from singularity theory. Building on Park’s framework relating (local) minimal exponents to GIT stability for hypersurfaces, we show that every isolated boundary singularity appearing here has local minimal exponent equal to
$\frac{7}{3}=\frac{n+1}{d}\qquad\text{for }(n,d)=(6,3),$
that is, it realizes the equality case at the critical threshold. In Park’s terminology, this is precisely the borderline regime where semistability can still occur while stability is no longer forced. Motivated by this, we package these six analytic types as a distinguished class and refer to them as extremal cubic fivefold singularities; see Definition 5.30 and Theorem C.
Adjacency of boundary families. Finally, we analyze adjacency relations among the boundary families arising from the closed-orbit normal forms. In a setting with many boundary families, the “connectivity” and “proximity” of families control the local structure of the moduli space and are closely tied to wall-crossing phenomena. Using Kirwan’s stratification and a detailed analysis of degenerations among the closed-orbit representatives, we show that the nontrivial adjacencies are extremely sparse: there are exactly eight adjacent pairs, as stated in Theorem D. This provides a minimal combinatorial skeleton for the boundary, and serves as a foundation for further investigations of local deformation models and related compactifications.
Idea of the proofs. Our approach combines standard GIT techniques—$1$-parameter-subgroup limits, computations of stabilizers and centralizers, and normalizations via residual group actions—with explicit calculations tailored to cubic fivefolds. For each boundary family we construct concrete normal forms and read off the structure of the singular locus directly from these models. The minimal exponent statement is obtained by applying Park’s criterion to our explicit list of isolated singularities and verifying the critical equality $\widetilde{\alpha}_{x}(X)=\frac{7}{3}$. The adjacency results follow from comparing degenerations along Kirwan strata and tracking how the boundary normal forms specialize.
Theorem A (Boundary families and components). We construct $21$ explicitly described $\mathrm{SL}(7)$-invariant families
$\Phi_{1},\Phi_{2},\ldots,\Phi_{21}\ \subset\ \mathbb{P}(W)^{ss}\setminus\mathbb{P}(W)^{s}\qquad(W=\mathrm{Sym}^{3}\mathbb{C}^{7}),$
together with explicit closed-orbit (polystable) representatives in normal form; see Table 2 (Section 4). Assuming Conjecture 3.5 (strongly supported by extensive numerical experiments; see Remark 3.6), these families are pairwise distinct and give exactly the $21$ irreducible components of $\mathbb{P}(W)^{ss}\setminus\mathbb{P}(W)^{s}$.
Theorem B (Closed-orbit test and normal-form construction). Let $G=\mathrm{SL}(7)$ act on $W=\mathrm{Sym}^{3}\mathbb{C}^{7}$. For each boundary family $\Phi_{k}$ $(k=1,\dots,21)$, choose a $1$-PS limit with stabilizer $H\subset G$ and consider the induced effective action of the centralizer $C_{G}(H)$ on $W^{H}$. Then closed orbits occurring in $\Phi_{k}$ can be represented by elements $\phi_{k}\in W^{H}$, and $\phi_{k}$ can be normalized to a normal form by the residual $C_{G}(H)$-action. Moreover, the closedness of $C_{G}(H)\cdot\phi_{k}$ (hence of $G\cdot\phi_{k}$) is decided uniformly by a dichotomy: in the toric case by the convex-hull criterion for weights, and otherwise by the Casimiro–Florentino criterion via a positive linear identity. In particular, Table 2 lists explicit closed-orbit normal forms associated with all $\Phi_{k}$.
Theorem C (Singularities on the boundary). Let $\phi^{\mathrm{nf}}_{k}$ be the closed–orbit representative associated with $\Phi_{k}$ (Table 2) and set $X_{k}:=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$. Then $\mathrm{Sing}(X_{k})$ is explicitly determined (Table 3): it is a union of low–degree curves (lines, smooth conics, complete intersections of two quadrics, or an elliptic quartic), possibly together with isolated points. Isolated points occur exactly for $k\in\{1,5,6,8,9,11,16,19,21\}$; they are quasi–homogeneous and fall into six analytic types (Definitions 5.1–5.6), including wild examples, and satisfy the extremal equality
$\widetilde{\alpha}_{x}(X_{k})=\frac{7}{3}=\frac{n+1}{d}\qquad\text{for }(n,d)=(6,3).$
Theorem D (Adjacency via wall-crossing). Let $\overline{\Phi}_{k}\subset\mathbb{P}(W)//\mathrm{SL}(7)$ denote the image of the boundary family $\Phi_{k}$, and write $\Sigma^{\mathrm{ss}}:=\mathbb{P}(W)^{\mathrm{ss}}\setminus\mathbb{P}(W)^{\mathrm{s}}$ for the strictly semistable locus. We record adjacencies of boundary components in the sense of Kirwan’s partial desingularization, i.e. via wall-crossings between Kirwan strata. Then the adjacency graph on $\{\overline{\Phi}_{k}\}_{k=1}^{21}$ has exactly eight edges, namely
$\{\overline{\Phi}_{1},\overline{\Phi}_{7}\},\ \{\overline{\Phi}_{2},\overline{\Phi}_{6}\},\ \{\overline{\Phi}_{3},\overline{\Phi}_{12}\},\ \{\overline{\Phi}_{8},\overline{\Phi}_{19}\},\ \{\overline{\Phi}_{9},\overline{\Phi}_{15}\},\ \{\overline{\Phi}_{10},\overline{\Phi}_{17}\},\ \{\overline{\Phi}_{11},\overline{\Phi}_{21}\},\ \{\overline{\Phi}_{14},\overline{\Phi}_{20}\}.$
In particular, the remaining five components
$\overline{\Phi}_{4},\ \overline{\Phi}_{5},\ \overline{\Phi}_{13},\ \overline{\Phi}_{16},\ \overline{\Phi}_{18}$
are isolated vertices (i.e. adjacent to no other component). Equivalently, for $i\neq j$ one has
$\overline{\Phi}_{i}\cap\overline{\Phi}_{j}\cap\Sigma^{\mathrm{ss}}\neq\varnothing$
if and only if $\{i,j\}$ is one of the eight unordered pairs listed above. See Theorem 6.1 and [Kir84, DH98, Tha96].
Ideas and methods
Our starting point is a convex-geometric analysis of Hilbert–Mumford weights for the natural $\mathrm{SL}(7)$-action on $W=\mathrm{Sym}^{3}(\mathbb{C}^{7})$ [MFK94, Dol03]. Fix a maximal torus $T\subset\mathrm{SL}(7)$, and for a cubic $f\in W$ write $\mathrm{Supp}(f)\subset I=\mathbb{Z}^{7}_{(3)}\cap\mathbb{Z}^{7}_{\geq 0}$ for its set of exponent vectors. We enumerate the maximal $T$-strictly semistable supports $I(r)_{\geq 0}$ whose convex hull contains the barycenter $\eta=(3/7,\dots,3/7)$. An algorithmic search over the faces of $\mathrm{Conv}(I)$ containing $\eta$ produces $23$ candidates up to permutation; exactly one is $T$-unstable and is discarded, leaving $22$ families with respect to $T$. Passing to the full $\mathrm{SL}(7)$-action, there is a single residual identification $f_{21}\sim f_{22}$. Extensive computations strongly suggest that no further identifications occur; see Conjecture 3.5 and Remark 3.6. This leads to the desired list of 2121 $\mathrm{SL}(7)$-inequivalent boundary families (Algorithm 2.2, Proposition 3.4, and Theorem 3.7).
For each family, we obtain closed $\mathrm{SL}(7)$-orbits by taking appropriate one-parameter subgroup limits and applying Luna’s centralizer reduction [Lun75]. If the centralizer is a torus, we apply the convex-hull criterion; if it is non-toric, we use the Casimiro–Florentino criterion [CF12]. This uniform dichotomy produces explicit closed-orbit normal forms, from which the family dimensions follow (Section 4; see also Table 2).
Organization of the paper
Section 1 recalls the numerical criterion for (semi)stability in convex-geometric language. Section 2 enumerates maximal strictly semistable supports for the maximal torus $T$ (Algorithm 2.2). Section 3 passes from $T$-data to the $\mathrm{SL}(7)$-action, eliminates redundancies, and arrives at the $21$ families $f_{1},\dots,f_{21}$ Section 4 constructs closed-orbit representatives, proves polystability (via the convex-hull or Casimiro–Florentino criterion), and records component dimensions. Section 5 computes singular loci and lists isolated types, including wild examples. Section 6 records adjacencies as wall-crossings in Kirwan’s stratification. Table 2 presents a compact summary of the normal forms and dimensions.
We present below a table comparing cubic threefolds and fourfolds.
Table 1: Cubic threefolds and fourfolds: moduli dimension, GIT boundary components, and stability-singularity profile.
Cubic threefolds in $\mathbb{P}^{4}$ (See [Yok02].)
Moduli dimension. Number of monomials $\binom{5+3-1}{3}=35$, projective dimension $34$; $\dim\mathrm{PGL}_{5}=24$. Thus $\dim\mathcal{M}^{\mathrm{GIT}}=34-24=\mathbf{10}$.
Components of the GIT boundary (polystable closed orbits). Two irreducible components: (i) a $\mathbb{P}^{1}$-family $\{\phi_{\alpha,\beta}\}$ with parameter $\alpha:\beta]$, and (ii) the isolated point represented by $\phi=vwz+x^{3}+y^{3}$. Singularity profile (stability). Stable $\Longleftrightarrow$ only double points of type $A_{n}$ with $n\leq 4$. Semistable $\Longleftrightarrow$ only $A_{n}$ $(n\leq 5)$, $D_{4}$, or $A_{\infty}$ double points. Inside the $\mathbb{P}^{1}$-component, the special member with $\alpha^{2}=4\beta$ is the secant threefold (the singular locus is the rational normal curve).
Adjacency/closures. Both boundary components consist of closed orbits; the $\mathbb{P}^{1}$-component and the isolated point are distinct boundary components (there is no specialization of one to the other).
Cubic fourfolds in $\mathbb{P}^{5}$ (See [Laz09, Yok08, Huy23].)
Moduli dimension. Number of monomials $\binom{6+3-1}{3}=56$, projective dimension $55$; $\dim\mathrm{PGL}_{6}=35$. Thus $\dim\mathcal{M}^{\mathrm{GIT}}=\mathbf{20}$.
Boundary (closed orbit) families and their dimensions. There are six types, denoted [C.1][C.1]-$[C.6]$, of respective dimensions $1,2,3,1,1,0$. A convenient set of normal forms is: [C.1] $u\,q_{1}(w,x,y,z)+v\,q_{2}(w,x,y,z)$ with $V(u,v,q_{1},q_{2})$ smooth. [C.2] $u(xy+xz+yz+\alpha z^{2})+v^{2}x+w^{2}y+2vwz$ (generic $\alpha$). [C.3] $uy^{2}+v^{2}z+l_{1}(w,x)\,uz+2\,l_{2}(w,x)\,vy+c(w,x)$ with $l_{2}^{2}\nmid c$, $l_{1}\nmid c$. [C.4] $uvw+c(x,y,z)$ with $V(u,v,w,c)$ smooth. [C.5] $\alpha\,uy^{2}+v^{2}z+w^{2}x-uxz+2vwy$ ($\alpha\neq 0$). [C.6] $uvw+xyz$. Stability via singularities. A cubic fourfold with only isolated simple (ADE) singularities is GIT stable. Conversely, non-stability occurs if any of the following conditions holds:  (1) $\mathrm{Sing}(X)$ contains a conic; (2) $\mathrm{Sing}(X)$ contains a line; (3) $\mathrm{Sing}(X)$ contains the intersection of two quadrics; (4) $X$ has a double point of rank $\leq 2$; (5) a rank 3 double point with a hyperplane section whose singular locus is a line with ranks $\leq 2$ along it; (6) a rank 3 double point whose tangent-cone singular locus is a $2$-plane in $X$.
Adjacency (specialization) among boundary strata. If we denote the families $[C.1]$ to $[C.6]$ by $C_{1},S_{2},V_{3},C_{4},C_{5},P_{6}$, then $P_{6}\subset\overline{S_{2}}\cap\overline{V_{3}},\qquad P_{6}\in\overline{C_{1}}\cap\overline{C_{4}}\cap\overline{C_{5}}.$ Cubic fivefolds in $\mathbb{P}^{6}$
Moduli dimension. Number of monomials $\binom{7+3-1}{3}=84$, projective dimension $83$; $\dim\mathrm{PGL}_{7}=48$. Thus $\dim\mathcal{M}^{\mathrm{GIT}}=\textbf{35}$.
Components of the GIT boundary (strictly semistable locus).
Remark 0.1 (Cubic threefolds and fourfolds rephrased via minimal exponents).
The classical GIT classifications of cubic threefolds and cubic fourfolds admit a succinct reinterpretation in terms of the (global) minimal exponent $\widetilde{\alpha}$ in the sense of Park. Cubic threefolds. Let $X\subset\mathbb{P}^{4}$ be a cubic threefold. The critical value in Park’s criterion is $\frac{n+1}{d}=\frac{5}{3}$. Park’s theorem gives the sufficient conditions $\widetilde{\alpha}(X)>\frac{5}{3}\Rightarrow X$ is GIT stable and $\widetilde{\alpha}(X)\geq\frac{5}{3}\Rightarrow X$ is GIT semistable [Par25, Theorem A]. In fact, the explicit GIT analysis of cubic threefolds (see [Yok02]) shows that $X$ is stable $\iff\widetilde{\alpha}(X)>\frac{5}{3}$ $X$ is semistable and not GIT-equivalent to the chordal cubic $\mathcal{T}$ $\iff\widetilde{\alpha}(X)\geq\frac{5}{3}$. The remaining strictly semistable closed orbit is the chordal cubic $\mathcal{T}$, which is GIT polystable and satisfies $\widetilde{\alpha}(\mathcal{T})=\frac{3}{2}$ [Par25, Remark 7.6(2)]. Cubic fourfolds. Let $X\subset\mathbb{P}^{5}$ be a cubic fourfold. Here $\frac{n+1}{d}=2$. The classical results of Yokoyama–Laza (see [Yok08, Laz09]) imply that a cubic fourfold with only isolated simple (ADE) singularities is GIT stable, hence in particular $\widetilde{\alpha}(X)>2$. Moreover, away from the one-parameter family $\chi$ of GIT polystable cubics, semistability is characterized by the threshold $\widetilde{\alpha}(X)\geq 2$: $X$ semistable and not GIT-equivalent to a member of $\chi$ $\Rightarrow\widetilde{\alpha}(X)\geq 2\Rightarrow$ $X$ semistable. [Par25, Remark 7.6(3)]. The exceptional family $\chi$ forms the polystable locus with $\widetilde{\alpha}(X)<2$: the secant to the Veronese surface $\omega\in\chi$ has $\widetilde{\alpha}(\omega)=\frac{3}{2}$, while for all other $X\in\chi\setminus\{\omega\}$ one has $\widetilde{\alpha}(X)=\frac{11}{6}$ [Par25, Remark 7.6(3)]. These lower-dimensional cases provide useful prototypes for cubic fivefolds, where the critical value is $\frac{n+1}{d}=\frac{7}{3}$ and our boundary isolated singularities realize the equality case at this threshold.
Scripts used in this paper
This article is a preliminary arXiv version intended to quickly disseminate the main results. The scripts used to produce the computational data (including Tables 2–4) will be made publicly available in the next version of this preprint. In the meantime, the corresponding outputs are fully recorded in the paper.
Use of AI Tools
The author used ChatGPT and Gemini as auxiliary tools to assist in exploring and organizing computations, including the derivation of candidate normal forms in Section 4 and candidate normal forms of isolated singularities in Section 5. They were also used to support coding related to these computations (e.g., drafting and debugging scripts). In addition, these tools were used to improve exposition, organization, and internal consistency of the manuscript through iterative discussions. Under the author’s supervision, they were used for language editing in English. All mathematical statements and results were ultimately checked and verified by explicit calculations carried out by the author, who takes full responsibility for the content.
Acknowledgments
The author would like to thank Professor T. Terasoma for his encouragement, advice, and valuable suggestions. He is deeply grateful to his parents for their support during difficult times. The author dedicates this work to the memory of Professor E. Horikawa.
1 Numerical criterion for cubic fivefolds
In this section, we review the numerical criterion for stability or semistability of cubic fivefolds. We use the following notations.
  • Let $\mathbb{C}[x_{0},\cdots,x_{6}]_{3}$ be the set of homogeneous polynomials of degree $3$.
  • For a vector $\mathbf{x}\in\mathbb{Q}^{7}$, $\mathrm{wt}(\mathbf{x})=\sum_{k=0}^{6}x_{k}$ is called the weight of $\mathbf{x}$.
  • We define $\mathbb{Z}^{7}_{\geq 0}=\{\mathbf{x}=(x_{0},x_{1},\cdots,x_{6})\in\mathbb{Z}^{7}|x_{k}\geq 0(k=0,1,\cdots,6)\}$,
    $\mathbb{Z}^{7}_{(d)}=\{\mathbf{x}\in\mathbb{Z}^{7}|\mathrm{wt}(\mathbf{x})=d\}$,
    $I=\mathbb{Z}^{7}_{(3)}\cap\mathbb{Z}^{7}_{\geq 0}$ and it is simply called the simplex.
  • For $\mathbf{r}\in\mathbb{Q}^{7}$, we define $I(\mathbf{r})_{\geq 0}=\{\mathbf{i}\in I|\mathbf{r}\cdot\mathbf{i}\geq 0\}$, $I(\mathbf{r})_{>0}=\{\mathbf{i}\in I|\mathbf{r}\cdot\mathbf{i}>0\}$ and $I(\mathbf{r})_{=0}=\{\mathbf{i}\in I|\mathbf{r}\cdot\mathbf{i}=0\}$, here $\cdot$ denotes the standard inner product of vectors.
  • For a polynomial $f=\sum_{\mathrm{wt}(\mathbf{i})=3}a_{\mathbf{i}}x^{\mathbf{i}}\in\mathbb{C}[x_{0},\cdots,x_{6}]_{3}$, we define the support of $f$ by $\mathrm{Supp}(f)=\{\mathbf{i}\in I|a_{\mathbf{i}}\neq 0\}$
  • We set $\eta=(3/7,3/7,3/7,3/7,3/7,3/7,3/7)\in\mathbb{Q}^{7}$ and it is called the barycenter of the simplex $I$.
  • A vector $\mathbf{r}\in\mathbb{Z}^{7}$ is said to be reduced when there is no integer $\alpha$ such that $|\alpha|\geq 2$ and $\frac{1}{\alpha}\mathbf{r}\in\mathbb{Z}^{7}$
We fix a maximal torus $T$ of $\mathrm{SL}(7)$. Consider a one-parameter subgroup (1-PS for short) $\lambda:\mathbb{G}_{m}\to\mathrm{SL}(7)$ whose image is contained in $T$. For a suitable basis of $\mathbb{C}^{7}$, $\lambda$ can be expressed as a diagonal matrix $\mathrm{diag}(t^{r_{0}},t^{r_{1}},\cdots,t^{r_{6}})$ where $t\neq 0$ is a parameter of $\mathbb{G}_{m}$. Let us choose and fix such basis. Then $\lambda$ corresponds to an element $\mathbf{r}=(r_{0},r_{1},\cdots,r_{6})$ in $\mathbb{Z}^{7}_{(0)}$. We can regard an element of $\mathbb{Z}^{7}_{(0)}$ as a 1-PS of $T$.
Definition 1.1.
Let $s$ be a subset of $I$. We say that $s$ is not stable (resp. unstable) with respect to $T$ when $s\subseteq I(\mathbf{r})_{\geq 0}$ (resp. $s\subseteq I(\mathbf{r})_{>0}$) for some 1-PS $\mathbf{r}$. For $0\neq f\in\mathbb{C}[x_{0},\cdots,x_{6}]_{3}$, we say that $f$ is not stable (resp. unstable) with respect to $T$ when $\mathrm{Supp}(f)\subseteq I$ is not stable (resp. unstable) with respect to $T$.
stablesemistableunstablestrictly semistablenot stable
Figure 1: various concepts of stability
The following theorem is the numerical criterion for stability via the language of convex geometry.
Theorem 1.2.
The cubic fivefold defined by $f\in\mathbb{C}[x_{0},\cdots,x_{6}]_{3}$ is not stable (resp. unstable) if and only if there exists an element $\sigma\in\mathrm{SL}(7)$ such that $f^{\sigma}$ is not stable (resp. unstable) with respect to $T$.
In particular, $f$ is strictly semistable if and only if
There exist $\sigma\in\mathrm{SL}(7)$ such that $f^{\sigma}$ is not stable with respect to $T$, and
For any $\sigma\in\mathrm{SL}(7)$, $f^{\sigma}$ is semistable with respect to $T$.
Proof.
See Theorem 9.1 of [Dol03]. ∎
2 Maximal strictly semistable cubic fivefolds with respect to the maximal torus $T$
In this section, we list the irreducible components corresponding to strictly semistable cubic fivefolds. For this purpose, we list all strictly semistable cubic fivefolds with respect to the maximal torus $T$. To solve this problem, we will consider the set of maximal strictly semistable subsets of $I$. The order in the set of subsets of $I$ is given by inclusion. For this purpose, we list the set of all maximal elements of $\mathcal{S}=\{I(\mathbf{r})_{\geq 0}|\mathbf{r}\in\mathbb{Z}_{(0)}^{7}\}$.
We solve this problem computationally. We need an algorithm which enables us to obtain them in finitely many steps. Before giving such an algorithm, we remark that $I(\mathbf{r})_{\geq 0}$ and $I(\mathbf{r^{\prime}})_{\geq 0}$ might be the same for two different vectors $\mathbf{r},\mathbf{r^{\prime}}\in\mathbb{Z}^{7}_{(0)}$.
Lemma 2.1.
Let $I(\mathbf{r})_{\geq 0}$ be a maximal element of $\mathcal{S}$, where $\mathbf{r}\in\mathbb{Z}^{7}_{(0)}.$ Then there exist $5$ elements $\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{5}\in I$ and a vector $\mathbf{r^{\prime}}\in\mathbb{Z}^{7}_{(0)}$ such that they satisfy the following three conditions:
The vector subspace $W$ of $\mathbb{Q}^{7}$ spanned by $\mathbf{x}_{1},\cdots,\mathbf{x}_{5},\eta$ over $\mathbb{Q}$ has dimension 66
The vector $\mathbf{r^{\prime}}$ is orthogonal to the subspace $W$ of $\mathbb{Q}^{7}$.
$I(\mathbf{r})_{\geq 0}=I(\mathbf{r^{\prime}})_{\geq 0}$
Proof.
Let us put $C=I(r)\cup\{\eta\}$. We consider the convex-hull $\check{C}$ of $C$ in $\mathbb{Q}^{7}$. Let $F$ be a face of $\check{C}$ containing the point $\eta$. There is a normal vector $\mathbf{r}^{\prime}$ of $F$ in $\mathbb{Z}^{7}_{(0)}$ such that $\check{C}\subseteq\{\mathbf{x}\in\mathbb{Q}^{7}|\mathbf{r}^{\prime}\cdot\mathbf{x}\geq 0\}$. We have $\mathrm{wt}(\mathbf{r^{\prime}})=0$ because the hyperplane defined by $\{\mathbf{x}\in\mathbb{Q}^{7}|\mathbf{r}^{\prime}\cdot\mathbf{x}=0\}$ passes through the point $\eta$. By the definition of the faces of a convex set in $\mathbb{Q}^{7}$, we can take 5 points $\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{5}$ from the set $I\cap F$ such that $\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{5},\eta$ are linearly independent over $\mathbb{Q}$. In general we have $I(\mathbf{r})_{\geq 0}\subseteq I(\mathbf{r}^{\prime})_{\geq 0}$, and by the assumption that $I(\mathbf{r})_{\geq 0}$ is maximal in $\mathcal{S}$, we conclude that $I(\mathbf{r})_{\geq 0}=I(\mathbf{r}^{\prime})_{\geq 0}$. ∎
By this lemma, we can determine the set of maximal elements of $\mathcal{S}$ up to permutations of coordinates in finite steps using the following algorithm.
Algorithm 2.2.
Let $\mathcal{F}$ be the set of five different points of $I$. We fix a total order on $\mathcal{F}$. As an initial data, we set $\mathcal{S}^{\prime}=\varnothing$ and $\mathbf{x}=(x_{0},\cdots,x_{5})$ be the minimum element of $\mathcal{F}$. We will modify $\mathcal{S}^{\prime}$ using the following algorithm.
  • Step 1. If the subspace $W$ spanned by $x_{0},\cdots,x_{5},\eta$ of $\mathbb{Q}^{7}$ has dimension $6$ then take a reduced normal vector $\mathbf{r}=(r_{0},\cdots,r_{6})\in\mathbb{Z}^{7}_{(0)}$ of $W$ and go to Step 2, else go to Step 5.
  • Step 2. If $r=(r_{0},\cdots,r_{6})$ satisfies the condition $r_{0}\geq\cdots\geq r_{6}$ or $r_{0}\leq\cdots\leq r_{6}$, then go to Step 3, else go to Step 5.
  • Step 3. If $r_{0}\geq\cdots\geq r_{6}$ (resp. $r_{0}\leq\cdots\leq r_{6}$) add $I(\mathbf{r})$ (resp. $I(\mathbf{-r})$) to $\mathcal{S}^{\prime}$ and go to Step 4.
  • Step 4. Delete all elements of $\mathcal{S}^{\prime}$ that are not maximal in $\mathcal{S}^{\prime}$ and go to Step 5.
  • Step 5. Replace the element $\mathbf{x}$ with the next element if $\mathbf{x}$ is not the maximum element, and go to Step 1. Otherwise, stop the algorithm.
We note that Step 2 removes the $S_{7}$ symmetry on the variables $x_{0},\cdots,x_{6}$. We also note that Step 4 is not essential but serves as technical measure to save memory. After running this algorithm with the aid of a computer, we find 23 elements $I(\mathbf{r}_{1})_{\geq 0},\cdots,I(\mathbf{r}_{23})_{\geq 0}$ in $\mathcal{S}^{\prime}$, where $\mathbf{r}_{k}=(r_{0},\cdots,r_{6})\in\mathbb{Z}^{7}_{(0)}$ is a reduced vector with $r_{0}\geq\cdots\geq r_{6}$. When we compute the convex-hulls of $I(\mathbf{r}_{1})_{\geq 0},\cdots,I(\mathbf{r}_{23})_{\geq 0}$ in $\mathbb{Q}^{7}$, only one of the convex-hulls of $I(\mathbf{r}_{k})_{\geq 0}$ does not contain $\eta$. We denote it as $I(\mathbf{r}_{23})_{\geq 0}$. As $I(\mathbf{r}_{23})_{\geq 0}$ is unstable with respect to $T$, we remove it from the list. Thus, we can conclude that there are 22 maximal strictly semistable cubic fivefolds for the fixed maximal torus $T$. Because of this algorithm, we have the following proposition.
Proposition 2.3.
The set $\mathcal{M}=\{I(\mathbf{r}_{1})_{\geq 0},\cdots,I(\mathbf{r}_{22})_{\geq 0}\}$ is given as follows.
$\mathbf{r}_{1}=(8,3,2,-1,-2,-4,-6)$ $\mathbf{r}_{2}=(6,4,1,-1,-2,-3,-5)$
$\mathbf{r}_{3}=(4,2,1,-1,-1,-2,-3)$ $\mathbf{r}_{4}=(3,2,1,0,-1,-2,-3)$
$\mathbf{r}_{5}=(4,2,1,0,-1,-2,-4)$ $\mathbf{r}_{6}=(5,3,2,1,-1,-4,-6)$
$\mathbf{r}_{7}=(6,4,2,1,-2,-3,-8)$ $\mathbf{r}_{8}=(4,1,1,0,-2,-2,-2)$
$\mathbf{r}_{9}=(2,2,0,0,-1,-1,-2)$ $\mathbf{r}_{10}=(2,1,0,0,-1,-1,-1)$
$\mathbf{r}_{11}=(2,0,0,0,0,-1,-1)$ $\mathbf{r}_{12}=(3,2,1,1,-1,-2,-4)$
$\mathbf{r}_{13}=(2,1,1,0,-1,-1,-2)$ $\mathbf{r}_{14}=(2,2,0,-1,-1,-1,-1)$
$\mathbf{r}_{15}=(2,1,1,0,0,-2,-2)$ $\mathbf{r}_{16}=(2,1,0,0,0,-1,-2)$
$\mathbf{r}_{17}=(1,1,1,0,0,-1,-2)$ $\mathbf{r}_{18}=(1,1,0,0,0,-1,-1)$
$\mathbf{r}_{19}=(2,2,2,0,-1,-1,-4)$ $\mathbf{r}_{20}=(1,1,1,1,0,-2,-2)$
$\mathbf{r}_{21}=(1,1,0,0,0,0,-2)$ $\mathbf{r}_{22}=(1,0,0,0,0,0,-1)$
For example, $I(\mathbf{r}_{1})_{\geq 0}$ is
$I(\mathbf{r}_{1})_{\geq 0}=\{x_{0}^{3},x_{0}^{2}x_{1},x_{0}^{2}x_{2},x_{0}^{2}x_{3},x_{0}^{2}x_{4},x_{0}^{2}x_{5},x_{0}^{2}x_{6},x_{0}x_{1}^{2},x_{0}x_{1}x_{2},x_{0}x_{1}x_{3},x_{0}x_{1}x_{4},\\ x_{0}x_{1}x_{5},x_{0}x_{1}x_{6},x_{0}x_{2}^{2},x_{0}x_{2}x_{3},x_{0}x_{2}x_{4},x_{0}x_{2}x_{5},x_{0}x_{2}x_{6},x_{0}x_{3}^{2},x_{0}x_{3}x_{4},x_{0}x_{3}x_{5},x_{0}x_{3}x_{6},\\ x_{0}x_{4}^{2},x_{0}x_{4}x_{5},x_{0}x_{4}x_{6},x_{0}x_{5}^{2},x_{1}^{3},x_{1}^{2}x_{2},x_{1}^{2}x_{3},x_{1}^{2}x_{4},x_{1}^{2}x_{5},x_{1}^{2}x_{6},x_{1}x_{2}^{2},x_{1}x_{2}x_{3},x_{1}x_{2}x_{4},\\ x_{1}x_{2}x_{5},x_{1}x_{3}^{2},x_{1}x_{3}x_{4},x_{2}^{3},x_{2}^{2}x_{3},x_{2}^{2}x_{4},x_{2}^{2}x_{5},x_{2}x_{3}^{2}\}.$
Here we use the notation $x_{0}^{i_{0}}x_{1}^{i_{1}}\cdots x_{6}^{i_{6}}$ for an element $(i_{0},i_{1},\cdots,i_{6})\in\mathbb{Z}^{7}_{(3)}$ in order to save space.
Remark 2.4.
The following vectors can serve as $r_{23}$, i.e., there are several vectors that yield the set $I(\mathbf{r}_{23})_{\geq 0}$. For example, we can take $\mathbf{r}_{23}=(8,5,3,2,-4,-4,-10)$.
Remark 2.5.
The algorithm in this section has been comprehensively generalized by [GMMS23].
3 21 maximal strictly semistable cubic fivefolds under action of $\mathrm{SL}(7)$
An element $I(\mathbf{r}_{k})_{\geq 0}$ of $\mathcal{M}$ represents a family of cubic fivefolds whose defining polynomial’s support is contained in $I(\mathbf{r}_{k})_{\geq 0}$. In this section, we analyze the inclusion relations among $I(\mathbf{r}_{k})_{\geq 0}$ under the action of $\mathrm{SL}(7)$. Let $f_{k}$ be a generic polynomial whose support is $I(\mathbf{r}_{k})_{\geq 0}$. $(k=1,2,\cdots,22)$. If we express $f_{k}$ directly, it becomes too long, so we introduce notation.
Definition 3.1.
The symbols $c,q,l,\alpha$ stand for a cubic form, a quadratic form, a linear form, and a constant term, respectively. Similarly, the symbols $q_{i},l_{i},\alpha_{i}$ denote the $i$-th quadratic form, a linear form, and a constant term, respectively.
The following theorem is a direct consequence of the list in Proposition 2.3.
Theorem 3.2.
Using the above notations, the generic polynomials of $f_{1},\cdots,f_{22}$ are the following forms.
  • $f_{1}=c(x_{0},x_{1},x_{2})+q_{1}(x_{0},x_{1},x_{2})x_{3}+l_{1}(x_{0},x_{1},x_{2})x_{3}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1})x_{3}\}x_{4}+\alpha_{1}x_{0}x_{4}^{2}+\{q_{3}(x_{0},x_{1},x_{2})+x_{0}l_{3}(x_{3},x_{4})\}x_{5}+\alpha_{2}x_{0}x_{5}^{2}+\{q_{4}(x_{0},x_{1})+x_{0}l_{4}(x_{2},x_{3},x_{4})\}x_{6}$
  • $f_{2}=c(x_{0},x_{1},x_{2})+q_{1}(x_{0},x_{1},x_{2})x_{3}+l_{1}(x_{0},x_{1})x_{3}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1})x_{3}\}x_{4}+l_{3}(x_{0},x_{1})x_{4}^{2}+\{q_{3}(x_{0},x_{1})+l_{4}(x_{0},x_{1})x_{2}+l_{5}(x_{0},x_{1})x_{3}+\alpha_{1}x_{0}x_{4}\}x_{5}+\alpha_{2}x_{0}x_{5}^{2}+\{q_{4}(x_{0},x_{1})+l_{6}(x_{0},x_{1})x_{2}+\alpha_{3}x_{0}x_{3}\}x_{6}$
  • $f_{3}=c(x_{0},x_{1},x_{2})+q_{1}(x_{0},x_{1},x_{2})x_{3}+l_{1}(x_{0},x_{1})x_{3}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1})x_{3}\}x_{4}+l_{3}(x_{0},x_{1})x_{4}^{2}+\{q_{3}(x_{0},x_{1},x_{2})+x_{0}l_{4}(x_{3},x_{4})\}x_{5}+\alpha_{1}x_{0}x_{5}^{2}+\{q_{4}(x_{0},x_{1})+l_{5}(x_{0},x_{1})x_{2}+x_{0}l_{6}(x_{3},x_{4})\}x_{6}$
  • $f_{4}=c(x_{0},x_{1},x_{2})+\{q_{1}(x_{0},x_{1},x_{2})+l_{1}(x_{0},x_{1},x_{2})x_{3}+\alpha_{1}x_{3}^{2}\}x_{3}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1},x_{2})x_{3}\}x_{4}+l_{3}(x_{0},x_{1})x_{4}^{2}+\{q_{3}(x_{0},x_{1},x_{2})+l_{4}(x_{0},x_{1})x_{3}+\alpha_{2}x_{0}x_{4}\}x_{5}+\{q_{4}(x_{0},x_{1})+l_{5}(x_{0},x_{1})x_{2}+\alpha_{3}x_{0}x_{3}\}x_{6}$
  • $f_{5}=c(x_{0},x_{1},x_{2},x_{3})+\{q_{1}(x_{0},x_{1},x_{2})+l_{1}(x_{0},x_{1},x_{2})x_{3}\}x_{4}+l_{2}(x_{0},x_{1})x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{3}(x_{0},x_{1})x_{3}+\alpha_{1}x_{0}x_{4}\}x_{5}+\alpha_{3}x_{0}x_{5}^{2}+\{q_{3}(x_{0},x_{1})+\alpha_{4}x_{0}x_{2}+\alpha_{2}x_{0}x_{3}\}x_{6}$
  • $f_{6}=c(x_{0},x_{1},x_{2},x_{3})+q_{1}(x_{0},x_{1},x_{2},x_{3})x_{4}+l_{1}(x_{0},x_{1},x_{2})x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1})x_{3}+\alpha_{1}x_{0}x_{4}\}x_{5}+\{q_{3}(x_{0},x_{1})+x_{0}l_{3}(x_{2},x_{3})\}x_{6}$
  • $f_{7}=c(x_{0},x_{1},x_{2},x_{3})+q_{1}(x_{0},x_{1},x_{2},x_{3})x_{4}+l_{1}(x_{0},x_{1})x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1},x_{2})x_{3}+\alpha_{1}x_{0}x_{4}\}x_{5}+\alpha_{2}x_{0}x_{5}^{2}+\{q_{3}(x_{0},x_{1})+\alpha_{3}x_{0}x_{2}\}x_{6}$
  • $f_{8}=c(x_{0},x_{1},x_{2},x_{3})+\{q_{1}(x_{0},x_{1},x_{2})+\alpha_{1}x_{0}x_{3}\}x_{4}+\alpha_{2}x_{0}x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+\alpha_{3}x_{0}x_{3}+\alpha_{4}x_{0}x_{4}\}x_{5}+\alpha_{5}x_{0}x_{5}^{2}+\{q_{3}(x_{0},x_{1},x_{2})+\alpha_{6}x_{0}x_{3}+\alpha_{7}x_{0}x_{4}+\alpha_{8}x_{0}x_{5}\}x_{6}+\alpha_{9}x_{0}x_{6}^{2}$
  • $f_{9}=c(x_{0},x_{1},x_{2},x_{3})+\{q_{1}(x_{0},x_{1})+l_{1}(x_{0},x_{1})x_{2}+l_{2}(x_{0},x_{1})x_{3}\}x_{4}+l_{3}(x_{0},x_{1})x_{4}^{2}+\{q_{2}(x_{0},x_{1})+l_{4}(x_{0},x_{1})x_{2}+l_{5}(x_{0},x_{1})x_{3}+l_{6}(x_{0},x_{1})x_{4}\}x_{5}+l_{7}(x_{0},x_{1})x_{5}^{2}+\{q_{3}(x_{0},x_{1})+l_{8}(x_{0},x_{1})x_{2}+l_{9}(x_{0},x_{1})x_{3}\}x_{6}$
  • $f_{10}=c(x_{0},x_{1},x_{2},x_{3})+\{q_{1}(x_{0},x_{1})+l_{1}(x_{0},x_{1})x_{2}+l_{2}(x_{0},x_{1})x_{3}\}x_{4}+\alpha_{1}x_{0}x_{4}^{2}+\{q_{2}(x_{0},x_{1})+l_{3}(x_{0},x_{1})x_{2}+l_{4}(x_{0},x_{1})x_{3}\}x_{5}+\alpha_{2}x_{0}x_{5}^{2}+\{q_{3}(x_{0},x_{1})+l_{5}(x_{0},x_{1})x_{2}+l_{6}(x_{0},x_{1})x_{3}\}x_{6}+\alpha_{3}x_{0}x_{6}^{2}+\alpha_{4}x_{0}x_{4}x_{5}+\alpha_{5}x_{0}x_{4}x_{6}+\alpha_{6}x_{0}x_{5}x_{6}$
  • $f_{11}=c(x_{0},x_{1},x_{2},x_{3},x_{4})+x_{0}l_{1}(x_{0},x_{1},x_{2},x_{3},x_{4})x_{5}+\alpha_{1}x_{0}x_{5}^{2}+x_{0}l_{2}(x_{0},x_{1},x_{2},x_{3},x_{4})x_{6}+\alpha_{2}x_{0}x_{6}^{2}+\alpha_{3}x_{0}x_{5}x_{6}$
  • $f_{12}=c(x_{0},x_{1},x_{2},x_{3})+q_{1}(x_{0},x_{1},x_{2},x_{3})x_{4}+l_{1}(x_{0},x_{1})x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2},x_{3})+\alpha_{1}x_{0}x_{4}\}x_{5}+\{q_{3}(x_{0},x_{1})+x_{0}l_{2}(x_{2},x_{3})\}x_{6}$
  • $f_{13}=c(x_{0},x_{1},x_{2},x_{3})+\{q_{1}(x_{0},x_{1},x_{2})+l_{1}(x_{0},x_{1},x_{2})x_{3}\}x_{4}+\alpha_{1}x_{0}x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{2}(x_{0},x_{1},x_{2})x_{3}+\alpha_{2}x_{0}x_{4}\}x_{5}+\alpha_{3}x_{0}x_{5}^{2}+\{q_{3}(x_{0},x_{1},x_{2})+\alpha_{4}x_{0}x_{3}\}x_{6}$
  • $f_{14}=\alpha x_{2}^{3}+x_{0}q_{1}(x_{0},\dots,x_{6})+x_{1}q_{2}(x_{1},\dots,x_{6})$
  • $f_{15}=c(x_{0},x_{1},x_{2},x_{3},x_{4})+\{x_{0}l_{1}(x_{0},x_{1},x_{2},x_{3},x_{4})+q_{1}(x_{1},x_{2})\}x_{5}+\{x_{0}l_{2}(x_{0},x_{1},x_{2},x_{3},x_{4})+q_{2}(x_{1},x_{2})\}x_{6}$
  • $f_{16}=c(x_{0},x_{1},x_{2},x_{3},x_{4})+\{q_{1}(x_{0},x_{1})+l_{1}(x_{0},x_{1})x_{2}+l_{2}(x_{0},x_{1})x_{3}+l_{3}(x_{0},x_{1})x_{4}\}x_{5}+\alpha_{1}x_{0}x_{5}^{2}+\{q_{2}(x_{0},x_{1})+x_{0}l_{4}(x_{2},x_{3},x_{4})\}x_{6}$
  • $f_{17}=c(x_{0},x_{1},x_{2},x_{3},x_{4})+\{q_{1}(x_{0},x_{1},x_{2})+l_{1}(x_{0},x_{1},x_{2})l_{2}(x_{3},x_{4})\}x_{5}+q_{2}(x_{0},x_{1},x_{2})x_{6}$
  • $f_{18}=c(x_{0},x_{1},x_{2},x_{3},x_{4})+\{x_{0}l_{1}(x_{0},x_{1},x_{2},x_{3},x_{4})+x_{1}l_{2}(x_{0},x_{1},x_{2},x_{3},x_{4})\}x_{5}+\{x_{0}l_{3}(x_{0},x_{1},x_{2},x_{3},x_{4})+x_{1}l_{4}(x_{0},x_{1},x_{2},x_{3},x_{4})\}x_{6}$
  • $f_{19}=c(x_{0},x_{1},x_{2},x_{3})+\{q_{1}(x_{0},x_{1},x_{2})+l_{1}(x_{0},x_{1},x_{2})x_{3}\}x_{4}+l_{2}(x_{0},x_{1},x_{2})x_{4}^{2}+\{q_{2}(x_{0},x_{1},x_{2})+l_{3}(x_{0},x_{1},x_{2})x_{3}+l_{4}(x_{0},x_{1},x_{2})x_{4}\}x_{5}+l_{5}(x_{0},x_{1},x_{2})x_{5}^{2}+q_{3}(x_{0},x_{1},x_{2})x_{6}$
  • $f_{20}=c(x_{0},x_{1},x_{2},x_{3},x_{4})+q_{1}(x_{0},x_{1},x_{2},x_{3})x_{5}+q_{2}(x_{0},x_{1},x_{2},x_{3})x_{6}$
  • $f_{21}=c(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})+q(x_{0},x_{1})x_{6}$
  • $f_{22}=c(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})+x_{0}l(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})x_{6}$
For an element $\sigma$ in $\mathrm{SL}(7)$ and $\mathbb{J}\subseteq I$, we set $\mathbb{J}^{\sigma}=\cup_{f}\mathrm{Supp}(f^{\sigma})$, where $f$ runs through all polynomials with $\mathrm{Supp}(f)\subseteq\mathbb{J}$.
Definition 3.3.
We denote
$I(\mathbf{r}_{k})_{\geq 0}\subseteq I(\mathbf{r}_{l})_{\geq 0}\ \mathrm{mod}\ \mathrm{SL}(7)$
when there exists $\sigma\in\mathrm{SL}(7)$ such that $I(\mathbf{r}_{k})_{\geq 0}^{\sigma}\subseteq I(\mathbf{r}_{l})_{\geq 0}$ and say that $I(\mathbf{r}_{k})_{\geq 0}$ is included in $I(\mathbf{r}_{l})_{\geq 0}$ modulo $\mathrm{SL}(7)$.
We construct a smaller subset $\mathcal{M}^{\prime}$ of $\mathcal{M}$ such that $(1)$ any element $I(\mathbf{r}_{k})_{\geq 0}$ in $\mathcal{M}$ is included in some element $I(\mathbf{r}_{l})_{\geq 0}$ in $\mathcal{M}^{\prime}$ mod $\mathrm{SL}(7)$, $(2)$ any element $I(\mathbf{r}_{k})_{\geq 0}$ in $\mathcal{M}^{\prime}$ is not included in any other $I(\mathbf{r}_{l})_{\geq 0}$ in $\mathcal{M}^{\prime}$ mod $\mathrm{SL}(7)$ $(1\leq l\leq 22)$
Proposition 3.4.
There are two relations
  • $I(\mathbf{r}_{21})_{\geq 0}\subseteq I(\mathbf{r}_{22})_{\geq 0}\ \mathrm{mod}\ \mathrm{SL}(7)$
  • $I(\mathbf{r}_{22})_{\geq 0}\subseteq I(\mathbf{r}_{21})_{\geq 0}\ \mathrm{mod}\ \mathrm{SL}(7)$.
Proof.
  • $f_{22}=c(x_{0},\cdots,x_{5})+x_{0}l(x_{0},\cdots,x_{5})x_{6}$
  • $\equiv c(x_{0},\cdots,x_{5})+x_{0}l(x_{0},x_{1})x_{6}$
  • $\equiv c(x_{0},\cdots,x_{5})+q(x_{0},x_{1})x_{6}$
  • $f_{21}$
Here $\equiv$ means equality after an $\mathrm{SL}(7)$ change of coordinates. ∎
From this proposition, we can remove $f_{22}$ from the list. Thus, we obtain a list of $21$ types of cubic fivefolds.
Conjecture 3.5.
For any $1\leq k,l\leq 21$ with $k\neq l$, there is no element $g\in\mathrm{SL}(7)$ such that
$g\cdot I(\mathbf{r}_{k})_{\geq 0}\subseteq I(\mathbf{r}_{l})_{\geq 0}.$
Remark 3.6 (Computational evidence).
Extensive numerical experiments strongly suggest that, for any $k\neq l$, there is no $g\in\mathrm{SL}(7)$ satisfying
$g\cdot I(\mathbf{r}_{k})_{\geq 0}\subseteq I(\mathbf{r}_{l})_{\geq 0}.$
Theorem 3.7.
Assuming Conjecture 3.5, the strictly semistable locus
$\mathbb{P}(W)^{ss}\setminus\mathbb{P}(W)^{s}$
has 2121 irreducible components, represented by $f_{1},\dots,f_{21}$.
4 Closed orbits of 21 families
In this section, we find the closed orbits in the 21 families of strictly semistable cubic fivefolds. We define $\lambda_{k}(t)\colon\mathbb{G}_{m}\to\mathrm{SL}(7)$ as 1-PSs corresponding to $\mathbf{r}_{k}$.
We shall use the following convex–geometric criterion repeatedly in this section, so we record it at the outset.
For the following theorem, see [PV94].
Theorem 4.1 (Convex-hull criterion).
Let $T$ be an algebraic torus acting linearly on a finite-dimensional vector space, $V$, and let $v\in V$. Subsequently, the following conditions are equivalent:
  1. (1)
    the $T$-orbit $T\cdot v$ is closed in $V$;
  2. (2)
    0 is an interior point of the convex-hull of $\mathrm{Supp}(v)$ in $X(T)_{\mathbb{R}}$.
Here, $\mathrm{Supp}(v)$ denotes the set of $T$-weights that occur in $v$, and $X(T)_{\mathbb{R}}$ is the real vector space spanned by the character lattice $X(T)$.
The following series of definitions and theorems will be extremely useful for determining polystability in cases where the centralizer is not a torus and the convex-hull criterion cannot be applied. Because they will be used repeatedly from this point on, we state them here beforehand.
Notation 4.2.
Let $G$ be a reductive algebraic group, and let $X$ be an affine variety equipped with an algebraic $G$-action.
  • $(1)$
    We denote by $Y(G)$ the set of one-parameter subgroups of $G$.
  • $(2)$
    When $x\in X$, we put $\Lambda_{x}=\{\lambda(t)\in Y(G)\colon\lim_{t\to 0}\lambda(t)\cdot x\text{ exists}\}$.
  • $(3)$
    When $\lambda\in Y(G)$, we define $P(\lambda)=\{g\in G\colon\lim_{t\to 0}\lambda(t)\cdot g\cdot\lambda(t)^{-1}\text{ exists}\}$. It is a parabolic subgroup of G.
Definition 4.3.
We say that a subset $\Lambda\subset Y(G)$ is symmetric if given any $\lambda\in\Lambda$, there is another 1-PS $\lambda^{\prime}\in\Lambda$ such that $P(\lambda)\cap P(\lambda^{\prime})$ is a Levi subgroup of both $P(\lambda)$ and $P(\lambda^{\prime})$.
We need the following theorem (Theorem 1.1 of [CF12]):
Theorem 4.4.
(Casimiro–Florentino criterion) Let $G$ be a reductive algebraic group and $X$ be an affine $G$-variety. Then, a point $x\in X$ is polystable if and only if $\Lambda_{x}$ is symmetric.
Lemma 4.5 (rank 1 symmetry for Casimiro–Florentino).
Let $G$ be a reductive algebraic group acting on an affine $G$-variety $X$, and let $x\in X$. Assume that there exists a one-parameter subgroup $\mu\in Y(G)$ such that
$\Lambda_{x}\;=\;\{\mu_{k}\mid k\in\mathbb{Z}\}\qquad\mu_{k}(t):=\mu(t^{k}).$
Then $\Lambda_{x}$ is symmetric in the sense of Definition 4.3. Consequently, by the Casimiro–Florentino criterion (Theorem 4.4), the point $x$ is polystable.
Proof.
Fix $k\neq 0$. It is standard that $P(\mu_{k})$ and $P(\mu_{-k})$ are opposite parabolic subgroups and that
$P(\mu_{k})\cap P(\mu_{-k})\;=\;C_{G}\!\bigl(\mu(\mathbb{G}_{m})\bigr)$
which is a Levi subgroup of both $P(\mu_{k})$ and $P(\mu_{-k})$. The inclusion $C_{G}\bigl(\mu(\mathbb{G}_{m})\bigr)\subset P(\mu_{k})\cap P(\mu_{-k})$ is immediate: if $g$ centralizes $\mu(\mathbb{G}_{m})$, then $\mu_{k}(t)\,g\,\mu_{k}(t)^{-1}=g$ for all $t\in\mathbb{G}_{m}$, so both limits as $t\to 0$ and $t\to\infty$ exist. For the reverse inclusion, take $g\in P(\mu_{k})\cap P(\mu_{-k})$ and consider
$\phi:\mathbb{G}_{m}\longrightarrow G,\qquad\phi(t):=\mu_{k}(t)\,g\,\mu_{k}(t)^{-1}.$
By assumption, the limits $\lim_{t\to 0}\phi(t)$ and $\lim_{t\to\infty}\phi(t)$ both exist (the latter because $g\in P(\mu_{-k})$ and $\phi(1/s)=\mu_{-k}(s)\,g\,\mu_{-k}(s)^{-1}$ for $s\to 0$). Hence, $\phi$ extends to a morphism $\tilde{\phi}:\mathbb{P}^{1}\to G$. As $G$ is affine, $\tilde{\phi}$ is constant; in particular,
$\mu_{k}(t)\,g\,\mu_{k}(t)^{-1}=\phi(t)=\phi(1)=g\quad\text{for all }t\in\mathbb{G}_{m}.$
Thus, $g$ centralizes $\mu_{k}(\mathbb{G}_{m})=\mu(\mathbb{G}_{m})$, i.e. $g\in C_{G}\bigl(\mu(\mathbb{G}_{m})\bigr)$. This proves $P(\mu_{k})\cap P(\mu_{-k})=C_{G}\bigl(\mu(\mathbb{G}_{m})\bigr)$.
Based on the hypothesis $\mu_{k},\mu_{-k}\in\Lambda_{x}$, so the requirement in Definition 4.3 is satisfied for every $\lambda=\mu_{k}$. Hence, $\Lambda_{x}$ is symmetric. The last assertion follows from Theorem 4.4 (Casimiro–Florentino). ∎
Convention 4.6.
Throughout this section, we fix the following setup and notation.
(1) Coordinates and the maximal torus. We work on $W=\mathrm{Sym}^{3}\mathbb{C}^{7}$ with homogeneous coordinates $(x_{0},\dots,x_{6})$. We fix the diagonal maximal torus $T\subset SL(7)$ acting by $\mathrm{diag}(\mu_{0},\dots,\mu_{6})$ with $\prod_{i}\mu_{i}=1$. A one-parameter subgroup (1-PS) is written as
$\lambda(t)=\mathrm{diag}\!\bigl(t^{a_{0}},\dots,t^{a_{6}}\bigr),\qquad\sum_{i=0}^{6}a_{i}=0.$
(2) 1-PS limits and centralizer reduction. For each $k\in\{1,\dots,21\}$ let $\lambda_{k}$ correspond to $r_{k}$ in Proposition 2.3, and let $f_{k}$ be the generic member from Theorem 3.2. We set the 1-PS limit
$\phi_{k}:=\lim_{t\to 0}\lambda_{k}(t)\cdot f_{k}.$
Put $H:=\lambda_{k}(\mathbb{G}_{m})$ and write $C_{G}(H)$ for the centralizer in $G=SL(7)$. Let $W^{H}\subset W$ be the $H$-fixed subspace. By Luna’s centralizer reduction [Lun75], closedness of $SL(7)\cdot\phi_{k}$ in $\mathbb{P}(W)^{ss}$ is equivalent to closedness of $C_{G}(H)\cdot\phi_{k}$ in $W^{H}$.
(3) Two criteria for polystability. If $C_{G}(H)$ is a torus, we certify closedness by the convex-hull criterion (Theorem 4.1). If $C_{G}(H)$ is non-toric, we use the Casimiro–Florentino criterion (Theorem 4.4) as follows: for $\lambda\in Y\!\left(C_{G}(H)\right)$ we write $\mathrm{wt}_{\lambda}(x_{i})$ for the $\lambda$-weight on $x_{i}$, and $w(m)$ for the induced weight of a monomial $m$. We denote by $S$ the linear constraint coming from $\det=1$ on $C_{G}(H)$ (the “trace” of block weights), so $S=0$ for every $\lambda$. We then exhibit a positive linear identity
$\sum_{j}c_{j}\,w(m_{j})\;=\;C\cdot S\qquad(c_{j}>0,\;C>0).$
If $\lambda\in\Lambda_{\phi_{k}}$, then every $w(m_{j})\geq 0$ and $S=0$; hence, all $w(m_{j})=0$; solving yields a symmetric $1$-PS $\mu_{k}$, so $\Lambda_{\phi_{k}}$ is symmetric and $\phi_{k}$ is polystable.
(4) Normal forms and coefficient normalizations. Passing to “normal form,” we are allowed to: (i) multiply by a nonzero scalar; (ii) act by $C_{G}(H)/H$ (e.g., block $GL(2)$, $GL(3)$ actions) to diagonalize blocks; and (iii) use diagonal elements of $T$ (with $\prod\mu_{i}=1$) to normalize nonzero coefficients to $1$. Parameters $(\alpha,\rho,\sigma,\dots)$ record the residual moduli.
(5) Dimension count. Component dimensions are computed as
$\dim(W^{H})-\dim_{\mathrm{eff}}\!\bigl(C_{G}(H)\bigr)-1,$
where $\dim_{\mathrm{eff}}$ is the dimension of the effective $C_{G}(H)$-action on $W^{H}$ (central tori acting trivially are subtracted). We state explicitly when a central factor acts trivially.
(6) Weights and symbols. We freely reuse symbols $a_{i}$ for nonzero coefficients of $\phi_{k}$ prior to normalization. For 1-PS families obtained in the CF-check, we write $\mu_{k}(t)$. All such conventions are in force throughout §4.
Let us now determine, for each $k=1,2,\dots,21$, a polynomial whose $\mathrm{SL}(7)$-orbit is closed. The procedure is uniform across all cases. First, we take a $1$-PS limit to produce a candidate $\phi_{k}$ for a polystable point. Next, to apply Luna’s criterion, we take the stabilizer $H\subset G=\mathrm{SL}(7)$ of $\phi_{k}$; we choose $H$ as large as possible so that its centralizer $C_{G}(H)$ is as small as possible, which simplifies the closedness check. Write $W^{H}$ for the $H$-fixed locus in the ambient representation $W=\mathrm{Sym}^{3}\mathbb{C}^{7}$. If $C_{G}(H)$ is a torus, we apply the convex-hull criterion (Theorem4.1) to show that $C_{G}(H)\cdot\phi_{k}$ is closed in $W^{H}$; if $C_{G}(H)$ is not a torus, we instead apply the Casimiro–Florentino criterion (Theorem 4.4) to obtain the same conclusion. In either case, $C_{G}(H)\cdot\phi_{k}$ is closed in $W^{H}$; hence, by Luna’s criterion, the orbit $\mathrm{SL}(7)\cdot\phi_{k}$ is closed in $\mathbb{P}(\mathrm{Sym}^{3}\mathbb{C}^{7})^{ss}$. Lastly, by determining a normal form of $\phi_{k}$ under the action of $C_{G}(H)/H$, we fix the dimension of the corresponding component of the moduli space for that $k$.
4.1 Case $k=1$
1-PS limit.
Set
$\lambda_{1}(t)=\mathrm{diag}\!\bigl(t^{8},\,t^{3},\,t^{2},\,t^{-1},\,t^{-2},\,t^{-4},\,t^{-6}\bigr),\quad t\in\mathbb{G}_{m}.$
For a generic $f_{1}$ as in Section 3, the $1$-PS limit is
$\phi_{1}:=\lim_{t\to 0}\lambda_{1}(t)\cdot f_{1}=a_{1}\,x_{2}x_{3}^{2}+a_{2}\,x_{1}x_{3}x_{4}+a_{3}\,x_{2}^{2}x_{5}+a_{4}\,x_{0}x_{5}^{2}+a_{5}\,x_{1}^{2}x_{6}+a_{6}\,x_{0}x_{4}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{1}(\mathbb{G}_{m})$. The diagonal weights on $\langle x_{0},\dots,x_{6}\rangle$ are pairwise distinct; hence, $C_{G}(H)=T$ (the maximal diagonal torus). Each monomial of $\phi_{1}$ has $H$-weight 0, so $\phi_{1}\in W^{H}$.
Polystability (Luna + convex-hull).
By Luna’s criterion (see [Lun75]), closedness of the $SL(7)$-orbit of $\phi_{1}$ is equivalent to closedness of the $T$-orbit in $W^{H}$. By the convex-hull criterion (Theorem 4.1), it suffices to check that 0 is an interior point of $\mathrm{Conv}(\mathrm{Supp}(\phi_{1}))\subset X(T)_{\mathbb{R}}\simeq\mathbb{R}^{7}/\mathbb{R}(1,\dots,1)$. This holds because the exponent vectors of the six monomials satisfy
$2(0,0,1,2,0,0,0)+2(0,1,0,1,1,0,0)+2(0,0,2,0,0,1,0)$ (1)
$\qquad+2(1,0,0,0,0,2,0)+2(0,2,0,0,0,0,1)+4(1,0,0,0,1,0,1)=6(1,1,1,1,1,1,1),$
Hence, $\phi_{1}$ is polystable and $SL(7)\cdot\phi_{1}$ is closed.
Normal form and component dimension.
Let $\mathrm{diag}(\mu_{0},\dots,\mu_{6})\in T$ with $\prod_{i=0}^{6}\mu_{i}=1$. Along with the overall projective scaling, this acts on the six coefficients of $\phi_{1}$ via the characters determined by the exponent vectors in (1); we may normalize all six coefficients to $1$ simultaneously. Thus, a normal form is
$\phi^{\mathrm{nf}}_{1}=x_{2}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{2}^{2}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{0}x_{4}x_{6}.$
Residual $T$-stabilizer.
Recall that $H=\lambda_{1}(\mathbb{G}_{m})$ and $C_{G}(H)=T$, and that every monomial of $\varphi_{1}$ has $H$-weight 0, so $H$ acts trivially on $W^{H}$ and the effective group on $\mathbb{P}(W^{H})$ is $T/H$. Let $\tau=\mathrm{diag}(\mu_{0},\ldots,\mu_{6})\in T$ with $\prod_{i=0}^{6}\mu_{i}=1$. For the normal form
$\varphi_{1}^{\mathrm{nf}}=x_{2}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{2}^{2}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{0}x_{4}x_{6},$
the induced $T$-characters on the six monomials are
$\chi_{1}$ $=\mu_{2}\mu_{3}^{2}$ $(x_{2}x_{3}^{2}),$ $\chi_{2}$ $=\mu_{1}\mu_{3}\mu_{4}$ $(x_{1}x_{3}x_{4}),$
$\chi_{3}$ $=\mu_{2}^{2}\mu_{5}$ $(x_{2}^{2}x_{5}),$ $\chi_{4}$ $=\mu_{0}\mu_{5}^{2}$ $(x_{0}x_{5}^{2}),$
$\chi_{5}$ $=\mu_{1}^{2}\mu_{6}$ $(x_{1}^{2}x_{6}),$ $\chi_{6}$ $=\mu_{0}\mu_{4}\mu_{6}$ $(x_{0}x_{4}x_{6}).$
Since we work projectively, $\tau$ stabilizes $\varphi_{1}^{\mathrm{nf}}]\in\mathbb{P}(W^{H})$ if and only if $\tau\cdot\varphi_{1}^{\mathrm{nf}}=c\,\varphi_{1}^{\mathrm{nf}}$ for some $c\in\mathbb{C}^{\times}$, equivalently
$\chi_{1}=\chi_{2}=\chi_{3}=\chi_{4}=\chi_{5}=\chi_{6}.$
Solving these equations (together with $\prod\mu_{i}=1$) yields $\mu_{4}^{7}=\mu_{3}^{14}$, hence $\mu_{4}=\mu_{3}^{2}\xi$ with $\xi^{7}=1$. Writing $t:=\mu_{3}^{-1}$, we can factor
$\tau=\lambda_{1}(t)\cdot\delta_{\xi},\qquad\delta_{\xi}:=\mathrm{diag}(\xi^{-9},\xi^{-4},\xi^{-3},1,\xi,\xi^{3},\xi^{5}).$
Therefore
$\mathrm{Stab}_{T}\!\bigl([\varphi_{1}^{\mathrm{nf}}]\bigr)=H\cdot\{\delta_{\xi}:\ \xi^{7}=1\}\ \cong\ H\times\mu_{7},$
so the induced stabilizer in the effective group $T/H$ is the finite group $\mu_{7}$. In particular, after normalizing all coefficients to $1$, no continuous torus symmetry remains; hence the corresponding boundary component is zero-dimensional.
4.2 Case $k=2$
1-PS limit.
Set
$\lambda_{2}(t)=\operatorname{diag}\!\bigl(t^{6},\,t^{4},\,t,\,t^{-1},\,t^{-2},\,t^{-3},\,t^{-5}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{2}$ as in Section 3, the $1$-PS limit is
$\phi_{2}:=\lim_{t\to 0}\lambda_{2}(t)\!\cdot\!f_{2}=a_{1}x_{2}^{2}x_{4}+a_{2}x_{1}x_{4}^{2}+a_{3}x_{1}x_{3}x_{5}+a_{4}x_{0}x_{5}^{2}+a_{5}x_{1}x_{2}x_{6}+a_{6}x_{0}x_{3}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{2}(\mathbb{G}_{m})$. The diagonal weights on $\langle x_{0},\dots,x_{6}\rangle$ are pairwise distinct; hence, $C_{G}(H)=T$ (the maximal diagonal torus). Each monomial of $\phi_{2}$ has $H$-weight 0, so $\phi_{2}\in W^{H}$.
Polystability (Luna + convex-hull).
By Luna’s criterion, closedness of the $SL(7)$-orbit of $\phi_{2}$ is equivalent to closedness of the $T$-orbit in the $H$-fixed subspace. By the convex-hull criterion (Theorem 4.1), it suffices to check that 0 lies in the interior of $\operatorname{Conv}(\operatorname{Supp}(\phi_{2}))\subset X(T)_{\mathbb{R}}\simeq\mathbb{R}^{7}/\mathbb{R}(1,\ldots,1)$. This holds because
$2(0,0,2,0,1,0,0)+2(0,1,0,0,2,0,0)+2(0,1,0,1,0,1,0)$ (2)
$\quad+2(1,0,0,0,0,2,0)+2(0,1,1,0,0,0,1)+4(1,0,0,1,0,0,1)$
$=6(1,1,1,1,1,1,1),$
written in terms of the exponent vectors of the six monomials of $\phi_{2}$. Hence, $\phi_{2}$ is polystable and the orbit $SL(7)\cdot\phi_{2}$ is closed.
Normal form and component dimension.
A diagonal scaling $\operatorname{diag}(\mu_{0},\dots,\mu_{6})\in T$ with $\prod\mu_{i}=1$, together with an overall scalar, acts on the six coefficients via the characters determined by the exponent vectors in (2). Thus we may normalize all six coefficients to $1$ simultaneously. Therefore, a normal form is
$\phi^{\mathrm{nf}}_{2}=x_{2}^{2}x_{4}+x_{1}x_{4}^{2}+x_{1}x_{3}x_{5}+x_{0}x_{5}^{2}+x_{1}x_{2}x_{6}+x_{0}x_{3}x_{6}.$
The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{2}$ of the moduli is zero-dimensional.
4.3 Case $k=3$
1-PS limit.
Set
$\lambda_{3}(t)=\operatorname{diag}\!\bigl(t^{4},\,t^{2},\,t,\,t^{-1},\,t^{-1},\,t^{-2},\,t^{-3}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{3}$ as in Section 3, the $1$-PS limit is
$\phi_{3}:=\lim_{t\to 0}\lambda_{3}(t)\cdot f_{3}={}$ $a_{1}\,x_{1}x_{3}^{2}+a_{2}\,x_{1}x_{3}x_{4}+a_{3}\,x_{1}x_{4}^{2}$
$\quad+a_{4}\,x_{2}^{2}x_{5}+a_{5}\,x_{0}x_{5}^{2}+a_{6}\,x_{1}x_{2}x_{6}+a_{7}\,x_{0}x_{3}x_{6}+a_{8}\,x_{0}x_{4}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{3}(\mathbb{G}_{m})$. The multiplicities of the diagonal weights on $\langle x_{0},\dots,x_{6}\rangle$ are $1$ on $x_{0},x_{1},x_{2},x_{5},x_{6}$ and $2$ on $\langle x_{3},x_{4}\rangle$; hence,
$C_{G}(H)=\left\{\begin{aligned} &\operatorname{diag}(\alpha_{0},\alpha_{1},\alpha_{2})\oplus A\oplus\operatorname{diag}(\alpha_{5},\alpha_{6}):\\[-1.0pt] &\quad\alpha_{i}\in\mathbb{G}_{m},\ A\in\mathrm{GL}(2),\ \alpha_{0}\alpha_{1}\alpha_{2}\,\det(A)\,\alpha_{5}\alpha_{6}=1\end{aligned}\right\}\cong\mathrm{SL}(2)\times\mathbb{G}_{m}^{5}.$
Each monomial of $\phi_{3}$ has $H$-weight 0, so $\phi_{3}$ is $H$-fixed.
Polystability (Luna + Casimiro–Florentino).
By Luna’s slice/centralizer reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{3}$ is equivalent to polystability for the $C_{G}(H)$-action on the $H$-fixed subspace. After conjugating inside the $\mathrm{SL}(2)$-block, any $\lambda\in Y(C_{G}(H))$ may be chosen with weights
$\mathrm{wt}(x_{0},\dots,x_{6})=(a_{0},\,a_{1},\,a_{2},\,c+n,\,c-n,\,a_{5},\,a_{6}),\qquad S:=a_{0}+a_{1}+a_{2}+2c+a_{5}+a_{6}=0$
(as in Convention 4.6. Let $w_{i}$ be the $\lambda$-weight of the $i$-th monomial of $\phi_{3}$. Then
$w_{1}$ $=a_{1}+2(c+n),$ $w_{2}$ $=a_{1}+2c,$ $w_{3}$ $=a_{1}+2(c-n),$
$w_{4}$ $=2a_{2}+a_{5},$ $w_{5}$ $=a_{0}+2a_{5},$ $w_{6}$ $=a_{1}+a_{2}+a_{6},$
$w_{7}$ $=a_{0}+(c+n)+a_{6},$ $w_{8}$ $=a_{0}+(c-n)+a_{6}.$
A direct computation yields the positive linear identity
$w_{1}+w_{2}+2w_{3}+2w_{4}+2w_{5}+2w_{6}+3w_{7}+w_{8}\;=\;6S.$ (3)
If $\lambda\in\Lambda_{\phi_{3}}$, then all $w_{i}\geq 0$ and $S=0$; by (3) they must all vanish. Solving gives
$n=0,\quad a_{1}=-2c,\quad a_{5}=-2a_{2},\quad a_{0}=4a_{2},\quad a_{6}=2c-a_{2},\quad a_{2}=-c.$
Thus, with $k\in\mathbb{Z}$,
$\mu_{k}(t):=\operatorname{diag}\!\bigl(t^{4k},\,t^{2k},\,t^{k},\,t^{-k},\,t^{-k},\,t^{-2k},\,t^{-3k}\bigr),\qquad\Lambda_{\phi_{3}}=\{\mu_{k}\mid k\in\mathbb{Z}\}.$
Therefore $\Lambda_{\phi_{3}}$ is symmetric, and by the Casimiro–Florentino criterion $\phi_{3}$ is polystable; in particular, $\mathrm{SL}(7)\cdot\phi_{3}$ is closed.
Normal form and component dimension.
On the $H$-fixed slice, the eight weight-zero monomials are
$x_{1}x_{3}^{2},\ x_{1}x_{3}x_{4},\ x_{1}x_{4}^{2},\ x_{2}^{2}x_{5},\ x_{0}x_{5}^{2},\ x_{1}x_{2}x_{6},\ x_{0}x_{3}x_{6},\ x_{0}x_{4}x_{6},$
Therefore, $W^{H}$ denotes their span. i.e.
$W^{H}=\underbrace{\langle x_{1}\rangle\otimes\mathrm{Sym}^{2}\langle x_{3},x_{4}\rangle}_{\mathrm{(I)}}\oplus\underbrace{\mathrm{Sym}^{2}\langle x_{2}\rangle\otimes x_{5}}_{\mathrm{(II)}}\oplus\underbrace{x_{0}\otimes\mathrm{Sym}^{2}\langle x_{5}\rangle}_{\mathrm{(III)}}\oplus\underbrace{x_{0}\otimes\langle x_{3},x_{4}\rangle\otimes x_{6}}_{\mathrm{(IV)}}\oplus\underbrace{x_{1}\otimes\langle x_{2}\rangle\otimes x_{6}}_{\mathrm{(V)}}.$
The centralizer is
$C_{G}(H)\cong\mathrm{SL}(2)\times\mathbb{G}_{m}^{5},$
acting by $\mathrm{SL}(2)$ on $\langle x_{3},x_{4}\rangle$ and by a diagonal torus on $\langle x_{0},x_{1},x_{2},x_{5},x_{6}\rangle$ (subject to the product-one condition).
(I) binary quadratic on $\langle x_{3},x_{4}\rangle$.
Write the $x_{1}$-part as a binary quadratic $Q=a_{1}x_{3}^{2}+a_{2}x_{3}x_{4}+a_{3}x_{4}^{2}\in\mathrm{Sym}^{2}\langle x_{3},x_{4}\rangle$. As $\Delta(Q)$ is $\mathrm{SL}(2)$-invariant while rescaling $x_{1}$ scales $Q$ (and hence $\Delta$) homogenously, for a generic (nondegenerate) $Q$, there exists $A\in\mathrm{SL}(2)$ and a rescaling of $x_{1}$ such that $Q\sim x_{3}^{2}+x_{3}x_{4}+x_{4}^{2}$. Thus, the block $\mathrm{(I)}$ is fixed to $x_{1}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{1}x_{4}^{2}$.
(II)(III)(IV)(V) torus normalizations.
Let $T^{\prime}=\{\mathrm{diag}(\mu_{0},\mu_{1},\mu_{2})\oplus I_{2}\oplus\mathrm{diag}(\mu_{5},\mu_{6}):\ \mu_{0}\mu_{1}\mu_{2}\mu_{5}\mu_{6}=1\}$. Under $T^{\prime}$, the coefficients transform by the characters determined by exponent vectors: $x_{2}^{2}x_{5}$ by $\mu_{2}^{2}\mu_{5}$, $x_{0}x_{5}^{2}$ by $\mu_{0}\mu_{5}^{2}$, $x_{1}x_{2}x_{6}$ by $\mu_{1}\mu_{2}\mu_{6}$, while $x_{0}x_{3}x_{6}$ and $x_{0}x_{4}x_{6}$ both by $\mu_{0}\mu_{6}$. Using $\mu_{0},\mu_{1},\mu_{2},\mu_{5},\mu_{6}$ (together with an overall scalar), we set $x_{2}^{2}x_{5}$, $x_{0}x_{5}^{2}$, and $x_{1}x_{2}x_{6}$ to have coefficient $1$. More precisely, we are imposing the conditions
$\mu_{2}^{2}\mu_{5}=1,\qquad\mu_{0}\mu_{5}^{2}=1,\qquad\mu_{1}\mu_{2}\mu_{6}=1,$
together with the product-one constraint
$\mu_{0}\mu_{1}\mu_{2}\mu_{5}\mu_{6}=1.$
We can solve these equations explicitly: choose $\mu_{2},\mu_{6}\in\mathbb{C}^{\times}$ freely, set
$\mu_{5}=\mu_{2}^{-2},\qquad\mu_{0}=\mu_{5}^{-2}=\mu_{2}^{4},\qquad\mu_{1}=(\mu_{2}\mu_{6})^{-1}.$
Then the product-one constraint $\mu_{0}\mu_{1}\mu_{2}\mu_{5}\mu_{6}=1$ is automatically satisfied. Consequently, we can normalize the coefficients of $x_{2}^{2}x_{5}$, $x_{0}x_{5}^{2}$, and $x_{1}x_{2}x_{6}$ to be all equal to $1$.
(IV)The remaining pair and the residual $\mathrm{SL}(2)$-symmetry. The pair $(x_{0}x_{3}x_{6},\;x_{0}x_{4}x_{6})$ is scaled by the same torus character $\mu_{0}\mu_{6}$, so the diagonal torus $T^{\prime}$ alone does not change their ratio. However, after fixing the nondegenerate binary quadratic
$Q\sim x_{3}^{2}+x_{3}x_{4}+x_{4}^{2},$
the stabilizer $\mathrm{Stab}_{\mathrm{SL}(2)}(Q)$ is not finite: it is a one-dimensional torus (isomorphic to $\mathrm{SO}(2,\mathbb{C})\cong\mathbb{G}_{m}$). Its induced action on $\mathbb{P}\langle x_{3},x_{4}\rangle$ has two fixed points (the two roots of $Q$) and a single open orbit. Hence, for a generic element (i.e. when the linear form in block (IV) is not proportional to a root of $Q$), we may act by $\mathrm{Stab}_{\mathrm{SL}(2)}(Q)$ to normalize that linear form to $x_{3}+x_{4}$. Using the remaining diagonal torus (and projective rescaling), we normalize its coefficient to $1$.
Consequently, no continuous parameter remains, and a convenient normal form is
$\varphi^{\mathrm{nf}}_{3}=x_{1}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{0}x_{5}^{2}+x_{1}x_{2}x_{6}+x_{0}x_{3}x_{6}+x_{0}x_{4}x_{6}.$
Finally, $\dim W^{H}=8$ and $\dim C_{G}(H)=8$, while $H\simeq\mathbb{G}_{m}$ acts trivially on $W^{H}$. Thus the effective group dimension is 81=78-1=7, and after projectivizing we obtain
$\dim(\Phi_{3})=8-7-1=0.$
4.4 Case $k=4$
1-PS limit.
$\lambda_{4}(t)=\mathrm{diag}\!\left(t^{3},\,t^{2},\,t,\,1,\,t^{-1},\,t^{-2},\,t^{-3}\right),\quad t\in\mathbb{G}_{m}.$
For a generic $f_{4}$ as in Section 3, the $1$-PS limit is
$\phi_{4}$ $=\lim_{t\to 0}\lambda_{4}(t)\cdot f_{4}$
$=a_{1}\,x_{3}^{3}+a_{2}\,x_{2}x_{3}x_{4}+a_{3}\,x_{1}x_{4}^{2}+a_{4}\,x_{2}^{2}x_{5}+a_{5}\,x_{1}x_{3}x_{5}+a_{6}\,x_{0}x_{4}x_{5}+a_{7}\,x_{1}x_{2}x_{6}+a_{8}\,x_{0}x_{3}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{4}(\mathbb{G}_{m})$. The diagonal weights on $\langle x_{0},\dots,x_{6}\rangle$ are pairwise distinct; hence $C_{G}(H)=T$ (the maximal diagonal torus). Each monomial of $\phi_{4}$ has $H$-weight 0, so $\phi_{4}\in W^{H}$.
Polystability (Luna + convex-hull).
By Luna’s criterion, closedness of the $SL(7)$-orbit of $\phi_{4}$ is equivalent to closedness of the $T$-orbit in $W^{H}$. By the convex-hull criterion (Theorem 4.1), it suffices to check that 0 is an interior point of
$\mathrm{Conv}(\mathrm{Supp}(\phi_{4}))\subset X(T)_{\mathbb{R}}\cong\mathbb{R}^{7}/\mathbb{R}(1,\dots,1),$
which holds because the exponent vectors satisfy
$(0,0,0,3,0,0,0)+(0,0,1,1,1,0,0)+(0,1,0,0,2,0,0)+(0,0,2,0,0,1,0)$ (4)
$\quad+2(0,1,0,1,0,1,0)+6(1,0,0,0,1,1,0)+6(0,1,1,0,0,0,1)$
$\quad+3(1,0,0,1,0,0,1)=9(1,1,1,1,1,1,1).$
Hence, $\phi_{4}$ is polystable and $SL(7)\cdot\phi_{4}$ is closed.
Normal form and component dimension.
Since every monomial of $\phi_{4}$ has $H$-weight 0, the subgroup $H\subset T$ acts trivially on $W^{H}$. Hence the $T$-action on $\mathbb{P}(W^{H})\cong\mathbb{P}^{7}$ factors through the effective torus $T/H$, which has dimension $5$.
For a point of $\mathbb{P}(W^{H})$ with all eight coefficients nonzero, the stabilizer in $T/H$ is finite (equivalently, the eight exponent vectors have affine rank $5$), so a general $(T/H)$-orbit has dimension $5$. Therefore the corresponding boundary component has dimension
$\dim\Phi_{4}=\dim\mathbb{P}(W^{H})-\dim(T/H)=7-5=2.$
Using the $5$ parameters of $T/H$ together with the overall projective scaling, we can normalize six of the eight coefficients to $1$, leaving two essential parameters. A convenient normal form is
$\phi^{\mathrm{nf}}_{4}(\alpha,\beta)=x_{3}^{3}+x_{2}x_{3}x_{4}+x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{4}x_{5}+\alpha\,x_{1}x_{2}x_{6}+\beta\,x_{0}x_{3}x_{6},\qquad(\alpha,\beta)\in(\mathbb{C}^{\times})^{2}.$
In particular, the corresponding component $\Phi_{4}$ of the moduli is two-dimensional.
4.5 Case $k=5$
1-PS limit.
Set
$\lambda_{5}(t)=\operatorname{diag}\!\bigl(t^{4},\,t^{2},\,t,\,1,\,t^{-1},\,t^{-2},\,t^{-4}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{5}$ as in Section 3, the $1$-PS limit is
$\phi_{5}:=\lim_{t\to 0}\lambda_{5}(t)\cdot f_{5}=a_{1}x_{3}^{3}+a_{2}x_{2}x_{3}x_{4}+a_{3}x_{1}x_{4}^{2}+a_{4}x_{2}^{2}x_{5}+a_{5}x_{1}x_{3}x_{5}+a_{6}x_{0}x_{5}^{2}+a_{7}x_{1}^{2}x_{6}+a_{8}x_{0}x_{3}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{5}(\mathbb{G}_{m})$. The diagonal weights on $\langle x_{0},\ldots,x_{6}\rangle$ are pairwise distinct; hence, $C_{G}(H)=T$ (the maximal diagonal torus). Each monomial of $\phi_{5}$ has $H$-weight 0, so $\phi_{5}\in W^{H}$.
Polystability (Luna + convex-hull).
By Luna’s criterion, closedness of the $SL(7)$-orbit of $\phi_{5}$ is equivalent to closedness of the $T$-orbit in the $H$-fixed subspace. By the convex-hull criterion (Theorem 4.1), it suffices to check that 0 lies in the interior of $\operatorname{Conv}(\operatorname{Supp}(\phi_{5}))\subset X(T)_{\mathbb{R}}\cong\mathbb{R}^{7}/\mathbb{R}(1,\ldots,1)$. This holds because the exponent vectors of the eight monomials satisfy
$(0,0,0,3,0,0,0)+(0,0,1,1,1,0,0)$ (5)
$10(0,0,2,0,0,1,0)\displaystyle\quad+$0(0,1,0,0,2,0,0)+0(0,0,2,0,0,1,0)
$\quad+(0,1,0,1,0,1,0)+5(1,0,0,0,0,2,0)$
[…]
Hence, $\phi_{5}$ is polystable and $SL(7)\cdot\phi_{5}$ is closed.
Normal form and component dimension.
Since every monomial of $\phi_{5}$ has $H$-weight 0, the subgroup $H\subset T$ acts trivially on $W^{H}$. Hence the $T$-action on $\mathbb{P}(W^{H})\cong\mathbb{P}^{7}$ factors through the effective torus $T/H$, which has dimension $5$.
For a point of $\mathbb{P}(W^{H})$ with all eight coefficients nonzero, the stabilizer in $T/H$ is finite (equivalently, the eight exponent vectors have affine rank $5$), so a general $(T/H)$-orbit has dimension $5$. Therefore the corresponding boundary component has dimension
$\dim\Phi_{5}=\dim\mathbb{P}(W^{H})-\dim(T/H)=7-5=2.$
Using the $5$ parameters of $T/H$ together with the overall projective scaling, we can normalize six of the eight coefficients to $1$, leaving two essential parameters. A convenient normal form is
$\phi^{\mathrm{nf}}_{5}(\alpha,\beta)=x_{3}^{3}+x_{2}x_{3}x_{4}+x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{5}^{2}+\alpha\,x_{1}^{2}x_{6}+\beta\,x_{0}x_{3}x_{6},\qquad(\alpha,\beta)\in(\mathbb{C}^{\times})^{2}.$
In particular, the corresponding component $\Phi_{5}$ of the moduli is two-dimensional.
4.6 Case $k=6$
1-PS limit.
Set
$\lambda_{6}(t)=\mathrm{diag}\!\left(t^{5},\,t^{3},\,t^{2},\,t,\,t^{-1},\,t^{-4},\,t^{-6}\right),\quad t\in\mathbb{G}_{m}.$
For a generic $f_{6}$ as in Section 3, the $1$-PS limit is
$\phi_{6}:=\lim_{t\to 0}\lambda_{6}(t)\cdot f_{6}=a_{1}\,x_{2}x_{4}^{2}+a_{2}\,x_{2}^{2}x_{5}+a_{3}\,x_{1}x_{3}x_{5}+a_{4}\,x_{0}x_{4}x_{5}+a_{5}\,x_{1}^{2}x_{6}+a_{6}\,x_{0}x_{3}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{6}(\mathbb{G}_{m})$. The diagonal weights on $\langle x_{0},\dots,x_{6}\rangle$ are pairwise distinct; hence $C_{G}(H)=T$ (the maximal diagonal torus). Each monomial of $\phi_{6}$ has $H$-weight 0, so $\phi_{6}\in W^{H}$.
Polystability (Luna + convex-hull).
Based on Luna’s criterion, closedness of the $SL(7)$-orbit of $\phi_{6}$ is equivalent to closedness of the $T$-orbit in the $H$-fixed subspace. By Theorem 4.1, it suffices to check that 0 is an interior point of
$\mathrm{Conv}(\mathrm{Supp}(\phi_{6}))\subset X(T)_{\mathbb{R}}\cong\mathbb{R}^{7}/\mathbb{R}(1,\dots,1),$
which holds because the exponent vectors of the six monomials of $\phi_{6}$ satisfy the positive relation
$(0,0,1,0,2,0,0)+(0,0,2,0,0,1,0)+(0,1,0,1,0,1,0)$ (6)
$\qquad+(1,0,0,0,1,1,0)+(0,2,0,0,0,0,1)+2(1,0,0,1,0,0,1)$
$=3(1,1,1,1,1,1,1).$
Hence, $\phi_{6}$ is polystable and $SL(7)\cdot\phi_{6}$ is closed.
Normal form and component dimension.
A diagonal scaling $\mathrm{diag}(\mu_{0},\cdots,\mu_{6})\in T$ with $\Pi\mu_{i}=1$, together with an overall scalar, acts on the six coefficients via the characters determined by the exponent vectors in (6). Hence we may normalize all six coefficients to $1$ simultaneously. A normal form is therefore
$\phi^{\mathrm{nf}}_{6}=x_{2}x_{4}^{2}+x_{2}^{2}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{4}x_{5}+x_{1}^{2}x_{6}+x_{0}x_{3}x_{6}.$
The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{6}$ of the moduli is zero-dimensional.
4.7 Case $k=7$
1-PS limit.
Set
$\lambda_{7}(t)=\mathrm{diag}\!\left(t^{6},\,t^{4},\,t^{2},\,t,\,t^{-2},\,t^{-3},\,t^{-8}\right),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{7}$ as in Section 3, the $1$-PS limit is
$\phi_{7}:=\lim_{t\to 0}\lambda_{7}(t)\cdot f_{7}=a_{1}\,x_{3}^{2}x_{4}+a_{2}\,x_{1}x_{4}^{2}+a_{3}\,x_{2}x_{3}x_{5}+a_{4}\,x_{0}x_{5}^{2}+a_{5}\,x_{1}^{2}x_{6}+a_{6}\,x_{0}x_{2}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{7}(\mathbb{G}_{m})$. The diagonal weights on $\langle x_{0},\ldots,x_{6}\rangle$ are pairwise distinct; hence $C_{G}(H)=T$ (the maximal diagonal torus). Each monomial of $\phi_{7}$ has $H$-weight 0, so $\phi_{7}$ lies in the $H$-fixed subspace.
Polystability (Luna + convex-hull).
By Luna’s criterion, closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{7}$ is equivalent to closedness of the $T$-orbit in the $H$-fixed subspace. By the convex-hull criterion (Theorem 4.1), it suffices to check that 0 is an interior point of
$\mathrm{Conv}(\mathrm{Supp}(\phi_{7}))\subset X(T)_{\mathbb{R}}\cong\mathbb{R}^{7}/\mathbb{R}(1,\ldots,1),$
which holds because
$(0,0,0,2,1,0,0)+(0,1,0,0,2,0,0)+(0,0,1,1,0,1,0)$ (7)
$\quad+(1,0,0,0,0,2,0)+(0,2,0,0,0,0,1)+2(1,0,1,0,0,0,1)$
$=3(1,1,1,1,1,1,1),$
written in terms of the exponent vectors of the six monomials of $\phi_{7}$. Hence, $\phi_{7}$ is polystable and $\mathrm{SL}(7)\cdot\phi_{7}$ is closed.
Normal form and component dimension.
A diagonal scaling $\mathrm{diag}(\mu_{0},\ldots,\mu_{6})\in T$ with $\prod\mu_{i}=1$, together with an overall scalar, acts on the six coefficients via the characters determined by the exponent vectors in (7). Thus we may normalize all six coefficients to $1$ simultaneously. Therefore, a normal form is
$\phi^{\mathrm{nf}}_{7}=x_{3}^{2}x_{4}+x_{1}x_{4}^{2}+x_{2}x_{3}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{0}x_{2}x_{6}.$
The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{7}$ of the moduli is zero-dimensional.
4.8 Case $k=8$
1-PS limit.
Set
$\lambda_{8}(t)=\mathrm{diag}\!\left(t^{4},\,t,\,t,\,1,\,t^{-2},\,t^{-2},\,t^{-2}\right),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{8}$ as in Section 3, the $1$-PS limit is
$\phi_{8}:=\lim_{t\to 0}\lambda_{8}(t)\cdot f_{8}$ $=a_{1}x_{3}^{3}+a_{2}x_{1}^{2}x_{4}+a_{3}x_{1}x_{2}x_{4}+a_{4}x_{2}^{2}x_{4}+a_{6}x_{1}^{2}x_{5}+a_{7}x_{1}x_{2}x_{5}+a_{8}x_{2}^{2}x_{5}$
$\quad+a_{11}x_{1}^{2}x_{6}+a_{12}x_{1}x_{2}x_{6}+a_{13}x_{2}^{2}x_{6}+a_{5}x_{0}x_{4}^{2}+a_{9}x_{0}x_{4}x_{5}+a_{14}x_{0}x_{4}x_{6}$
$\quad+a_{10}x_{0}x_{5}^{2}+a_{15}x_{0}x_{5}x_{6}+a_{16}x_{0}x_{6}^{2}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{8}(\mathbb{G}_{m})$. The weights on $\langle x_{0},\dots,x_{6}\rangle$ are (4,1,1,0,2,2,2)(4,1,1,0,-2,-2,-2) with multiplicities $(1,2,1,3)$; hence,
$C_{G}(H)=\Bigl\{\mathrm{diag}(\alpha)\oplus A\oplus\mathrm{diag}(\beta)\oplus B:\ \alpha,\beta\in\mathbb{G}_{m},\ A\in\mathrm{GL}(2),\ B\in\mathrm{GL}(3),\ \alpha\beta\det(A)\det(B)=1\Bigr\}.$
Thus $C_{G}(H)\cong\bigl(\mathbb{G}_{m}\times\mathrm{GL}(2)\times\mathbb{G}_{m}\times\mathrm{GL}(3)\bigr)\cap\mathrm{SL}(7)$ and $\dim C_{G}(H)=14$. Every monomial of $\phi_{8}$ has $H$-weight 0; hence, $\phi_{8}\in W^{H}$.
Polystability (Luna + Casimiro–Florentino).
Based on Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{8}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating inside the $\mathrm{GL}(2)$- and $\mathrm{GL}(3)$-blocks, any $\lambda\in Y(C_{G}(H))$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha,\,s+u,\,s-u,\,\beta,\,\gamma+v_{1},\,\gamma+v_{2},\,\gamma+v_{3}),$
where $\alpha,\beta,s,u,\gamma,v_{i}\in\mathbb{Z}$, $v_{1}+v_{2}+v_{3}=0$, and
$S:=\alpha+\beta+2s+3\gamma=0$
is the $\mathrm{SL}$-constraint fixed in Convention 4.6. Let $w(\cdot)$ be the weight of a monomial. A direct computation yields the positive linear identity
$\sum_{j=4}^{6}\Bigl(w(x_{1}^{2}x_{j})+w(x_{2}^{2}x_{j})+2\,w(x_{1}x_{2}x_{j})\Bigr)+3w(x_{0}x_{4}^{2})+3w(x_{0}x_{5}^{2})+3w(x_{0}x_{6}^{2})$
$+w(x_{0}x_{4}x_{5})+w(x_{0}x_{4}x_{6})+w(x_{0}x_{5}x_{6})+4\,w(x_{3}^{3})=12S.$ (8)
If $\lambda\in\Lambda_{\phi_{8}}$, then all the above weights are $\geq 0$ and $S=0$; therefore, by (8) they must all vanish. Solving gives
$\beta=0,\quad v_{1}=v_{2}=v_{3}=0,\quad u=0,\quad 2s+\gamma=0,\quad\alpha+2\gamma=0.$
Putting $s=k\in\mathbb{Z}$, we obtain
$\mu_{k}(t)=\mathrm{diag}\!\left(t^{4k},\,t^{k},\,t^{k},\,1,\,t^{-2k},\,t^{-2k},\,t^{-2k}\right),\qquad\Lambda_{\phi_{8}}=\{\mu_{k}\mid k\in\mathbb{Z}\},$
which is symmetric. Hence, by the Casimiro–Florentino criterion, $\phi_{8}$ is polystable.
Normal form and component dimension.
On the $H$-fixed subspace
$W^{H}=\underbrace{\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle\otimes\langle x_{4},x_{5},x_{6}\rangle}_{\mathrm{(I)}}\ \oplus\ \underbrace{x_{0}\otimes\mathrm{Sym}^{2}\langle x_{4},x_{5},x_{6}\rangle}_{\mathrm{(II)}}\ \oplus\underbrace{\langle x_{3}^{3}\rangle}_{\mathrm{(III)}},$
use $\mathrm{GL}(3)$ on $\langle x_{4},x_{5},x_{6}\rangle$ to diagonalize the quadratic, then the left/right actions on $\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle\otimes\langle x_{4},x_{5},x_{6}\rangle$ to diagonalize the $3\times 3$ block (SVD-type reduction under left $\mathrm{Sym}^{2}\mathrm{GL}(2)$ and right $\mathrm{O}(U,q)$, here $U=\langle x_{4},x_{5},x_{6}\rangle$ and $q=x_{4}^{2}+x_{5}^{2}+x_{6}^{2}$) and normalize one diagonal entry to $1$; the remaining two appear as parameters $\rho,\sigma$. A normal form is
$\phi^{\mathrm{nf}}_{8}(\rho,\sigma)=x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{5}^{2}+x_{0}x_{6}^{2}+x_{1}^{2}x_{4}+\rho\,x_{1}x_{2}x_{5}+\sigma\,x_{2}^{2}x_{6},\qquad(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}.$
As $\dim W^{H}=16$ and the effective action has dimension $13$ (after projectivizing), the closed component has dimension 16131=216-13-1=2.
Remark 4.7.
Here $\mathrm{O}(U,q)$ denotes the orthogonal group of the quadratic space $(U,q)$, i.e. $\mathrm{O}(U,q)=\{\,g\in\mathrm{GL}(U)\mid q(gu)=q(u)\ \forall u\in U\}$. In the chosen basis $U=\langle x_{4},x_{5},x_{6}\rangle$ with $q=x_{4}^{2}+x_{5}^{2}+x_{6}^{2}$, this is $\{\,B\in\mathrm{GL}_{3}\mid B^{\mathsf{T}}B=I_{3}\,\}$.
The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{8}$ of the moduli is two-dimensional.
4.9 Case $k=9$
1-PS limit.
Set
$\lambda_{9}(t)=\operatorname{diag}\!\bigl(t^{2},\,t^{2},\,1,\,1,\,t^{-1},\,t^{-1},\,t^{-2}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{9}$ as in Section 3, the $1$-PS limit is
$\phi_{9}:=\lim_{t\to 0}\lambda_{9}(t)\cdot f_{9}=\begin{aligned} &a_{1}x_{2}^{3}+a_{2}x_{2}^{2}x_{3}+a_{3}x_{2}x_{3}^{2}+a_{4}x_{3}^{3}\\ &{}+a_{5}x_{0}x_{4}^{2}+a_{6}x_{1}x_{4}^{2}+a_{7}x_{0}x_{4}x_{5}+a_{8}x_{1}x_{4}x_{5}\\ &{}+a_{9}x_{0}x_{5}^{2}+a_{10}x_{1}x_{5}^{2}\\ &{}+a_{11}x_{0}x_{2}x_{6}+a_{12}x_{1}x_{2}x_{6}+a_{13}x_{0}x_{3}x_{6}+a_{14}x_{1}x_{3}x_{6}.\end{aligned}$
(All monomials have $H$-weight 0.)
$H$ and $C_{G}(H)$.
Let $H=\lambda_{9}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,2,0,0,1,1,2)(2,2,0,0,-1,-1,-2) with block decomposition $\langle x_{0},x_{1}\rangle$, $\langle x_{2},x_{3}\rangle$, $\langle x_{4},x_{5}\rangle$, $\langle x_{6}\rangle$. Hence,
$C_{G}(H)=\Bigl\{A\oplus B\oplus C\oplus\gamma:\ A,B,C\in\mathrm{GL}(2),\ \gamma\in\mathbb{G}_{m},\ \det(A)\det(B)\det(C)\gamma=1\Bigr\}$
$\cong\ \bigl(\mathrm{GL}(2)^{3}\times\mathbb{G}_{m}\bigr)\cap\mathrm{SL}(7).$
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{9}$ is equivalent to polystability for the $C_{G}(H)$-action on the $H$-fixed locus. After conjugating inside each $\mathrm{GL}(2)$-block, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(a+u,\,a-u,\,b+v,\,b-v,\,c+w,\,c-w,\,d),\qquad S:=2a+2b+2c+d=0.$
A direct computation yields the positive linear identity
$\bigl[w(x_{2}^{3})+w(x_{2}^{2}x_{3})+w(x_{2}x_{3}^{2})+w(x_{3}^{3})\bigr]$
$\quad+2\bigl[w(x_{0}x_{4}^{2})+w(x_{0}x_{4}x_{5})+w(x_{0}x_{5}^{2})+w(x_{1}x_{4}^{2})+w(x_{1}x_{4}x_{5})+w(x_{1}x_{5}^{2})\bigr]$
$\quad+3\bigl[w(x_{0}x_{2}x_{6})+w(x_{1}x_{2}x_{6})+w(x_{0}x_{3}x_{6})+w(x_{1}x_{3}x_{6})\bigr]=12\,S.$ (9)
If $\lambda\in\Lambda_{\phi_{9}}$, then all the above weights are $\geq 0$ and $S=0$; hence, by $(9)$ they all vanish. Solving gives
$b=v=u=w=0,\qquad a+2c=0,\qquad a+d=0.$
Writing $a=2k$ with $k\in\mathbb{Z}$ we obtain
$\mu_{k}(t):=\operatorname{diag}\!\bigl(t^{2k},\,t^{2k},\,1,\,1,\,t^{-k},\,t^{-k},\,t^{-2k}\bigr),\qquad\Lambda_{\phi_{9}}=\{\mu_{k}\mid k\in\mathbb{Z}\}.$
Thus $\Lambda_{\phi_{9}}$ is symmetric, and by the Casimiro–Florentino criterion $\phi_{9}$ is polystable; in particular $\mathrm{SL}(7)\cdot\phi_{9}$ is closed.
Normal form and component dimension.
From now on, we normalize coefficients under the $C_{G}(H)$-action. Decompose
$W^{H}=\underbrace{\mathrm{Sym}^{3}\langle x_{2},x_{3}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{\langle x_{0},x_{1}\rangle\otimes\mathrm{Sym}^{2}\langle x_{4},x_{5}\rangle}_{(\mathrm{II})}\ \oplus\ \underbrace{\langle x_{0},x_{1}\rangle\otimes\langle x_{2},x_{3}\rangle\otimes\langle x_{6}\rangle}_{(\mathrm{III})}.$
(I) Binary cubic $\mathrm{Sym}^{3}\langle x_{2},x_{3}\rangle$.
A general element is $\mathrm{GL}(\langle x_{2},x_{3}\rangle)$-equivalent (up to overall scaling) to
$x_{2}^{2}x_{3}+\tau\,x_{2}x_{3}^{2},\qquad\tau\in\mathbb{C}^{\times},$
leaving one parameter $\tau$.
Indeed, over $\mathbb{C}$ any binary cubic $F\in\mathrm{Sym}^{3}\langle x_{2},x_{3}\rangle$ factors as
$F=L_{1}L_{2}L_{3},\qquad L_{i}\in\langle x_{2},x_{3}\rangle,$
and for a general $F$ the three linear factors (equivalently, the three roots on $\mathbb{P}^{1}=\mathbb{P}(\langle x_{2},x_{3}\rangle)$) are distinct. Since $\mathrm{PGL}(\langle x_{2},x_{3}\rangle)$ acts 33–transitively on $\mathbb{P}^{1}$, after a change of basis we may send two of the roots to [1:0][1\!:\!0] and [0:1][0\!:\!1], i.e. arrange $L_{1}=x_{2}$ and $L_{2}=x_{3}$. Thus
$F=x_{2}x_{3}(\alpha x_{2}+\beta x_{3}),\qquad\alpha,\beta\in\mathbb{C}^{\times}.$
Up to an overall scalar we may take $\alpha=1$ and set $\tau:=\beta/\alpha$, giving
$F\sim x_{2}^{2}x_{3}+\tau\,x_{2}x_{3}^{2}.$
(II) $\langle x_{0},x_{1}\rangle\otimes\mathrm{Sym}^{2}\langle x_{4},x_{5}\rangle$ ($2\times 3$ block).
Via the right $\mathrm{GL}(\langle x_{4},x_{5}\rangle)$ (through $\mathrm{Sym}^{2}$) diagonalize a reference quadratic to $q=x_{4}^{2}+x_{5}^{2}$; the residual right group is $O(U,q)$ on $U=\langle x_{4},x_{5}\rangle$. Using the left $\mathrm{GL}(\langle x_{0},x_{1}\rangle)$ together with this right orthogonal action (an SVD-type reduction), eliminate the $x_{4}x_{5}$ cross term and equalize the $x_{4}^{2}$ entries. After central torus/projective scalings,
$x_{0}(x_{4}^{2}+\rho\,x_{5}^{2})+x_{1}(x_{4}^{2}+x_{5}^{2}),\qquad\rho\in\mathbb{C}^{\times}.$
(III) $\langle x_{0},x_{1}\rangle\otimes\langle x_{2},x_{3}\rangle\otimes\langle x_{6}\rangle$ ($2\times 2$ block).
With $\mathrm{GL}(\langle x_{0},x_{1}\rangle)$, $\mathrm{GL}(\langle x_{2},x_{3}\rangle)$ (respecting the choice in (I)), and scaling $x_{6}$, we diagonalize to
$x_{0}x_{2}x_{6}+x_{1}x_{3}x_{6},$
and normalize the coefficients to $1$.
Combining the three steps yields the normal form
$\varphi^{\mathrm{nf}}_{9}(\tau,\rho)=x_{2}^{2}x_{3}+\tau\,x_{2}x_{3}^{2}+x_{0}x_{4}^{2}+\rho\,x_{0}x_{5}^{2}+x_{1}x_{4}^{2}+x_{1}x_{5}^{2}+x_{0}x_{2}x_{6}+x_{1}x_{3}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$
As $\dim W^{H}=14$ and $\dim C_{G}(H)=12$, the effective action has dimension $11$; after projectivizing, we obtain 14111=214-11-1=2. The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{9}$ of the moduli is two-dimensional.
4.10 Case $k=10$
1-PS limit.
Set
$\lambda_{10}(t)=\operatorname{diag}\!\bigl(t^{2},\,t,\,1,\,1,\,t^{-1},\,t^{-1},\,t^{-1}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{10}$ as in Section 3, the $1$-PS limit is
$\phi_{10}=\lim_{t\to 0}\lambda_{10}(t)\cdot f_{10}$ $=a_{1}x_{2}^{3}+a_{2}x_{2}^{2}x_{3}+a_{3}x_{2}x_{3}^{2}+a_{4}x_{3}^{3}$
$\quad+a_{5}x_{1}x_{2}x_{4}+a_{6}x_{1}x_{3}x_{4}+a_{8}x_{1}x_{2}x_{5}+a_{9}x_{1}x_{3}x_{5}$
$\quad+a_{12}x_{1}x_{2}x_{6}+a_{13}x_{1}x_{3}x_{6}$
$\quad+a_{7}x_{0}x_{4}^{2}+a_{10}x_{0}x_{4}x_{5}+a_{11}x_{0}x_{5}^{2}+a_{14}x_{0}x_{4}x_{6}+a_{15}x_{0}x_{5}x_{6}+a_{16}x_{0}x_{6}^{2}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{10}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,1,0,0,1,1,1)(2,1,0,0,-1,-1,-1) with block multiplicities $(1,1,2,3)$; hence,
$C_{G}(H)$ $=\Bigl\{\ \operatorname{diag}(\alpha)\oplus\operatorname{diag}(\beta)\oplus A\oplus B:\ \alpha,\beta\in\mathbb{G}_{m},\ A\in\mathrm{GL}(2),\ B\in\mathrm{GL}(3),\ \alpha\beta\det(A)\det(B)=1\ \Bigr\}$
$\cong\bigl(\mathbb{G}_{m}\times\mathbb{G}_{m}\times\mathrm{GL}(2)\times\mathrm{GL}(3)\bigr)\cap\mathrm{SL}(7).$ (10)
Every monomial of $\phi_{10}$ has $H$-weight 0, so $\phi_{10}\in W^{H}$.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{10}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating within the $\mathrm{GL}(2)$- and $\mathrm{GL}(3)$-blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken as
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha,\,\beta,\,s+u,\,s-u,\,\gamma+v_{1},\,\gamma+v_{2},\,\gamma+v_{3}),$
where $\alpha,\beta,s,u,\gamma,v_{i}\in\mathbb{Z}$ with $v_{1}+v_{2}+v_{3}=0$, and by Convention 4.6 the $\mathrm{SL}$-constraint is
$S:=\alpha+\beta+2s+3\gamma=0.$
A direct computation gives the positive linear identity
$\bigl[w(x_{2}^{3})+w(x_{2}^{2}x_{3})+w(x_{2}x_{3}^{2})+w(x_{3}^{3})\bigr]$ (11)
$\quad+2\bigl[w(x_{0}x_{4}^{2})+w(x_{0}x_{4}x_{5})+w(x_{0}x_{5}^{2})+w(x_{0}x_{4}x_{6})+w(x_{0}x_{5}x_{6})+w(x_{0}x_{6}^{2})\bigr]$
$\quad+2\bigl[w(x_{1}x_{2}x_{4})+w(x_{1}x_{3}x_{4})+w(x_{1}x_{2}x_{5})+w(x_{1}x_{3}x_{5})+w(x_{1}x_{2}x_{6})+w(x_{1}x_{3}x_{6})\bigr]\;=12\,S.$
If $\lambda\in\Lambda_{\phi_{10}}$, then all the weights above are $\geq 0$ and $S=0$; hence, by (11), they all vanish. Solving yields
$\beta=-\gamma,\qquad s=0,\qquad u=0,\qquad v_{1}=v_{2}=v_{3}=0,\qquad\alpha=-2\gamma.$
Writing $\gamma=-k$ with $k\in\mathbb{Z}$ we obtain
$\mu_{k}(t):=\operatorname{diag}\!\bigl(t^{2k},\,t^{k},\,1,\,1,\,t^{-k},\,t^{-k},\,t^{-k}\bigr),\qquad\Lambda_{\phi_{10}}=\{\mu_{k}\mid k\in\mathbb{Z}\}.$
Thus $\Lambda_{\phi_{10}}$ is symmetric, and by the Casimiro–Florentino criterion, $\phi_{10}$ is polystable; in particular $\mathrm{SL}(7)\cdot\phi_{10}$ is closed.
Normal form and component dimension.
Set
$U:=\langle x_{2},x_{3}\rangle,\qquad V:=\langle x_{4},x_{5},x_{6}\rangle.$
Then the $H$-fixed subspace decomposes as
$W^{H}=\underbrace{\mathrm{Sym}^{3}U}_{(\mathrm{I})}\ \oplus\ \underbrace{x_{1}\otimes U\otimes V}_{(\mathrm{II})}\ \oplus\ \underbrace{x_{0}\otimes\mathrm{Sym}^{2}V}_{(\mathrm{III})}.$
We normalize a general element of $W^{H}$ under the $C_{G}(H)$-action, taking care that the $\mathrm{GL}(U)$-freedom is shared by blocks (I) and (II).
(III) The quadratic block $x_{0}\otimes\mathrm{Sym}^{2}V$.
For a general element, the quadratic form on $V$ is nondegenerate. Acting by $\mathrm{GL}(V)\subset C_{G}(H)$ and absorbing an overall scalar into the $x_{0}$-factor (using the central torus/projective scaling), we may assume the quadratic block is
$x_{0}\,q,\qquad q=x_{4}x_{5}+x_{6}^{2}.$
After this normalization, the residual right group on $V$ is the orthogonal group
$\mathrm{O}(V,q)=\{\,B\in\mathrm{GL}(V)\mid q(Bv)=q(v)\,\}.$
(II) The mixed block $x_{1}\otimes U\otimes V$.
Write the mixed term as $x_{1}m$ with $m\in U\otimes V$, and represent $m$ by a $2\times 3$ matrix after choosing bases of $U$ and $V$. For a general element, $\mathrm{rank}(m)=2$ and the restriction $q|_{\mathrm{Im}(m)}$ is nondegenerate. Equivalently, pulling back $q$ along $m$ yields a nondegenerate binary quadratic form on $U$:
$q_{m}(u):=q(m(u))\in\mathrm{Sym}^{2}U^{\vee}.$
Using $\mathrm{GL}(U)$ we may put $q_{m}$ into split form. Then, by Witt’s extension theorem, there exists $B\in\mathrm{O}(V,q)$ sending the resulting hyperbolic pair in $\mathrm{Im}(m)$ to $\langle x_{4},x_{5}\rangle$. Hence, after the $\mathrm{GL}(U)\times\mathrm{O}(V,q)$-action and scaling $x_{1}$, we may assume
$x_{1}m=x_{1}(x_{2}x_{5}+x_{3}x_{4}).$
(If $\mathrm{rank}(m)=1$ or $q|_{\mathrm{Im}(m)}$ is degenerate, one lands in a proper closed subset; for the generic closed orbit in this component we work on the open set above.)
Residual symmetry after (II) and (III).
After fixing (II) and (III) as above, the stabilizer contains the one-dimensional torus
$T=\Bigl\{\ x_{2}\mapsto t\,x_{2},\ x_{3}\mapsto t^{-1}x_{3},\ x_{4}\mapsto t\,x_{4},\ x_{5}\mapsto t^{-1}x_{5},\ x_{6}\mapsto x_{6}\ \Bigr\}\subset C_{G}(H),\qquad t\in\mathbb{G}_{m},$
which preserves both $x_{2}x_{5}+x_{3}x_{4}$ and $x_{4}x_{5}+x_{6}^{2}$.
(I) The cubic block $\mathrm{Sym}^{3}U$.
A general element of $\mathrm{Sym}^{3}U$ is
$c_{0}x_{2}^{3}+c_{1}x_{2}^{2}x_{3}+c_{2}x_{2}x_{3}^{2}+c_{3}x_{3}^{3}.$
Under $T$, these monomials have weights $t^{3},t^{1},t^{-1},t^{-3}$, respectively. On the dense open set where $c_{0}c_{3}\neq 0$, using the $T$-parameter and projective scaling we normalize $c_{0}=c_{3}=1$. Writing $\tau:=c_{1}$ and $\rho:=c_{2}$, we obtain
$x_{2}^{3}+\tau\,x_{2}^{2}x_{3}+\rho\,x_{2}x_{3}^{2}+x_{3}^{3}.$
Combining (I)–(III), we may take the closed-orbit representative in Case $k=10$ to be
$\varphi^{\mathrm{nf}}_{10}(\tau,\rho)=x_{2}^{3}+\tau\,x_{2}^{2}x_{3}+\rho\,x_{2}x_{3}^{2}+x_{3}^{3}+x_{1}x_{2}x_{5}+x_{1}x_{3}x_{4}+x_{0}x_{4}x_{5}+x_{0}x_{6}^{2},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$
As $\dim W^{H}=16$ and $\dim C_{G}(H)=14$, the effective action has dimension $13$; after projectivizing, we obtain 16131=216-13-1=2. The residual $T$-stabilizer of a general representative is finite; hence, the corresponding component $\Phi_{10}$ of the moduli is two-dimensional.
4.11 Case $k=11$
1-PS limit.
Set
$\lambda_{11}(t)=\operatorname{diag}\!\bigl(t^{2},\,1,\,1,\,1,\,1,\,t^{-1},\,t^{-1}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{11}$ as in Section 3, the $1$-PS limit is
$\phi_{11}:=\lim_{t\to 0}\lambda_{11}(t)\cdot f_{11}=F(x_{1},x_{2},x_{3},x_{4})+x_{0}\,Q(x_{5},x_{6}),$
where $F\in\mathrm{Sym}^{3}\langle x_{1},x_{2},x_{3},x_{4}\rangle$ is a generic cubic form and $Q\in\mathrm{Sym}^{2}\langle x_{5},x_{6}\rangle$ is a generic quadratic form.
$H$ and $C_{G}(H)$.
Let $H=\lambda_{11}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,0,0,0,0,1,1)(2,0,0,0,0,-1,-1) with block decomposition $\langle x_{0}\rangle\oplus\langle x_{1},x_{2},x_{3},x_{4}\rangle\oplus\langle x_{5},x_{6}\rangle$. Hence,
$C_{G}(H)$ $=\Bigl\{\ \operatorname{diag}(\alpha)\oplus A\oplus B\ :\ \alpha\in\mathbb{G}_{m},\ A\in\mathrm{GL}(4),\ B\in\mathrm{GL}(2),\ \alpha\,\det(A)\det(B)=1\ \Bigr\}$
$\cong\bigl(\mathbb{G}_{m}\times\mathrm{GL}(4)\times\mathrm{GL}(2)\bigr)\cap\mathrm{SL}(7).$ (12)
Every monomial of $\phi_{11}$ has $H$-weight 0, so $\phi_{11}\in W^{H}$.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{11}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating inside the $\mathrm{GL}(4)$- and $\mathrm{GL}(2)$-blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha,\,s_{1},\,s_{2},\,s_{3},\,s_{4},\,t+u,\,t-u),\qquad S:=\alpha+(s_{1}+s_{2}+s_{3}+s_{4})+2t=0$
(Convention 4.6). Summing the $\lambda$-weights of the $20$ cubic monomials in $F$ gives $15(s_{1}+s_{2}+s_{3}+s_{4})$, while the weights of the quadratic terms $x_{0}Q$ sum to $3\alpha+6t$. Using the relation $S=0$, we obtain the positive linear identity
$w(F)+5\,w(x_{0}Q)\;=\;15(s_{1}+s_{2}+s_{3}+s_{4})+15(\alpha+2t)\;=\;15\,S.$ (13)
If $\lambda\in\Lambda_{\phi_{11}}$, then all weights are $\geq 0$ and $S=0$; by (13) they all vanish. Solving this implies that $\Lambda_{\phi_{11}}$ is symmetric, and by the Casimiro–Florentino criterion, $\phi_{11}$ is polystable.
Normal form and component dimension.
We normalize generic elements under the action of the centralizer $C_{G}(H)$. The $H$-fixed subspace decomposes as
$W^{H}=\mathrm{Sym}^{3}\langle x_{1},x_{2},x_{3},x_{4}\rangle\ \oplus\ x_{0}\otimes\mathrm{Sym}^{2}\langle x_{5},x_{6}\rangle.$
(1) Quadratic part. Under the $\mathrm{GL}(2)$-action on $\langle x_{5},x_{6}\rangle$, a generic binary quadratic is normalized to $x_{5}^{2}+x_{6}^{2}$. The stabilizer of this form contains the orthogonal group $\mathrm{O}(2)$, contributing $1$ dimension to the stabilizer of the generic point.
(2) Cubic part. Under the $\mathrm{GL}(4)$-action on $\langle x_{1},x_{2},x_{3},x_{4}\rangle$, a generic cubic surface can be written as the sum of five cubes of linear forms (Sylvester’s Pentahedron Theorem). Let $L_{0},\dots,L_{4}$ be fixed general linear forms in $x_{1},\dots,x_{4}$. Then the cubic part is $\sum_{i=0}^{4}c_{i}L_{i}^{3}$.
Thus, a closed-orbit representative is:
$\varphi_{11}^{\mathrm{nf}}(c)=\sum_{i=0}^{4}c_{i}L_{i}(x_{1},x_{2},x_{3},x_{4})^{3}+x_{0}(x_{5}^{2}+x_{6}^{2}),$
where the parameters are $c_{0}:\dots:c_{4}]\in\mathbb{P}^{4}$.
The dimension of the component is computed as follows: $\dim W^{H}=20(\text{cubic})+3(\text{quadratic})=23$. The effective group dimension is $\dim C_{G}(H)-1=19$ (since $H$ acts trivially). Since the generic stabilizer has dimension $1$ (due to $\mathrm{O}(2)$), the orbit dimension is 191=1819-1=18.
$\dim(\Phi_{11})=23-18-1=4.$
(Note: This case is structurally isomorphic to Case $k=21$.)
4.12 Case $k=12$
1-PS limit.
Set
$\lambda_{12}(t)=\operatorname{diag}\!\bigl(t^{3},\,t^{2},\,t,\,t,\,t^{-1},\,t^{-2},\,t^{-4}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{12}$ as in Section 3, the $1$-PS limit is
$\phi_{12}:=\lim_{t\to 0}\lambda_{12}(t)\cdot f_{12}=a_{1}x_{1}x_{4}^{2}+a_{2}x_{2}^{2}x_{5}+a_{3}x_{2}x_{3}x_{5}+a_{4}x_{3}^{2}x_{5}+a_{5}x_{0}x_{4}x_{5}+a_{6}x_{1}^{2}x_{6}+a_{7}x_{0}x_{2}x_{6}+a_{8}x_{0}x_{3}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{12}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (3,2,1,1,1,2,4)(3,2,1,1,-1,-2,-4) with multiplicities $(1,1,2,1,1,1)$; hence,
$C_{G}(H)$ $=\Bigl\{\ \operatorname{diag}(\alpha)\oplus\operatorname{diag}(\beta)\oplus A\oplus\operatorname{diag}(\gamma,\delta,\varepsilon)\ :\ \alpha,\beta,\gamma,\delta,\varepsilon\in\mathbb{G}_{m},\ A\in\mathrm{GL}(2),$
$\qquad\alpha\beta\det(A)\gamma\delta\varepsilon=1\ \Bigr\}\ \cong\ \bigl(\mathbb{G}_{m}^{5}\times\mathrm{GL}(2)\bigr)\cap\mathrm{SL}(7),\qquad\dim C_{G}(H)=8.$ (14)
Every monomial of $\phi_{12}$ has $H$-weight 0.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{12}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating inside the $\mathrm{GL}(2)$-block on $\langle x_{2},x_{3}\rangle$, any $\lambda\in Y(C_{G}(H))$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha,\,\beta,\,s+u,\,s-u,\,\gamma,\,\delta,\,\varepsilon),\qquad S:=\alpha+\beta+2s+\gamma+\delta+\varepsilon=0.$
A direct computation yields the positive identity
$2w(x_{1}x_{4}^{2})+w(x_{2}^{2}x_{5})+w(x_{2}x_{3}x_{5})+2w(x_{3}^{2}x_{5})+2w(x_{0}x_{4}x_{5})+2w(x_{1}^{2}x_{6})$
$\qquad+3w(x_{0}x_{2}x_{6})+w(x_{0}x_{3}x_{6})=6S.$ (15)
If $\lambda\in\Lambda_{\phi_{12}}$, then all the weights on the left are $\geq 0$ and $S=0$; by (15) they all vanish. Solving gives
$u=0,\qquad\beta=-2\gamma,\qquad\delta=-2s,\qquad\varepsilon=-2\beta=4\gamma,\qquad\alpha=-3\gamma,\qquad s=-\gamma.$
Writing $k=-\gamma\in\mathbb{Z}$, we obtain
$\mu_{k}(t)=\operatorname{diag}\!\bigl(t^{3k},\,t^{2k},\,t^{k},\,t^{k},\,t^{-k},\,t^{-2k},\,t^{-4k}\bigr),\qquad\Lambda_{\phi_{12}}=\{\mu_{k}\mid k\in\mathbb{Z}\}$
Thus $\Lambda_{\phi_{12}}$ is symmetric, and by the Casimiro–Florentino criterion $\phi_{12}$ is polystable; in particular, $\mathrm{SL}(7)\cdot\phi_{12}$ is closed.
Normal form and component dimension.
Work on the $H$-fixed slice
$W^{H}=\underbrace{x_{1}\otimes\mathrm{Sym}^{2}\langle x_{4}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{\mathrm{Sym}^{2}\langle x_{2},x_{3}\rangle\otimes x_{5}}_{(\mathrm{II})}\ \oplus\ \underbrace{x_{0}\otimes\langle x_{4}\rangle\otimes x_{5}}_{(\mathrm{III})}\ \oplus\ \underbrace{\mathrm{Sym}^{2}\langle x_{1}\rangle\otimes x_{6}}_{(\mathrm{IV})}$
[…]
Diagonalize the binary quadratic in the $(\mathrm{II})$-block.
Using $\mathrm{GL}(\langle x_{2},x_{3}\rangle)$, bring $\mathrm{Sym}^{2}\langle x_{2},x_{3}\rangle\otimes x_{5}$ to $x_{2}^{2}x_{5}+x_{3}^{2}x_{5}$; the cross term $x_{2}x_{3}x_{5}$ is eliminated.
Align the $(\mathrm{V})$-block.
Within $x_{0}\otimes\langle x_{2},x_{3}\rangle\otimes x_{6}$, use the same $\mathrm{GL}(2)$ to align this block to $x_{0}x_{2}x_{6}$ (so the $x_{0}x_{3}x_{6}$ entry vanishes).
Normalize coefficients by torus scalings and projective scaling in $(\mathrm{I})$, $(\mathrm{III})$, $(\mathrm{V})$.
Use the $1$-dimensional tori on the $1$-dimensional weight spaces and the overall projective scaling to set the remaining nonzero coefficients to $1$.
This yields the closed orbit normal form
$\varphi^{\mathrm{nf}}_{12}=x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{3}^{2}x_{5}+x_{0}x_{4}x_{5}+x_{1}^{2}x_{6}+x_{0}x_{2}x_{6}.$
Finally, $\dim W^{H}=8$ and $\dim C_{G}(H)=8$; the effective action has dimension $7$ (with $H$ acting trivially). After projectivizing, we obtain 871=08-7-1=0; hence, the corresponding boundary component is zero-dimensional.
4.13 Case $k=13$
1-PS limit.
Set
$\lambda_{13}(t)=\operatorname{diag}\!\bigl(t^{2},\,t,\,t,\,1,\,t^{-1},\,t^{-1},\,t^{-2}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{13}$ as in Section 3, the $1$-PS limit is
$\phi_{13}=\lim_{t\to 0}\lambda_{13}(t)\cdot f_{13}$ $=a_{1}x_{3}^{3}+a_{2}x_{1}x_{3}x_{4}+a_{3}x_{2}x_{3}x_{4}+a_{4}x_{0}x_{4}^{2}+a_{5}x_{1}x_{3}x_{5}+a_{6}x_{2}x_{3}x_{5}$
$\quad+a_{7}x_{0}x_{4}x_{5}+a_{8}x_{0}x_{5}^{2}+a_{9}x_{1}^{2}x_{6}+a_{10}x_{1}x_{2}x_{6}+a_{11}x_{2}^{2}x_{6}+a_{12}x_{0}x_{3}x_{6}.$
This is $H$-fixed for $H=\lambda_{13}(\mathbb{G}_{m})$.
$H$ and $C_{G}(H)$.
The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,1,1,0,1,1,2)(2,1,1,0,-1,-1,-2) with blocks $\langle x_{0}\rangle\oplus\langle x_{1},x_{2}\rangle\oplus\langle x_{3}\rangle\oplus\langle x_{4},x_{5}\rangle\oplus\langle x_{6}\rangle$. Hence,
$C_{G}(H)$ $=\Bigl\{\ \operatorname{diag}(\alpha)\ \oplus\ A\ \oplus\ \operatorname{diag}(\beta)\ \oplus\ B\ \oplus\ \operatorname{diag}(\gamma):\ \alpha,\beta,\gamma\in\mathbb{G}_{m},\ A,B\in\mathrm{GL}(2),$
$\qquad\qquad\qquad\qquad\qquad\alpha\,\det(A)\,\beta\,\det(B)\,\gamma=1\,\Bigr\}$
$\cong\bigl(\mathbb{G}_{m}\times\mathrm{GL}(2)\times\mathbb{G}_{m}\times\mathrm{GL}(2)\times\mathbb{G}_{m}\bigr)\cap\mathrm{SL}(7),\qquad\dim C_{G}(H)=10.$ (16)
Each monomial of $\phi_{13}$ has $H$-weight 0.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{13}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating within the two $\mathrm{GL}(2)$-blocks, any $\lambda\in Y(C_{G}(H))$ may be taken as
[…]
A direct computation yields the positive identity
$\bigl[w(x_{1}^{2}x_{6})+w(x_{2}^{2}x_{6})+2\,w(x_{1}x_{2}x_{6})\bigr]+\bigl[w(x_{0}x_{4}^{2})+2\,w(x_{0}x_{4}x_{5})+w(x_{0}x_{5}^{2})\bigr]$
$\qquad+\bigl[w(x_{1}x_{3}x_{4})+w(x_{2}x_{3}x_{4})+w(x_{1}x_{3}x_{5})+w(x_{2}x_{3}x_{5})\bigr]+2\,w(x_{0}x_{3}x_{6})=6S.$ (17)
If $\lambda\in\Lambda_{\phi_{13}}$, then the twelve weights on the left are $\geq 0$ and $S=0$; hence, they all vanish. Solving gives
$u=0,\quad v=0,\quad\alpha+2t=0,\quad 2s+\gamma=0,\quad s+\beta+t=0,\quad\alpha+\beta+\gamma=0.$
Writing $s=k\in\mathbb{Z}$ yields
$(\alpha,s,\beta,t,\gamma)=(2k,\,k,\,0,\,-k,\,-2k),\qquad\mu_{k}(t):=\operatorname{diag}\!\bigl(t^{2k},\,t^{k},\,t^{k},\,1,\,t^{-k},\,t^{-k},\,t^{-2k}\bigr),$
so $\Lambda_{\phi_{13}}=\{\mu_{k}\mid k\in\mathbb{Z}\}$ is symmetric; by the Casimiro–Florentino criterion $\phi_{13}$ is polystable; hence, $\mathrm{SL}(7)\cdot\phi_{13}$ is closed.
Normal form and component dimension.
We display
$W^{H}$ $=\underbrace{\mathrm{Sym}^{3}\langle x_{3}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{x_{0}\otimes\mathrm{Sym}^{2}\langle x_{4},x_{5}\rangle}_{(\mathrm{II})}\ \oplus\ \underbrace{\langle x_{1},x_{2}\rangle\otimes\langle x_{4},x_{5}\rangle\otimes\langle x_{3}\rangle}_{(\mathrm{III})}$
$\ \ \ \oplus\ \underbrace{\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle\otimes\langle x_{6}\rangle}_{(\mathrm{IV})}\ \oplus\ \underbrace{x_{0}\otimes\langle x_{3}\rangle\otimes x_{6}}_{(\mathrm{V})}.$
Normalize block by block under the $C_{G}(H)$-action:
(II) $x_{0}\otimes\mathrm{Sym}^{2}\langle x_{4},x_{5}\rangle$.
Diagonalize the binary quadratic to
$x_{0}x_{4}^{2}+x_{0}x_{5}^{2}.$
(III) $\langle x_{1},x_{2}\rangle\otimes\langle x_{4},x_{5}\rangle\otimes\langle x_{3}\rangle$.
View the four $x_{3}$-bilinear terms as a $2\times 2$ matrix on $\langle x_{1},x_{2}\rangle\otimes\langle x_{4},x_{5}\rangle$ and bring it to diagonal form,
$x_{1}x_{3}x_{4}+\rho\,x_{2}x_{3}x_{5},\qquad\rho\in\mathbb{C}^{\times}.$
(IV) $\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle\otimes\langle x_{6}\rangle$.
Diagonalize the symmetric $2\times 2$ form to
$x_{1}^{2}x_{6}+\sigma\,x_{2}^{2}x_{6},\qquad\sigma\in\mathbb{C}^{\times}.$
(I) and (V).
Use the three torus factors on $x_{0},x_{3},x_{6}$ together with projective scaling to normalize the remaining nonzero coefficients to $1$.
This yields the convenient normal form
$\varphi^{\mathrm{nf}}_{13}(\rho,\sigma)=x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{5}^{2}+x_{1}x_{3}x_{4}+\rho\,x_{2}x_{3}x_{5}+x_{1}^{2}x_{6}+\sigma\,x_{2}^{2}x_{6}+x_{0}x_{3}x_{6},\qquad(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}.$
Here, $\dim(W^{H})=12$ and $\dim C_{G}(H)=10$; as $H$ acts trivially, the effective group dimension is $9$. After projectivizing, we obtain 1291=212-9-1=2. The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{13}$ of the moduli is two-dimensional.
4.14 Case $k=14$
1-PS limit.
Set
$\lambda_{14}(t)=\operatorname{diag}\!\bigl(t^{2},\,t^{2},\,1,\,t^{-1},\,t^{-1},\,t^{-1},\,t^{-1}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{14}$ as in Section 3, the $1$-PS limit is
$\phi_{14}=\lim_{t\to 0}\lambda_{14}(t)\cdot f_{14}$ $=a_{1}x_{2}^{3}+a_{2}x_{0}x_{3}^{2}+a_{3}x_{1}x_{3}^{2}+a_{4}x_{0}x_{3}x_{4}+a_{5}x_{1}x_{3}x_{4}+a_{6}x_{0}x_{4}^{2}+a_{7}x_{1}x_{4}^{2}$
$\quad+a_{8}x_{0}x_{3}x_{5}+a_{9}x_{1}x_{3}x_{5}+a_{10}x_{0}x_{4}x_{5}+a_{11}x_{1}x_{4}x_{5}+a_{12}x_{0}x_{5}^{2}+a_{13}x_{1}x_{5}^{2}$
$\quad+a_{14}x_{0}x_{3}x_{6}+a_{15}x_{1}x_{3}x_{6}+a_{16}x_{0}x_{4}x_{6}+a_{17}x_{1}x_{4}x_{6}$
$\quad+a_{18}x_{0}x_{5}x_{6}+a_{19}x_{1}x_{5}x_{6}+a_{20}x_{0}x_{6}^{2}+a_{21}x_{1}x_{6}^{2}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{14}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,2,0,1,1,1,1)(2,2,0,-1,-1,-1,-1) with block multiplicities $(2,1,4)$. Hence,
$C_{G}(H)$ $=\Bigl\{\,A\oplus\beta\oplus B:\ A\in\mathrm{GL}(2),\ \beta\in\mathbb{G}_{m},\ B\in\mathrm{GL}(4),\ \det(A)\,\beta\,\det(B)=1\,\Bigr\}$
[…]
Every monomial of $\phi_{14}$ has $H$-weight 0; hence, $\phi_{14}\in W^{H}$.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{14}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating inside the $\mathrm{GL}(2)$- and $\mathrm{GL}(4)$-blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(a+u,\,a-u,\,b,\,c+v_{1},\,c+v_{2},\,c+v_{3},\,c+v_{4}),$
where $v_{1}+v_{2}+v_{3}+v_{4}=0$, and with the $\mathrm{SL}$-constraint (Convention 4.6)
$S:=2a+b+4c=0.$
A direct computation yields the positive identity
[…]
If $\lambda\in\Lambda_{\phi_{14}}$, then all $21$ weights are $\geq 0$ and $S=0$; hence, by (18) they all vanish. Solving gives
$b=0,\qquad u=0,\qquad v_{1}=v_{2}=v_{3}=v_{4}=0,\qquad a+2c=0.$
Writing c=kc=-k with $k\in\mathbb{Z}$, we obtain
$\mu_{k}(t)=\operatorname{diag}\!\bigl(t^{2k},\,t^{2k},\,1,\,t^{-k},\,t^{-k},\,t^{-k},\,t^{-k}\bigr),\qquad\Lambda_{\phi_{14}}=\{\mu_{k}\mid k\in\mathbb{Z}\}.$
Thus $\Lambda_{\phi_{14}}$ is symmetric, and by the Casimiro–Florentino criterion $\phi_{14}$ is polystable; in particular $\mathrm{SL}(7)\cdot\phi_{14}$ is closed.
Normal form and component dimension.
We display $W^{H}$ in block form:
$W^{H}$ $=\underbrace{\mathrm{Sym}^{3}\langle x_{2}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{\langle x_{0}\rangle\otimes\mathrm{Sym}^{2}\langle x_{3},x_{4},x_{5},x_{6}\rangle}_{(\mathrm{II})}\ \oplus\ \underbrace{\langle x_{1}\rangle\otimes\mathrm{Sym}^{2}\langle x_{3},x_{4},x_{5},x_{6}\rangle}_{(\mathrm{III})}.$
Equivalently, writing $U:=\langle x_{3},x_{4},x_{5},x_{6}\rangle$ and $V:=\langle x_{0},x_{1}\rangle$, any element of $W^{H}$ can be written as
$\phi\;=\;a\,x_{2}^{3}\;+\;x_{0}Q_{0}(x_{3},x_{4},x_{5},x_{6})\;+\;x_{1}Q_{1}(x_{3},x_{4},x_{5},x_{6}),\qquad Q_{0},Q_{1}\in\mathrm{Sym}^{2}U^{\vee}.$
Using the $\mathbb{G}_{m}$-factor on $x_{2}$ together with projective scaling, we normalize $a=1$.
Diagonalization of the quadratic pencil. Consider the pencil $\langle Q_{0},Q_{1}\rangle$, i.e. the binary quartic
$\Delta(s,t):=\det(sQ_{0}+tQ_{1}),\qquad[s:t]\in\mathbb{P}^{1}.$
On the dense open set where $\Delta$ has four distinct roots (equivalently, the pencil is regular, i.e. it lies off the discriminant locus), we may assume $Q_{0}$ is nondegenerate. Acting by $\mathrm{GL}(U)\cong\mathrm{GL}(4)$, we take
$Q_{0}\sim q:=x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}.$
With respect to $q$, the residual group is $O(U,q)$, and for a regular pencil the $q$-selfadjoint endomorphism $A$ defined by $Q_{1}(u,v)=q(Au,v)$ has four distinct eigenvalues. Hence, after an $O(U,q)$-change of coordinates,
$Q_{1}\sim\lambda_{3}x_{3}^{2}+\lambda_{4}x_{4}^{2}+\lambda_{5}x_{5}^{2}+\lambda_{6}x_{6}^{2},\qquad\lambda_{3},\lambda_{4},\lambda_{5},\lambda_{6}\ \text{pairwise distinct}.$
Reduction to a one-parameter normal form (cross-ratio). The $\mathrm{GL}(V)\cong\mathrm{GL}(2)$-action on $(x_{0},x_{1})$ replaces $(Q_{0},Q_{1})$ by
$(Q_{0}^{\prime},Q_{1}^{\prime})=(\alpha Q_{0}+\beta Q_{1},\ \gamma Q_{0}+\delta Q_{1}),\qquad\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}\in\mathrm{GL}(2),$
and (after re-normalizing $Q_{0}^{\prime}$ back to $q$ by a diagonal rescaling in $U$) this transforms the eigenvalues by the fractional-linear map
$\lambda\ \longmapsto\ \frac{\gamma+\delta\lambda}{\alpha+\beta\lambda}.$
Thus the regular pencil is determined, up to permutation of the diagonal coordinates in $U$, by the unordered set $\{\lambda_{3},\lambda_{4},\lambda_{5},\lambda_{6}\}\subset\mathbb{P}^{1}$ modulo $\mathrm{PGL}_{2}$, i.e. by a single cross-ratio parameter.
To match the normal form displayed below, we normalize three eigenvalues to 1,1,21,-1,2 and denote the fourth by $\tau$. Using the $3$-transitivity of $\mathrm{PGL}_{2}$ on $\mathbb{P}^{1}$ and then permuting $(x_{3},x_{4},x_{5},x_{6})$, we may assume
$(\lambda_{3},\lambda_{4},\lambda_{5},\lambda_{6})=(1,\tau,-1,2).$
Consequently (on the dense open set $\tau\notin\{1,-1,2\}$; and restricting further to $\tau\in\mathbb{C}^{\times}$ if we also require $Q_{1}$ nondegenerate), a convenient normal form is
$\varphi^{\mathrm{nf}}_{14}(\tau)=x_{2}^{3}+x_{0}\bigl(x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)+x_{1}\bigl(x_{3}^{2}+\tau x_{4}^{2}-x_{5}^{2}+2x_{6}^{2}\bigr),\qquad\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}.$
(For the special values $\tau\in\{1,-1,2\}$ the pencil acquires repeated eigenvalues and the stabilizer jumps; these form a proper closed subset of the same one-dimensional family.)
As $\dim W^{H}=21$ and $\dim C_{G}(H)=20$, and since $H\simeq\mathbb{G}_{m}$ acts trivially on $W^{H}$, the effective action has dimension $19$; after projectivizing we obtain 21191=121-19-1=1. For a general parameter $\tau$ the residual stabilizer is finite; hence, the corresponding component $\Phi_{14}$ of the moduli is one-dimensional.
4.15 Case $k=15$
1-PS limit.
Set
$\lambda_{15}(t)=\operatorname{diag}\!\bigl(t^{2},\,t,\,t,\,1,\,1,\,t^{-2},\,t^{-2}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{15}$ as in Section 3, the $1$-PS limit is
$\phi_{15}=\lim_{t\to 0}\lambda_{15}(t)\cdot f_{15}$ $=a_{1}x_{3}^{3}+a_{2}x_{3}^{2}x_{4}+a_{3}x_{3}x_{4}^{2}+a_{4}x_{4}^{3}+a_{5}x_{1}^{2}x_{5}+a_{6}x_{1}x_{2}x_{5}+a_{7}x_{2}^{2}x_{5}$
$\quad+a_{8}x_{0}x_{3}x_{5}+a_{9}x_{0}x_{4}x_{5}+a_{10}x_{1}^{2}x_{6}+a_{11}x_{1}x_{2}x_{6}+a_{12}x_{2}^{2}x_{6}$
$\quad+a_{13}x_{0}x_{3}x_{6}+a_{14}x_{0}x_{4}x_{6}.$
All monomials have $H$-weight 0, so $\phi_{15}\in W^{H}$ for $H=\lambda_{15}(\mathbb{G}_{m})$.
$H$ and $C_{G}(H)$.
The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,1,1,0,0,2,2)(2,1,1,0,0,-2,-2) with block decomposition $\langle x_{0}\rangle\oplus\langle x_{1},x_{2}\rangle\oplus\langle x_{3},x_{4}\rangle\oplus\langle x_{5},x_{6}\rangle$. Hence,
$C_{G}(H)$ $=\Bigl\{\ \operatorname{diag}(\alpha)\oplus A\oplus B\oplus C:\ \alpha\in\mathbb{G}_{m},\ A,B,C\in\mathrm{GL}(2),\ \alpha\,\det(A)\det(B)\det(C)=1\ \Bigr\}$
[…]
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{15}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating inside the three $\mathrm{GL}(2)$-blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha,\,s+u,\,s-u,\,t+v,\,t-v,\,\gamma+w,\,\gamma-w),\qquad S:=\alpha+2s+2t+2\gamma=0.$
A direct computation yields the positive identity
$2\bigl[w(x_{3}^{3})+w(x_{3}^{2}x_{4})+w(x_{3}x_{4}^{2})+w(x_{4}^{3})\bigr]+3\Bigl[\bigl(w(x_{1}^{2}x_{5})+2w(x_{1}x_{2}x_{5})+w(x_{2}^{2}x_{5})\bigr)$
$\qquad\qquad\qquad\qquad\qquad\qquad+\bigl(w(x_{1}^{2}x_{6})+2w(x_{1}x_{2}x_{6})+w(x_{2}^{2}x_{6})\bigr)\Bigr]$
$\qquad\qquad\qquad\qquad\qquad\qquad+6\bigl[w(x_{0}x_{3}x_{5})+w(x_{0}x_{4}x_{5})+w(x_{0}x_{3}x_{6})+w(x_{0}x_{4}x_{6})\bigr]\;=\;24\,S.$ (19)
If $\lambda\in\Lambda_{\phi_{15}}$, then all $14$ weights above are $\geq 0$ and $S=0$; hence, by (19) they all vanish. Solving gives
$u=0,\quad v=0,\quad w=0,\quad t=0,\quad\alpha+\gamma=0,\quad 2s+\gamma=0.$
Writing $s=k\in\mathbb{Z}$ we obtain
$\mu_{k}(t):=\operatorname{diag}\!\bigl(t^{2k},\,t^{k},\,t^{k},\,1,\,1,\,t^{-2k},\,t^{-2k}\bigr),\qquad\Lambda_{\phi_{15}}=\{\mu_{k}\mid k\in\mathbb{Z}\}.$
Thus $\Lambda_{\phi_{15}}$ is symmetric; by the Casimiro–Florentino criterion $\phi_{15}$ is polystable, so $\mathrm{SL}(7)\cdot\phi_{15}$ is closed.
Normal form and component dimension.
We have
$W^{H}=\underbrace{\mathrm{Sym}^{3}\langle x_{3},x_{4}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{\bigl(\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle\bigr)\otimes\langle x_{5},x_{6}\rangle}_{(\mathrm{II})}\ \oplus\ \underbrace{x_{0}\otimes\langle x_{3},x_{4}\rangle\otimes\langle x_{5},x_{6}\rangle}_{(\mathrm{III})},$
of respective dimensions $4$, 66, and $4$ (total $\dim W^{H}=14$).
Reduction to normal form.
We now normalize $\varphi_{15}$ under the action of $C_{G}(H)$ on $W^{H}$, using only: (i) the left $\mathrm{GL}(2)$ on $\langle x_{3},x_{4}\rangle$, (ii) the left $\mathrm{Sym}^{2}\mathrm{GL}(2)$ on $\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle$, (iii) the right $\mathrm{GL}(2)$ on $\langle x_{5},x_{6}\rangle$, (iv) diagonal tori (subject to $\det=1$) and projective rescaling. We proceed block by block.
(I) The binary cubic block $\mathrm{Sym}^{3}\langle x_{3},x_{4}\rangle$.
A general binary cubic is $\mathrm{GL}(2)$-equivalent (after one overall scalar) to
$x_{3}^{2}x_{4}+\tau\,x_{3}x_{4}^{2},\qquad\tau\in\mathbb{C}^{\times},$
which fixes the $\mathrm{Sym}^{3}$-part up to the single modulus $\tau$.
(III) The bilinear $2\times 2$ block $x_{0}\otimes\langle x_{3},x_{4}\rangle\otimes\langle x_{5},x_{6}\rangle$.
Write this part as $x_{0}\,(x_{3},x_{4})\,M\,(x_{5},x_{6})^{\!\top}$ with $M\in M_{2\times 2}$. Using the left $\mathrm{GL}(2)$ action on $\langle x_{3},x_{4}\rangle$ and the right $\mathrm{GL}(2)$ action on $\langle x_{5},x_{6}\rangle$ (an SVD-type reduction), we bring $M$ to the identity; a diagonal torus and projective rescaling normalize the two coefficients to $1$:
$x_{0}x_{3}x_{5}+x_{0}x_{4}x_{6}.$
(II) The $3\times 2$ block $\bigl(\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle\bigr)\otimes\langle x_{5},x_{6}\rangle$.
Choose bases $\{x_{1}^{2},\ x_{1}x_{2},\ x_{2}^{2}\}$ and $\{x_{5},x_{6}\}$. The left action of $\mathrm{Sym}^{2}\mathrm{GL}(2)$ on $\mathrm{Sym}^{2}\langle x_{1},x_{2}\rangle$ together with the right action of $\mathrm{GL}(2)$ on $\langle x_{5},x_{6}\rangle$ allows a simultaneous reduction that eliminates the mixed $x_{1}x_{2}$ row and diagonalizes the remaining two rows. After using diagonal tori and an overall scale, we obtain
$x_{1}^{2}x_{5}+\rho\,x_{2}^{2}x_{6},\qquad\rho\in\mathbb{C}^{\times}.$
Normal form.
Hence, a closed-orbit representative is
$\varphi^{\mathrm{nf}}_{15}(\tau,\rho)\;=\;x_{3}^{2}x_{4}\;+\;\tau\,x_{3}x_{4}^{2}\;+\;x_{0}x_{3}x_{5}\;+\;x_{0}x_{4}x_{6}\;+\;x_{1}^{2}x_{5}\;+\;\rho\,x_{2}^{2}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$ (20)
For general $(\tau,\rho)$ the residual stabilizer in $C_{G}(H)$ is finite, so the parameters $(\tau,\rho)$ record genuine moduli on the closed stratum.
4.16 Case $k=16$
1-PS limit.
Set
$\lambda_{16}(t)=\operatorname{diag}\!\bigl(t^{2},\,t,\,1,\,1,\,1,\,t^{-1},\,t^{-2}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{16}$ as in Section 3, the $1$-PS limit is
$\phi_{16}=\lim_{t\to 0}\lambda_{16}(t)\cdot f_{16}$ $=a_{1}x_{2}^{3}+a_{2}x_{2}^{2}x_{3}+a_{3}x_{2}x_{3}^{2}+a_{4}x_{3}^{3}+a_{5}x_{2}^{2}x_{4}+a_{6}x_{2}x_{3}x_{4}+a_{7}x_{3}^{2}x_{4}$
$\quad+a_{8}x_{2}x_{4}^{2}+a_{9}x_{3}x_{4}^{2}+a_{10}x_{4}^{3}+a_{11}x_{1}x_{2}x_{5}+a_{12}x_{1}x_{3}x_{5}+a_{13}x_{1}x_{4}x_{5}$
$\quad+a_{14}x_{0}x_{5}^{2}+a_{15}x_{1}^{2}x_{6}+a_{16}x_{0}x_{2}x_{6}+a_{17}x_{0}x_{3}x_{6}+a_{18}x_{0}x_{4}x_{6}.$
(All monomials have $H$-weight 0, so $\phi_{16}\in W^{H}$ for $H=\lambda_{16}(\mathbb{G}_{m})$.)
$H$ and $C_{G}(H)$.
The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,1,0,0,0,1,2)(2,1,0,0,0,-1,-2) with block decomposition $\langle x_{0}\rangle\oplus\langle x_{1}\rangle\oplus\langle x_{2},x_{3},x_{4}\rangle\oplus\langle x_{5}\rangle\oplus\langle x_{6}\rangle$. Hence,
$C_{G}(H)$ $=\Bigl\{\ \operatorname{diag}(\alpha)\oplus\operatorname{diag}(\beta)\oplus B\oplus\operatorname{diag}(\delta)\oplus\operatorname{diag}(\varepsilon)\ :\ \alpha,\beta,\delta,\varepsilon\in\mathbb{G}_{m},\ B\in\mathrm{GL}(3),$
$\qquad\qquad\alpha\,\beta\,\det(B)\,\delta\,\varepsilon=1\ \Bigr\}$
[…]
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{16}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating within the $\mathrm{GL}(3)$-block on $\langle x_{2},x_{3},x_{4}\rangle$, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha,\,\beta,\,c+v_{1},\,c+v_{2},\,c+v_{3},\,\delta,\,\varepsilon),\qquad v_{1}+v_{2}+v_{3}=0,$
and, by Convention 4.6, with the $\mathrm{SL}$-constraint
$S:=\alpha+\beta+3c+\delta+\varepsilon=0.$
Writing $w(\cdot)$ for the $\lambda$-weight of a monomial, a direct calculation yields the positive identity
$\sum_{\begin{subarray}{c}i+j+k=3\\ i,j,k\geq 0\end{subarray}}w(x_{2}^{i}x_{3}^{j}x_{4}^{k})\;+\;6\,w(x_{0}x_{5}^{2})\;+\;6\,w(x_{1}^{2}x_{6})\;+\;2\sum_{i=2}^{4}w(x_{0}x_{i}x_{6})\;=\;12S.$ (21)
If $\lambda\in\Lambda_{\phi_{16}}$, then all $18$ weights on the left are $\geq 0$ and $S=0$; hence, by (21) they all vanish. From the ten cubics, we obtain
$c=0,\qquad v_{1}=v_{2}=v_{3}=0,$
and from the remaining terms
$\alpha+2\delta=0,\qquad 2\beta+\varepsilon=0,\qquad\alpha+\varepsilon=0.$
Thus $\alpha=2\beta,\ \delta=-\beta,\ \varepsilon=-2\beta$. Writing $\beta=k\in\mathbb{Z}$ gives
$\mu_{k}(t)=\operatorname{diag}\!\bigl(t^{2k},\,t^{k},\,1,\,1,\,1,\,t^{-k},\,t^{-2k}\bigr),\qquad\Lambda_{\phi_{16}}=\{\mu_{k}\mid k\in\mathbb{Z}\}.$
As $\Lambda_{\phi_{16}}$ is symmetric, $\phi_{16}$ is polystable by the Casimiro–Florentino criterion; in particular, $\mathrm{SL}(7)\cdot\phi_{16}$ is closed.
Normal form and component dimension.
On the HH–fixed slice we have the block decomposition
$W^{H}=\underbrace{\mathrm{Sym}^{3}\langle x_{2},x_{3},x_{4}\rangle}_{(\mathrm{I})}\;\oplus\;\underbrace{x_{1}\otimes\langle x_{2},x_{3},x_{4}\rangle\otimes x_{5}}_{(\mathrm{II})}\;\oplus\;\underbrace{\langle x_{0}x_{5}^{2}\rangle}_{(\mathrm{III})}\;\oplus\;\underbrace{\langle x_{1}^{2}x_{6}\rangle}_{(\mathrm{IV})}\;\oplus\;\underbrace{x_{0}\otimes\langle x_{2},x_{3},x_{4}\rangle\otimes x_{6}}_{(\mathrm{V})}.$
The centralizer is
$CG(H)\simeq(\mathbb{G}_{m})^{4}\times\mathrm{GL}\!\left(\langle x_{2},x_{3},x_{4}\rangle\right),$
acting blockwise (subject to the determinant-one condition). We normalize the blocks one by one.
Step 1: Normalization of the rank-one blocks $(\mathrm{II})$ and $(\mathrm{V})$.
The blocks
$(\mathrm{II})=\{x_{1}x_{2}x_{5},\,x_{1}x_{3}x_{5},\,x_{1}x_{4}x_{5}\},\qquad(\mathrm{V})=\{x_{0}x_{2}x_{6},\,x_{0}x_{3}x_{6},\,x_{0}x_{4}x_{6}\}$
correspond to two linear functionals on
$U^{\vee}=\langle x_{2},x_{3},x_{4}\rangle^{\vee}.$
For a general member, these two $3$-vectors are linearly independent; hence there exists a single $g\in\mathrm{GL}(U)$ sending them to the coordinate covectors:
$(\mathrm{II})\;\sim\;x_{1}x_{2}x_{5},\qquad(\mathrm{V})\;\sim\;x_{0}x_{3}x_{6}.$
Using the diagonal tori on $(x_{1},x_{5})$ and $(x_{0},x_{6})$ (and projective scaling), we normalize the coefficients to $1$.
Step 2: Normalization of the one-dimensional blocks $(\mathrm{III})$ and $(\mathrm{IV})$.
The $\mathbb{G}_{m}$-factors on $x_{0}$ and $x_{1}$ allow us to set
$(\mathrm{III})\sim x_{0}x_{5}^{2},\qquad(\mathrm{IV})\sim x_{1}^{2}x_{6}.$
Step 3: Normalization of the ternary cubic in $(\mathrm{I})$.
Let
$F(x_{2},x_{3},x_{4})\in\mathrm{Sym}^{3}\langle x_{2},x_{3},x_{4}\rangle.$
After Step 1, the residual subgroup of $\mathrm{GL}(\langle x_{2},x_{3},x_{4}\rangle)$ preserving $(\mathrm{II})$ and $(\mathrm{V})$ consists of
$x_{2}\mapsto s\,x_{2},\qquad x_{3}\mapsto t\,x_{3},\qquad x_{4}\mapsto u\,x_{4}+\alpha x_{2}+\beta x_{3},\qquad(s,t,u)\in(\mathbb{G}_{m})^{3},\;(\alpha,\beta)\in\mathbb{A}^{2}.$
(a) Elimination of $x_{2}x_{4}^{2}$ and $x_{3}x_{4}^{2}$. The substitution $x_{4}\mapsto x_{4}+\alpha x_{2}+\beta x_{3}$ yields
$(x_{4}+\alpha x_{2}+\beta x_{3})^{3}=x_{4}^{3}+3\alpha\,x_{2}x_{4}^{2}+3\beta\,x_{3}x_{4}^{2}+\cdots,$
so choosing $(\alpha,\beta)$ suitably (with the $x_{4}^{3}$ coefficient nonzero), we kill both coefficients.
(b) Normalization of pure cubes. Scaling by $(s,t,u)$ (and compensating using the diagonal tori on $(x_{1},x_{5})$ and $(x_{0},x_{6})$ so that the coefficients fixed in Steps 1–2 remain equal to $1$), we may assume
$F=x_{2}^{3}+x_{3}^{3}+\sigma x_{4}^{3}+\cdots,\qquad\sigma\in\mathbb{C}^{\times}.$
(c) Remaining mixed terms. After (a) and (b), the possible mixed terms are
$x_{2}^{2}x_{3},\;x_{2}x_{3}^{2},\;x_{2}^{2}x_{4},\;x_{3}^{2}x_{4},\;x_{2}x_{3}x_{4}.$
Note that the residual subgroup preserving the normalizations in Step 1 does not contain the shear $x_{3}\mapsto x_{3}+\gamma x_{2}$ (it would reintroduce an $x_{0}x_{2}x_{6}$ term in the $(\mathrm{V})$-block). Hence the coefficient of $x_{2}x_{3}^{2}$ cannot be eliminated in general. Therefore we write
$F(x_{2},x_{3},x_{4})=x_{2}^{3}+x_{3}^{3}+\sigma x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+\kappa\,x_{2}^{2}x_{3}+\nu\,x_{2}x_{3}^{2}+\mu\,x_{3}^{2}x_{4}+\lambda\,x_{2}^{2}x_{4},$
with $\rho,\kappa,\nu,\mu,\lambda\in\mathbb{C}^{\times}$.
Collecting all blocks.
Hence a closed-orbit representative is
$\phi^{\mathrm{nf}}_{16}$ $=x_{2}^{3}+x_{3}^{3}+\sigma x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+\kappa\,x_{2}^{2}x_{3}+\nu\,x_{2}x_{3}^{2}+\mu\,x_{3}^{2}x_{4}+\lambda\,x_{2}^{2}x_{4}$
$\qquad+x_{1}x_{2}x_{5}+x_{0}x_{3}x_{6}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6},$
with six free parameters
$(\sigma,\rho,\kappa,\nu,\mu,\lambda)\in(\mathbb{C}^{\times})^{6}.$
Component dimension.
We compute
$\dim W^{H}=10+3+1+1+3=18.$
Since $C_{G}(H)$ has dimension $12$ and the $H\simeq\mathbb{G}_{m}$ factor acts trivially on $W^{H}$, the effective dimension is
$\dim_{\mathrm{eff}}C_{G}(H)=11.$
Projectivization subtracts one more dimension, giving
$\dim\Phi_{16}=18-11-1=6,$
agreeing with the six independent parameters of $\phi^{\mathrm{nf}}_{16}$.
4.17 Case $k=17$
1-PS limit.
Set
$\lambda_{17}(t)=\mathrm{diag}\!\bigl(t,\,t,\,t,\,1,\,1,\,t^{-1},\,t^{-2}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{17}$ as in Section 3, the $1$-PS limit is
$\phi_{17}=\lim_{t\to 0}\lambda_{17}(t)\cdot f_{17}$ $=a_{1}\,x_{3}^{3}+a_{2}\,x_{3}^{2}x_{4}+a_{3}\,x_{3}x_{4}^{2}+a_{4}\,x_{4}^{3}$
$\quad+a_{5}\,x_{0}x_{3}x_{5}+a_{6}\,x_{1}x_{3}x_{5}+a_{7}\,x_{2}x_{3}x_{5}$
$\quad+a_{8}\,x_{0}x_{4}x_{5}+a_{9}\,x_{1}x_{4}x_{5}+a_{10}\,x_{2}x_{4}x_{5}$
$\quad+a_{11}\,x_{0}^{2}x_{6}+a_{12}\,x_{0}x_{1}x_{6}+a_{13}\,x_{1}^{2}x_{6}$
$\quad+a_{14}\,x_{0}x_{2}x_{6}+a_{15}\,x_{1}x_{2}x_{6}+a_{16}\,x_{2}^{2}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{17}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (1,1,1,0,0,1,2)(1,1,1,0,0,-1,-2) with block multiplicities $\langle x_{0},x_{1},x_{2}\rangle$, $\langle x_{3},x_{4}\rangle$, $\langle x_{5}\rangle$, $\langle x_{6}\rangle$. Thus
$C_{G}(H)$ $=\Bigl\{\,A\oplus B\oplus\mathrm{diag}(\gamma)\oplus\mathrm{diag}(\delta):\ A\in\mathrm{GL}(3),\ B\in\mathrm{GL}(2),\ \gamma,\delta\in\mathbb{G}_{m},$
$\hskip 72.00008pt\det(A)\det(B)\gamma\delta=1\,\Bigr\}$
$\cong\bigl(\mathrm{GL}(3)\times\mathrm{GL}(2)\times\mathbb{G}_{m}\times\mathbb{G}_{m}\bigr)\cap\mathrm{SL}(7).$
Each monomial of $\phi_{17}$ has $H$-weight 0. Moreover, $\dim C_{G}(H)=14$.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{17}$ is equivalent to polystability for the $C_{G}(H)$-action on the $H$-fixed subspace. After conjugating inside the $\mathrm{GL}(3)$- and $\mathrm{GL}(2)$-blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\dots,x_{6})=(a+u_{1},\,a+u_{2},\,a+u_{3},\,b+v,\,b-v,\,c,\,d),\quad u_{1}+u_{2}+u_{3}=0,$
and the $\mathrm{SL}(7)$-condition (Convention 4.6)
$S:=3a+2b+c+d=0.$
A direct calculation yields the positive linear identity
$\bigl[w(x_{3}^{3})+w(x_{3}^{2}x_{4})+w(x_{3}x_{4}^{2})+w(x_{4}^{3})\bigr]$
$\quad+2\bigl[w(x_{0}x_{3}x_{5})+w(x_{1}x_{3}x_{5})+w(x_{2}x_{3}x_{5})+w(x_{0}x_{4}x_{5})+w(x_{1}x_{4}x_{5})+w(x_{2}x_{4}x_{5})\bigr]$
$\quad+2\bigl[w(x_{0}^{2}x_{6})+w(x_{0}x_{1}x_{6})+w(x_{1}^{2}x_{6})+w(x_{0}x_{2}x_{6})+w(x_{1}x_{2}x_{6})+w(x_{2}^{2}x_{6})\bigr]$
$\qquad=12\,S.$ (22)
If $\lambda\in\Lambda_{\phi_{17}}$, then all $16$ monomial weights are $\geq 0$ and $S=0$; by (22) they must all vanish. Solving gives
$b=0,\quad v=0,\quad u_{1}=u_{2}=u_{3}=0,\quad c=-a,\quad d=-2a.$
Writing $a=k$ ($k\in\mathbb{Z}$) yields
$\Lambda_{\phi_{17}}=\{\mu_{k}\mid k\in\mathbb{Z}\},\qquad\mu_{k}(t)=\mathrm{diag}\!\bigl(t^{k},\,t^{k},\,t^{k},\,1,\,1,\,t^{-k},\,t^{-2k}\bigr).$
Thus, $\Lambda_{\phi_{17}}$ is symmetric, and by the Casimiro–Florentino criterion, $\phi_{17}$ is polystable; in particular, $\mathrm{SL}(7)\cdot\phi_{17}$ is closed.
Normal form and component dimension.
On the $H$-fixed slice
$W^{H}=\underbrace{\mathrm{Sym}^{3}\langle x_{3},x_{4}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{\langle x_{0},x_{1},x_{2}\rangle\otimes\langle x_{3},x_{4}\rangle\otimes x_{5}}_{(\mathrm{II})}\ \oplus\ \underbrace{\mathrm{Sym}^{2}\langle x_{0},x_{1},x_{2}\rangle\otimes x_{6}}_{(\mathrm{III})},$
of total dimension $4+6+6=16$.
$(\mathrm{III})$ The quadratic block $\mathrm{Sym}^{2}\langle x_{0},x_{1},x_{2}\rangle\otimes x_{6}$.
By the $\mathrm{GL}_{3}$-action, one diagonalizes the ternary quadratic, and using diagonal tori together with projective rescaling, the coefficients of $x_{0}^{2}x_{6}$ and $x_{1}^{2}x_{6}$ are normalized to $1$, leaving
$x_{0}^{2}x_{6}\;+\;x_{1}^{2}x_{6}\;+\;\rho\,x_{2}^{2}x_{6},\qquad\rho\in\mathbb{C}^{\times}.$
Importantly, after the reduction of $(\mathrm{II})$ below, the subgroup of $\mathrm{GL}_{3}$ that preserves $(\mathrm{II})$ is block-diagonal on $\langle x_{0},x_{1}\rangle\oplus\langle x_{2}\rangle$ (up to an $\mathrm{O}(2)$ on $\langle x_{0},x_{1}\rangle$ and an independent scaling on $x_{2}$). Hence, the relative scale of $x_{2}$ cannot be absorbed, and the parameter $\rho$ cannot be removed.
$(\mathrm{II})$ The $3\times 2$ block $\langle x_{0},x_{1},x_{2}\rangle\otimes\langle x_{3},x_{4}\rangle\otimes x_{5}$.
Using the left $\mathrm{GL}_{3}$ and right $\mathrm{GL}_{2}$ actions (an SVD-type reduction), a generic rank-$2$ element is brought to diagonal form; scaling $x_{5}$ then fixes the two nonzero entries to $1$:
$x_{0}x_{3}x_{5}\;+\;x_{1}x_{4}x_{5}.$
$(\mathrm{I})$ The binary cubic $\mathrm{Sym}^{3}\langle x_{3},x_{4}\rangle$.
Acting by $\mathrm{GL}_{2}$ on $\langle x_{3},x_{4}\rangle$ puts a general binary cubic into a two-term form (sending three roots to $\{0,\infty,-\tau\}$), and a residual diagonal scaling fixes the first coefficient to $1$:
$x_{3}^{2}x_{4}\;+\;\tau\,x_{3}x_{4}^{2},\qquad\tau\in\mathbb{C}^{\times}.$
Collecting $(\mathrm{I})$-$(\mathrm{III})$, we obtain the closed-orbit representative
$\phi^{\mathrm{nf}}_{17}(\tau,\rho)=x_{3}^{2}x_{4}+\tau\,x_{3}x_{4}^{2}+x_{0}x_{3}x_{5}+x_{1}x_{4}x_{5}+x_{0}^{2}x_{6}+x_{1}^{2}x_{6}+\rho\,x_{2}^{2}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$
Lastly, $\dim W^{H}=16$ and $\dim C_{G}(H)=14$. As the $H\simeq\mathbb{G}_{m}$-factor acts trivially on $W^{H}$, the effective group dimension is 141=1314-1=13. After projectivizing, we get
$\dim(\text{component})=\dim W^{H}-13-1=2,$
so $(\tau,\rho)$ are the two genuine moduli of the closed stratum. Hence, the corresponding component $\Phi_{17}$ of the moduli is two-dimensional.
4.18 Case $k=18$
1-PS limit.
Set
$\lambda_{18}(t)=\mathrm{diag}\!\bigl(t,\,t,\,1,\,1,\,1,\,t^{-1},\,t^{-1}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{18}$ as in Section 3, the $1$-PS limit is
$\phi_{18}=\lim_{t\to 0}\lambda_{18}(t)\cdot f_{18}$ $=a_{1}\,x_{2}^{3}+a_{2}\,x_{2}^{2}x_{3}+a_{3}\,x_{2}x_{3}^{2}+a_{4}\,x_{3}^{3}+a_{5}\,x_{2}^{2}x_{4}+a_{6}\,x_{2}x_{3}x_{4}+a_{7}\,x_{3}^{2}x_{4}$
$\quad+a_{8}\,x_{2}x_{4}^{2}+a_{9}\,x_{3}x_{4}^{2}+a_{10}\,x_{4}^{3}+a_{11}\,x_{0}x_{2}x_{5}+a_{12}\,x_{1}x_{2}x_{5}+a_{13}\,x_{0}x_{3}x_{5}$
$\quad+a_{14}\,x_{1}x_{3}x_{5}+a_{15}\,x_{0}x_{4}x_{5}+a_{16}\,x_{1}x_{4}x_{5}+a_{17}\,x_{0}x_{2}x_{6}+a_{18}\,x_{1}x_{2}x_{6}$
$\quad+a_{19}\,x_{0}x_{3}x_{6}+a_{20}\,x_{1}x_{3}x_{6}+a_{21}\,x_{0}x_{4}x_{6}+a_{22}\,x_{1}x_{4}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{18}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (1,1,0,0,0,1,1)(1,1,0,0,0,-1,-1) with block decomposition
$\langle x_{0},x_{1}\rangle\oplus\langle x_{2},x_{3},x_{4}\rangle\oplus\langle x_{5},x_{6}\rangle.$
Hence
$C_{G}(H)$ $=\Bigl\{\,A\oplus B\oplus C:\ A\in\mathrm{GL}(2),\ B\in\mathrm{GL}(3),\ C\in\mathrm{GL}(2),$
$\hskip 75.0001pt\det(A)\det(B)\det(C)=1\,\Bigr\}$
$\cong\bigl(\mathrm{GL}(2)\times\mathrm{GL}(3)\times\mathrm{GL}(2)\bigr)\cap\mathrm{SL}(7),$
so $\dim C_{G}(H)=16$. Each monomial of $\phi_{18}$ has $H$-weight 0; hence $\phi_{18}\in W^{H}$.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{18}$ is equivalent to polystability for the $C_{G}(H)$-action on the $H$-fixed subspace. After conjugating within the three blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\dots,x_{6})=(a+u,\,a-u,\,b+v_{1},\,b+v_{2},\,b+v_{3},\,c+w,\,c-w),$
where $a,b,c,u,w,v_{i}\in\mathbb{Z}$ and $v_{1}+v_{2}+v_{3}=0$, and with the $\mathrm{SL}$-constraint
$S:=2a+3b+2c=0.$
Let $w(\cdot)$ denote the $\lambda$-weight of a monomial. Then
$\sum_{\text{all 10 cubics in }x_{2},x_{3},x_{4}}w\;=\;30\,b,$
and
$\sum_{i=2}^{4}[w(x_{0}x_{i}x_{5})+w(x_{1}x_{i}x_{5})]=6(a+b+c+w),\qquad\sum_{i=2}^{4}[w(x_{0}x_{i}x_{6})+w(x_{1}x_{i}x_{6})]=6(a+b+c-w).$
Hence the positive identity
$\Bigl[\sum_{\text{all 10 cubics in }x_{2},x_{3},x_{4}}w\Bigr]+5\sum_{i=2}^{4}\bigl[w(x_{0}x_{i}x_{5})+w(x_{1}x_{i}x_{5})\bigr]$
$\quad+5\sum_{i=2}^{4}\bigl[w(x_{0}x_{i}x_{6})+w(x_{1}x_{i}x_{6})\bigr]\;=\;30\,S.$ (23)
If $\lambda\in\Lambda_{\phi_{18}}$, then all $22$ monomial weights are $\geq 0$ and $S=0$; hence, by (23) they must all vanish. Solving yields
$b=0,\quad v_{1}=v_{2}=v_{3}=0,\quad u=0,\quad w=0,\quad a+c=0.$
Writing $a=k$ ($k\in\mathbb{Z}$) gives
$\Lambda_{\phi_{18}}=\{\mu_{k}\mid k\in\mathbb{Z}\},\qquad\mu_{k}(t)=\mathrm{diag}\!\bigl(t^{k},\,t^{k},\,1,\,1,\,1,\,t^{-k},\,t^{-k}\bigr).$
Thus $\Lambda_{\phi_{18}}$ is symmetric, and by the Casimiro–Florentino criterion, $\phi_{18}$ is polystable.
Normal form and component dimension.
Put
$V:=\langle x_{0},x_{1}\rangle,\qquad U:=\langle x_{2},x_{3},x_{4}\rangle,\qquad W:=\langle x_{5},x_{6}\rangle.$
Then the HH–fixed slice decomposes as
$W^{H}=\mathrm{Sym}^{3}U\;\oplus\;(V\otimes U\otimes x_{5})\;\oplus\;(V\otimes U\otimes x_{6})\;\cong\;\mathrm{Sym}^{3}U\;\oplus\;(V\otimes U\otimes W),$
and $C_{G}(H)=\{A\oplus B\oplus C:\ A\in GL(V),\ B\in GL(U),\ C\in GL(W),\ \det(A)\det(B)\det(C)=1\}$ acts blockwise (with the same $B\in GL(U)$ acting simultaneously on the cubic block and on the mixed block).
To avoid any ambiguity about this shared GL(U)GL(U)–action, we normalize in two stages: first we fix a convenient basis of $U$ using the cubic block, and then we reduce the mixed block using only those UU–changes that preserve the chosen cubic shape (namely permutations and diagonal scalings of $x_{2},x_{3},x_{4}$).
(I) The ternary cubic block (fixing the UU–basis). Write the cubic part as $F\in\mathrm{Sym}^{3}U$. On the dense open subset where $F$ is a smooth plane cubic with nonvanishing Hessian, the classical Hesse normal form gives a change of basis $B\in GL(U)$ and a scalar such that (after rescaling and renaming the coordinates on $U$) we may assume
$F\;\sim\;x_{2}^{3}+\sigma_{1}x_{3}^{3}+\sigma_{2}x_{4}^{3}+\rho\,x_{2}x_{3}x_{4},\qquad(\sigma_{1},\sigma_{2},\rho)\in(\mathbb{C}^{\times})^{3}.$ (24)
From this point on, the basis $(x_{2},x_{3},x_{4})$ is fixed. The remaining GL(U)GL(U)–freedom that preserves the 44–term shape in (24) is the monomial subgroup: permutations of $(x_{2},x_{3},x_{4})$ and diagonal scalings $x_{i}\mapsto t_{i}x_{i}$ (which only rescale $(\sigma_{1},\sigma_{2},\rho)$).
(II)+(III) The mixed block as a $2\times 3$ matrix pencil (reduced relative to the fixed UU–basis). Write the mixed part as
$x_{5}\,L_{5}\;+\;x_{6}\,L_{6},\qquad L_{5},L_{6}\in V\otimes U,$
and represent $L_{5},L_{6}$ by $2\times 3$ matrices $M_{5},M_{6}$ with respect to the fixed bases $(x_{0},x_{1})$ of $V$ and $(x_{2},x_{3},x_{4})$ of $U$:
$L_{j}=(x_{0},x_{1})\,M_{j}\,(x_{2},x_{3},x_{4})^{\mathsf{T}}\qquad(j=5,6).$
The GL(V)GL(V)–factor acts by left multiplication on both $M_{5},M_{6}$, the GL(W)GL(W)–factor changes the basis of the pencil (so it replaces $(M_{5},M_{6})$ by another spanning pair), and the residual monomial subgroup in $GL(U)$ permutes/rescales the three columns.
We work on the dense open set where the pencil $\langle M_{5},M_{6}\rangle$ contains a rank-$2$ matrix and where suitable $2\times 2$ minors in the fixed column basis are nonzero (an open condition). Using $GL(W)$ we choose the basis $(x_{5},x_{6})$ so that $M_{5}$ has rank $2$. After possibly permuting $(x_{2},x_{3},x_{4})$ (allowed by the monomial subgroup), we may assume the $2\times 2$ minor of $M_{5}$ in the $(x_{2},x_{3})$–columns is invertible. Then, multiplying on the left by its inverse (an element of $GL(V)$), we achieve
$M_{5}\;\sim\;\begin{pmatrix}1&0&\alpha\\ 0&1&0\end{pmatrix},\qquad\alpha\in\mathbb{C}^{\times},$
where the remaining entry in the $(x_{4})$–column is recorded as the modulus $\alpha$ (we do not use column additions, only the allowed row operations, column permutation/diagonal scaling, and basis change in $W$).
Next, keeping $M_{5}$ fixed, we use the subgroup of $GL(W)$ stabilizing the first basis vector $x_{5}$ (i.e. upper–triangular matrices) to replace $M_{6}$ by $M_{6}+\lambda M_{5}$, and use GL(V)GL(V)–row operations that preserve the pivot normalization of $M_{5}$. On the same open set one can arrange that $M_{6}$ takes the form
$M_{6}\;\sim\;\begin{pmatrix}1&0&0\\ 0&\beta&\gamma\end{pmatrix},\qquad(\beta,\gamma)\in(\mathbb{C}^{\times})^{2},$
after additionally allowing diagonal rescalings of $(x_{2},x_{3},x_{4})$ (which remain inside the 44–term family (24) and merely rescale $(\sigma_{1},\sigma_{2},\rho)$).
Translating back to monomials, the mixed block becomes
$x_{0}x_{2}x_{5}+x_{1}x_{3}x_{5}+\alpha\,x_{0}x_{4}x_{5}\;+\;x_{0}x_{2}x_{6}+\beta\,x_{1}x_{3}x_{6}+\gamma\,x_{1}x_{4}x_{6},\qquad(\alpha,\beta,\gamma)\in(\mathbb{C}^{\times})^{3}.$
Collecting (I)–(III). Combining (24) with the above reduction of the mixed block, we obtain the convenient normal form
$\varphi^{\mathrm{nf}}_{18}=x_{2}^{3}+\sigma_{1}x_{3}^{3}+\sigma_{2}x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+x_{0}x_{2}x_{5}+x_{1}x_{3}x_{5}+\alpha\,x_{0}x_{4}x_{5}+x_{0}x_{2}x_{6}+\beta\,x_{1}x_{3}x_{6}+\gamma\,x_{1}x_{4}x_{6},$
with $(\sigma_{1},\sigma_{2},\rho,\alpha,\beta,\gamma)\in(\mathbb{C}^{\times})^{6}$. The corresponding component $\Phi_{18}$ of the moduli is six-dimensional.
4.19 Case $k=19$
1-PS limit.
Set
$\lambda_{19}(t)=\mathrm{diag}\!\bigl(t^{2},\,t^{2},\,t^{2},\,1,\,t^{-1},\,t^{-1},\,t^{-4}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{19}$ as in Section 3, the $1$-PS limit is
$\phi_{19}=\lim_{t\to 0}\lambda_{19}(t)\cdot f_{19}$ $=a_{1}\,x_{3}^{3}+a_{2}\,x_{0}x_{4}^{2}+a_{3}\,x_{1}x_{4}^{2}+a_{4}\,x_{2}x_{4}^{2}$
$\quad+a_{5}\,x_{0}x_{4}x_{5}+a_{6}\,x_{1}x_{4}x_{5}+a_{7}\,x_{2}x_{4}x_{5}$
$\quad+a_{8}\,x_{0}x_{5}^{2}+a_{9}\,x_{1}x_{5}^{2}+a_{10}\,x_{2}x_{5}^{2}$
$\quad+a_{11}\,x_{0}^{2}x_{6}+a_{12}\,x_{0}x_{1}x_{6}+a_{13}\,x_{1}^{2}x_{6}$
$\quad+a_{14}\,x_{0}x_{2}x_{6}+a_{15}\,x_{1}x_{2}x_{6}+a_{16}\,x_{2}^{2}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{19}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (2,2,2,0,1,1,4)(2,2,2,0,-1,-1,-4) with block decomposition
$\langle x_{0},x_{1},x_{2}\rangle\ \oplus\ \langle x_{3}\rangle\ \oplus\ \langle x_{4},x_{5}\rangle\ \oplus\ \langle x_{6}\rangle.$
Hence,
$C_{G}(H)$ $=\Bigl\{\,A\oplus\mathrm{diag}(\beta)\oplus B\oplus\mathrm{diag}(\delta):\ A\in\mathrm{GL}(3),\ B\in\mathrm{GL}(2),$
$\hskip 98.00017pt\beta,\delta\in\mathbb{G}_{m},\ \det(A)\,\beta\,\det(B)\,\delta=1\,\Bigr\}$
$\cong\bigl(\mathrm{GL}(3)\times\mathrm{GL}(2)\times\mathbb{G}_{m}\times\mathbb{G}_{m}\bigr)\cap\mathrm{SL}(7),$
so $\dim C_{G}(H)=14$. Each monomial of $\phi_{19}$ has $H$-weight 0.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{19}$ is equivalent to polystability for the $C_{G}(H)$-action on the $H$-fixed subspace. After conjugating inside the blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\dots,x_{6})=(s+u_{1},\,s+u_{2},\,s+u_{3},\,b,\,t+v,\,t-v,\,c),$
where $s,t,b,c,u_{i},v\in\mathbb{Z}$, $u_{1}+u_{2}+u_{3}=0$, and the $\mathrm{SL}(7)$-condition (Convention 4.6) is
$S:=3s+b+2t+c=0.$
A direct computation yields the positive identity
$\sum_{i=0}^{2}\Bigl[w\bigl(x_{i}x_{4}^{2}\bigr)+2\,w\bigl(x_{i}x_{4}x_{5}\bigr)+w\bigl(x_{i}x_{5}^{2}\bigr)\Bigr]$
$\quad+2\Bigl[w(x_{0}^{2}x_{6})+w(x_{0}x_{1}x_{6})+w(x_{1}^{2}x_{6})+w(x_{0}x_{2}x_{6})+w(x_{1}x_{2}x_{6})+w(x_{2}^{2}x_{6})\Bigr]$
$\quad+4\,w(x_{3}^{3})\;=\;12\,S.$ (25)
If $\lambda\in\Lambda_{\phi_{19}}$, then all $16$ monomial weights are $\geq 0$ and $S=0$; by (25) they must all vanish. Solving gives
$b=0,\quad v=0,\quad s+u_{r}+2t=0\ (r=1,2,3),\quad 2s+c=0.$
Using $u_{1}+u_{2}+u_{3}=0$ we obtain $s+2t=0$ and $u_{1}=u_{2}=u_{3}=0$. Writing $s=2k$ with $k\in\mathbb{Z}$ yields
$\Lambda_{\phi_{19}}=\{\mu_{k}\mid k\in\mathbb{Z}\},\qquad\mu_{k}(t)=\mathrm{diag}\!\bigl(t^{2k},\,t^{2k},\,t^{2k},\,1,\,t^{-k},\,t^{-k},\,t^{-4k}\bigr).$
Thus $\Lambda_{\phi_{19}}$ is symmetric, and by the Casimiro–Florentino criterion $\phi_{19}$ is polystable; in particular $\mathrm{SL}(7)\cdot\phi_{19}$ is closed.
Normal form and component dimension.
In the $H$-fixed subspace
$W^{H}=\ \langle x_{3}^{3}\rangle\ \oplus\ \langle x_{0},x_{1},x_{2}\rangle\otimes\mathrm{Sym}^{2}\langle x_{4},x_{5}\rangle\ \oplus\ \mathrm{Sym}^{2}\langle x_{0},x_{1},x_{2}\rangle\otimes x_{6},$
using the $\mathrm{GL}(3)$-action on $\langle x_{0},x_{1},x_{2}\rangle$ to identify the basis of $\langle x_{0},x_{1},x_{2}\rangle$ with the standard basis $\{x_{4}^{2},x_{4}x_{5},x_{5}^{2}\}$ of $\mathrm{Sym}^{2}\langle x_{4},x_{5}\rangle$ (ensuring full rank), and using the residual stabilizer (isomorphic to $\mathrm{GL}(2)$) to normalize the third block, a generic element is brought to
$\phi_{19}^{\mathrm{nf}}(\rho,\sigma)=x_{3}^{3}+(x_{0}x_{4}^{2}+x_{1}x_{4}x_{5}+x_{2}x_{5}^{2})+x_{6}(x_{0}^{2}+x_{2}^{2}+\rho\,x_{1}^{2}+\sigma\,x_{0}x_{2}),$
with $(\rho,\sigma)\in{\mathbb{C}^{\times}}^{2}$ general. As $\dim W^{H}=16$ and $\dim C_{G}(H)=14$ (with $H$ acting trivially), the effective action has dimension $13$; after projectivizing we obtain
$\dim\Phi_{19}=16-13-1=2.$
The residual $T$-stabilizer is finite; hence, the corresponding component $\Phi_{19}$ of the moduli is two-dimensional.
4.20 Case $k=20$
1-PS limit.
Set
$\lambda_{20}(t)=\mathrm{diag}\!\bigl(t,\,t,\,t,\,t,\,1,\,t^{-2},\,t^{-2}\bigr),\quad t\in\mathbb{G}_{m}.$
For a generic $f_{20}$ as in Section 3, the $1$-PS limit is
$\phi_{20}=\lim_{t\to 0}\lambda_{20}(t)\cdot f_{20}$ $=a_{1}\,x_{4}^{3}+a_{2}\,x_{0}^{2}x_{5}+a_{3}\,x_{0}x_{1}x_{5}+a_{4}\,x_{1}^{2}x_{5}+a_{5}\,x_{0}x_{2}x_{5}+a_{6}\,x_{1}x_{2}x_{5}$
$\quad+a_{7}\,x_{2}^{2}x_{5}+a_{8}\,x_{0}x_{3}x_{5}+a_{9}\,x_{1}x_{3}x_{5}+a_{10}\,x_{2}x_{3}x_{5}+a_{11}\,x_{3}^{2}x_{5}$
$\quad+a_{12}\,x_{0}^{2}x_{6}+a_{13}\,x_{0}x_{1}x_{6}+a_{14}\,x_{1}^{2}x_{6}+a_{15}\,x_{0}x_{2}x_{6}+a_{16}\,x_{1}x_{2}x_{6}$
$\quad+a_{17}\,x_{2}^{2}x_{6}+a_{18}\,x_{0}x_{3}x_{6}+a_{19}\,x_{1}x_{3}x_{6}+a_{20}\,x_{2}x_{3}x_{6}+a_{21}\,x_{3}^{2}x_{6}.$
$H$ and $C_{G}(H)$.
Let $H=\lambda_{20}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (1,1,1,1,0,2,2)(1,1,1,1,0,-2,-2) with block multiplicities $\langle x_{0},x_{1},x_{2},x_{3}\rangle$, $\langle x_{4}\rangle$, $\langle x_{5},x_{6}\rangle$. Hence,
$C_{G}(H)$ $=\Bigl\{A\oplus\mathrm{diag}(\beta)\oplus C:\ A\in\mathrm{GL}(4),\ \beta\in\mathbb{G}_{m},\ C\in\mathrm{GL}(2),$
$\hskip 134.00014pt\det(A)\,\beta\,\det(C)=1\Bigr\}$
$\cong\bigl(\mathrm{GL}(4)\times\mathbb{G}_{m}\times\mathrm{GL}(2)\bigr)\cap\mathrm{SL}(7),$
so $\dim C_{G}(H)=20$. Every monomial of $\phi_{20}$ has $H$-weight 0.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{20}$ is equivalent to polystability for the $C_{G}(H)$-action on the $H$-fixed subspace. After conjugating inside the $\mathrm{GL}(4)$- and $\mathrm{GL}(2)$-blocks, any $\lambda\in Y(C_{G}(H))$ may be taken with weights
$\mathrm{wt}(x_{0},\dots,x_{6})=\bigl(a+u_{0},\,a+u_{1},\,a+u_{2},\,a+u_{3},\,b,\,c+w,\,c-w\bigr),$
where $a,b,c,u_{i},w\in\mathbb{Z}$, $u_{0}+u_{1}+u_{2}+u_{3}=0$, and, by Convention 4.6,
$S:=4a+b+2c=0.$
Let $w(\cdot)$ denote the $\lambda$-weight of a monomial. Then
$w(x_{4}^{3})=3b,$
$w(x_{i}^{2}x_{5})=2(a+u_{i})+c+w,\qquad w(x_{i}^{2}x_{6})=2(a+u_{i})+c-w\quad(i=0,1,2,3),$
$w(x_{i}x_{j}x_{5})=2a+(u_{i}+u_{j})+c+w,\quad w(x_{i}x_{j}x_{6})=2a+(u_{i}+u_{j})+c-w\ (0\leq i<j\leq 3).$
A direct computation yields the positive identity
$3\sum_{i=0}^{3}\bigl[w(x_{i}^{2}x_{5})+w(x_{i}^{2}x_{6})\bigr]+3\sum_{0\leq i<j\leq 3}\bigl[w(x_{i}x_{j}x_{5})+w(x_{i}x_{j}x_{6})\bigr]$
$\qquad\qquad\qquad\qquad\qquad\qquad+10\,w(x_{4}^{3})\;=\;30\,S.$ (26)
If $\lambda\in\Lambda_{\phi_{20}}$, then all $21$ monomial weights are $\geq 0$ and $S=0$; by (26) they must all vanish. Solving gives
$b=0,\quad w=0,\quad 2(a+u_{i})+c=0\ (i=0,1,2,3),\quad 2a+(u_{i}+u_{j})+c=0\ (i<j),$
Hence, $u_{0}=u_{1}=u_{2}=u_{3}=0$ and $2a+c=0$. Writing $a=k\in\mathbb{Z}$ yields
$\Lambda_{\phi_{20}}=\{\mu_{k}\mid k\in\mathbb{Z}\},\qquad\mu_{k}(t)=\mathrm{diag}\!\bigl(t^{k},\,t^{k},\,t^{k},\,t^{k},\,1,\,t^{-2k},\,t^{-2k}\bigr).$
As $\Lambda_{\phi_{20}}$ is symmetric, $\phi_{20}$ is polystable by the Casimiro–Florentino criterion; in particular, $\mathrm{SL}(7)\cdot\phi_{20}$ is closed.
Normal form and component dimension.
We have
$W^{H}=\underbrace{\langle x_{4}^{3}\rangle}_{(\mathrm{I})}\ \oplus\ \underbrace{\mathrm{Sym}^{2}\langle x_{0},x_{1},x_{2},x_{3}\rangle\otimes\langle x_{5}\rangle}_{(\mathrm{II})}\ \oplus\ \underbrace{\mathrm{Sym}^{2}\langle x_{0},x_{1},x_{2},x_{3}\rangle\otimes\langle x_{6}\rangle}_{(\mathrm{III})},$
so a general element of $W^{H}$ is
$\varphi=a\,x_{4}^{3}+x_{5}Q_{5}(x_{0},\dots,x_{3})+x_{6}Q_{6}(x_{0},\dots,x_{3}),\qquad Q_{5},Q_{6}\in\mathrm{Sym}^{2}\langle x_{0},\dots,x_{3}\rangle^{\vee}.$
Using the central torus factor together with projective scaling, we normalize $a=1$.
Diagonalization of the quadratic pencil. Consider the pencil $\langle Q_{5},Q_{6}\rangle$ and its discriminant
$\Delta(s,t):=\det(sQ_{5}+tQ_{6}),\qquad[s:t]\in\mathbb{P}^{1}.$
On the dense open set where $\Delta$ has four distinct roots (equivalently, the pencil is regular, i.e. it lies off the discriminant locus), we may assume $Q_{5}$ is nondegenerate. Acting by $\mathrm{GL}\langle x_{0},x_{1},x_{2},x_{3}\rangle\cong\mathrm{GL}(4)$, we take
$Q_{5}\sim q:=x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}.$
With respect to $q$, the residual group on $\langle x_{0},\dots,x_{3}\rangle$ is $O(4)$. For a regular pencil, the $q$-selfadjoint endomorphism $A$ defined by $Q_{6}(u,v)=q(Au,v)$ has four distinct eigenvalues, hence after an $O(4)$-change of coordinates
$Q_{6}\sim\lambda_{0}x_{0}^{2}+\lambda_{1}x_{1}^{2}+\lambda_{2}x_{2}^{2}+\lambda_{3}x_{3}^{2},\qquad\lambda_{i}\ \text{pairwise distinct}.$
Reduction to a one-parameter normal form (cross-ratio). The $\mathrm{GL}\langle x_{5},x_{6}\rangle\cong\mathrm{GL}(2)$-action replaces $(Q_{5},Q_{6})$ by
$(Q_{5}^{\prime},Q_{6}^{\prime})=(\alpha Q_{5}+\beta Q_{6},\ \gamma Q_{5}+\delta Q_{6}),\qquad\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}\in\mathrm{GL}(2),$
and, after re-normalizing $Q_{5}^{\prime}$ back to $q$, this acts on the eigenvalues by the fractional-linear transformation
$\lambda\ \longmapsto\ \frac{\gamma+\delta\lambda}{\alpha+\beta\lambda}.$
Thus a regular pencil is determined (up to permuting the diagonal coordinates) by the unordered set $\{\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3}\}\subset\mathbb{P}^{1}$ modulo $\mathrm{PGL}_{2}$, i.e. by one cross-ratio parameter. Using $3$-transitivity of $\mathrm{PGL}_{2}$ on $\mathbb{P}^{1}$, we may normalize three eigenvalues to 1,1,21,-1,2 and denote the fourth by $\tau$. After permuting $(x_{0},x_{1},x_{2},x_{3})$ if necessary, we assume
$(\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3})=(1,\tau,-1,2),\qquad\tau\notin\{1,-1,2\}.$
Consequently, on the open set $\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}$, a convenient normal form is
$\varphi^{\mathrm{nf}}_{20}(\tau)=x_{4}^{3}+x_{5}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})+x_{6}(x_{0}^{2}+\tau x_{1}^{2}-x_{2}^{2}+2x_{3}^{2}),\qquad\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}.$
(For the excluded values, the pencil acquires repeated eigenvalues and the stabilizer jumps; these form a proper closed subset.)
Component dimension. Here $\dim W^{H}=21$ and $\dim C_{G}(H)=20$, and $C_{G}(H)$ contains a one-dimensional central torus acting trivially on $W^{H}$. Hence the effective action has dimension $19$, and after projectivizing we obtain
$\dim(\Phi_{20})=\dim W^{H}-19-1=1,$
in agreement with the single modulus $\tau$ for a regular pencil.
4.21 Case $k=21$
1-PS limit.
Set
$\lambda_{21}(t)=\operatorname{diag}\!\bigl(t,\,t,\,1,\,1,\,1,\,1,\,t^{-2}\bigr),\qquad t\in\mathbb{G}_{m}.$
For a generic $f_{21}$ as in Section 3, the $1$-PS limit is
$\phi_{21}:=\lim_{t\to 0}\lambda_{21}(t)\cdot f_{21}=F(x_{2},x_{3},x_{4},x_{5})+x_{6}\,Q(x_{0},x_{1}),$
where $F\in\mathrm{Sym}^{3}\langle x_{2},x_{3},x_{4},x_{5}\rangle$ and $Q\in\mathrm{Sym}^{2}\langle x_{0},x_{1}\rangle$.
$H$ and $C_{G}(H)$.
Let $H=\lambda_{21}(\mathbb{G}_{m})$. The $H$-weights on $\langle x_{0},\dots,x_{6}\rangle$ are (1,1,0,0,0,0,2)(1,1,0,0,0,0,-2) with block decomposition $\langle x_{0},x_{1}\rangle\oplus\langle x_{2},x_{3},x_{4},x_{5}\rangle\oplus\langle x_{6}\rangle$. Hence,
$C_{G}(H)$ $=\Bigl\{\ A\oplus B\oplus\gamma\ :\ A\in\mathrm{GL}(2),\ B\in\mathrm{GL}(4),\ \gamma\in\mathbb{G}_{m},\ \det(A)\det(B)\gamma=1\ \Bigr\}$
[…]
Every monomial of $\phi_{21}$ has $H$-weight 0, so $\phi_{21}\in W^{H}$.
Polystability (Luna + Casimiro–Florentino).
By Luna’s reduction, the closedness of the $\mathrm{SL}(7)$-orbit of $\phi_{21}$ is equivalent to polystability for the $C_{G}(H)$-action on $W^{H}$. After conjugating inside the $\mathrm{GL}(2)$- and $\mathrm{GL}(4)$-blocks, any $\lambda\in Y\!\bigl(C_{G}(H)\bigr)$ may be taken with
$\mathrm{wt}(x_{0},\ldots,x_{6})=(\alpha+u,\,\alpha-u,\,\beta+v_{1},\,\dots,\,\beta+v_{4},\,\gamma),$
where $\sum v_{i}=0$, and with the $\mathrm{SL}$-constraint $S:=2\alpha+4\beta+\gamma=0$. The weight of any monomial in $F$ is $3\beta$, and the weight of any monomial in $x_{6}Q$ is $\gamma+2\alpha$. Using $S=0$, we have $\gamma+2\alpha=-4\beta$. Thus we satisfy the positive linear identity
$3\,w(x_{6}Q)\;+\;4\,w(F)\;=\;3(-4\beta)+4(3\beta)\;=\;0.$ (27)
If $\lambda\in\Lambda_{\phi_{21}}$, weights are non-negative, so they must vanish: $\beta=0$ (implies $v_{i}=0$) and $\gamma+2\alpha=0$. The condition $w(x_{6}Q)=0$ with $u=0$ (from generic stability of $Q$) implies $\lambda$ reduces to the form of $\lambda_{21}$ (up to scaling). Thus $\Lambda_{\phi_{21}}$ is symmetric, and by the Casimiro–Florentino criterion, $\phi_{21}$ is polystable.
Normal form and component dimension.
We normalize generic elements under the action of the centralizer $C_{G}(H)$.
(1) Quadratic part. Under the $\mathrm{GL}(2)$-action on $\langle x_{0},x_{1}\rangle$, a generic binary quadratic $Q$ is normalized to $x_{0}^{2}+x_{1}^{2}$. The stabilizer of this form contains the orthogonal group $\mathrm{O}(2)$, contributing 1 dimension to the stabilizer of the generic point.
(2) Cubic part. Under the $\mathrm{GL}(4)$-action on $\langle x_{2},x_{3},x_{4},x_{5}\rangle$, a generic cubic surface $F$ can be written as the sum of five cubes of linear forms (Sylvester’s Pentahedron Theorem). Let $L_{0},\dots,L_{4}$ be fixed linear forms defining a generic pentahedron in $\mathbb{C}^{4}$. Then $F=\sum_{i=0}^{4}c_{i}L_{i}^{3}$.
Thus, the normal form is:
$\varphi_{21}^{\mathrm{nf}}(c)=\sum_{i=0}^{4}c_{i}L_{i}(x_{2},x_{3},x_{4},x_{5})^{3}+x_{6}(x_{0}^{2}+x_{1}^{2}),$
where the parameters are $c_{0}:\dots:c_{4}]\in\mathbb{P}^{4}$.
The dimension of the component is computed as follows: $\dim W^{H}=20(\text{cubic})+3(\text{quadratic})=23$. The effective group dimension is $\dim C_{G}(H)-1=19$ (since $H$ acts trivially). Since the generic stabilizer has dimension $1$ (due to $\mathrm{O}(2)$), the orbit dimension is 191=1819-1=18.
$\dim(\Phi_{21})=23-18-1=4.$
Table 2: Summary of closed orbit representatives, criteria used, and component dimensions for $k=1,\dots,21$. C-H means the Convex-hull criterion and C-F means the Casimiro–Florentino criterion.
$k$ Criterion Normal form Parameters Dim
1 C-H $x_{2}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{2}^{2}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{0}x_{4}x_{6}$ none 0
2 C-H $x_{2}^{2}x_{4}+x_{1}x_{4}^{2}+x_{1}x_{3}x_{5}+x_{0}x_{5}^{2}+x_{1}x_{2}x_{6}+x_{0}x_{3}x_{6}$ none 0
3 C-F $x_{1}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{0}x_{5}^{2}+x_{1}x_{2}x_{6}+x_{0}x_{3}x_{6}+x_{0}x_{4}x_{6}$ none 0
4 C-H $x_{3}^{3}+x_{2}x_{3}x_{4}+x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{4}x_{5}+\alpha x_{1}x_{2}x_{6}+\beta x_{0}x_{3}x_{6}$ $\alpha,\beta\in\mathbb{C}^{\times}$ 22
5 C-H $x_{3}^{3}+x_{2}x_{3}x_{4}+x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{5}^{2}+\alpha x_{1}^{2}x_{6}+\beta x_{0}x_{3}x_{6}$ $\alpha,\beta\in\mathbb{C}^{\times}$ 22
6 C-H $x_{2}x_{4}^{2}+x_{2}^{2}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{4}x_{5}+x_{1}^{2}x_{6}+x_{0}x_{3}x_{6}$ none 0
7 C-H $x_{3}^{2}x_{4}+x_{1}x_{4}^{2}+x_{2}x_{3}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{0}x_{2}x_{6}$ none 0
8 C-F $x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{5}^{2}+x_{0}x_{6}^{2}+x_{1}^{2}x_{4}+\rho\,x_{1}x_{2}x_{5}+\sigma\,x_{2}^{2}x_{6}$ $(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}$ 22
9 C-F $x_{2}^{2}x_{3}+\tau\,x_{2}x_{3}^{2}+x_{0}x_{4}^{2}+\rho\,x_{0}x_{5}^{2}+x_{1}x_{4}^{2}+x_{1}x_{5}^{2}+x_{0}x_{2}x_{6}+x_{1}x_{3}x_{6}$ $(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$ 22
10 C-F $x_{2}^{3}+\tau\,x_{2}^{2}x_{3}+\rho\,x_{2}x_{3}^{2}+x_{3}^{3}+x_{1}x_{2}x_{5}+x_{1}x_{3}x_{4}+x_{0}x_{4}x_{5}+x_{0}x_{6}^{2}$ $(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$ 22
11 C-F $\sum_{i=0}^{4}c_{i}L_{i}(x_{1},x_{2},x_{3},x_{4})^{3}+x_{0}(x_{5}^{2}+x_{6}^{2})$ $(c_{0}:c_{1}:c_{2}:c_{3}:c_{4})\in\mathbb{P}^{4}$ 44
12 C-F $x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{3}^{2}x_{5}+x_{0}x_{4}x_{5}+x_{1}^{2}x_{6}+x_{0}x_{2}x_{6}$ none 0
13 C-F $x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{5}^{2}+x_{1}x_{3}x_{4}+\rho\,x_{2}x_{3}x_{5}+x_{1}^{2}x_{6}+\sigma\,x_{2}^{2}x_{6}+x_{0}x_{3}x_{6}$ $(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}$ 22
14 C-F $x_{2}^{3}+x_{0}\bigl(x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)+x_{1}\bigl(x_{3}^{2}+\tau x_{4}^{2}-x_{5}^{2}+2x_{6}^{2}\bigr)$ $\tau\in\mathbb{C}^{\times}$, $\tau\neq\pm 1,2$ 11
15 C-F $x_{3}^{2}x_{4}+\tau\,x_{3}x_{4}^{2}+x_{0}x_{3}x_{5}+x_{0}x_{4}x_{6}+x_{1}^{2}x_{5}+\rho\,x_{2}^{2}x_{6}$ $(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$ 22
16 C-F $x_{2}^{3}+x_{3}^{3}+\sigma x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+x_{1}x_{2}x_{5}+x_{0}x_{3}x_{6}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+\kappa\,x_{2}^{2}x_{3}+\nu\,x_{2}x_{3}^{2}+\mu\,x_{3}^{2}x_{4}+\lambda\,x_{2}^{2}x_{4}$ $(\sigma,\rho,\kappa,\nu,\mu,\lambda)\in(\mathbb{C}^{\times})^{6}$ 66
17 C-F $x_{3}^{2}x_{4}+\tau\,x_{3}x_{4}^{2}+x_{0}x_{3}x_{5}+x_{1}x_{4}x_{5}+x_{0}^{2}x_{6}+x_{1}^{2}x_{6}+\rho\,x_{2}^{2}x_{6}$ $(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$ 22
18 C-F $x_{2}^{3}+\sigma_{1}x_{3}^{3}+\sigma_{2}x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+x_{0}x_{2}x_{5}+x_{1}x_{3}x_{5}+\alpha\,x_{0}x_{4}x_{5}+x_{0}x_{2}x_{6}+\beta\,x_{1}x_{3}x_{6}+\gamma\,x_{1}x_{4}x_{6}$ $(\sigma_{1},\sigma_{2},\rho,\alpha,\beta,\gamma)\in(\mathbb{C}^{\times})^{6}$ 66
19 C-F $x_{3}^{3}+(x_{0}x_{4}^{2}+x_{1}x_{4}x_{5}+x_{2}x_{5}^{2})+x_{6}(x_{0}^{2}+x_{2}^{2}+\rho\,x_{1}^{2}+\sigma\,x_{0}x_{2})$ $(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}$ 22
20 C-F $x_{4}^{3}+x_{5}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})+x_{6}(x_{0}^{2}+\tau x_{1}^{2}-x_{2}^{2}+2x_{3}^{2})$ $\tau\in\mathbb{C}^{\times}$, $\tau\neq\pm 1,2$ 11
21 C-F $\sum_{i=0}^{4}c_{i}L_{i}(x_{2},x_{3},x_{4},x_{5})^{3}+x_{6}(x_{0}^{2}+x_{1}^{2})$ $(c_{0}:c_{1}:c_{2}:c_{3}:c_{4})\in\mathbb{P}^{4}$ 44
5 Singular loci of 21 polystable cubic fivefolds
In this section, we determine the singular loci of the closed–orbit representatives $\phi^{\mathrm{nf}}_{k}$ constructed in Section 4. For each $k=1,\dots,21$, we set
$X_{k}:=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$
and compute
$\mathrm{Sing}(X_{k})\;=\;V\big(J(\phi^{\mathrm{nf}}_{k})\big)\subset\mathbb{P}^{6},\qquad J(\varphi):=\left(\frac{\partial\varphi}{\partial x_{0}},\dots,\frac{\partial\varphi}{\partial x_{6}}\right):(x_{0},\dots,x_{6})^{\infty},$
i.e. the scheme cut out by the saturated Jacobian ideal. We give a set–theoretic description of $\mathrm{Sing}(X_{k})$ in each case, and when isolated points occur, we also record the corank and the local invariants (Milnor and Tjurina numbers, which agree for our quasi–homogeneous normal forms). Table 3 presents a compact summary—listing the type and degree of the top–dimensional part and indicating the presence (or absence) of isolated points. Detailed case–by–case statements are recorded as Propositions 5.85.28.
Isolated singularities appearing in this paper
We define the isolated singularities that appear in the investigation of the singularities of $\varphi_{k}^{\mathrm{nf}}$. Since we could not locate an explicit normal form in the standard tables (e.g. [AGV12]), so we fix notation here.
Definition 5.1.
We write $\mathrm{QH}(9,6,5;24)_{19}$ for the isolated hypersurface singularity analytically equivalent to $X^{2}Y+Y^{4}+XZ^{3}$. It is quasi-homogeneous of total degree $24$ with respect to weights $(w_{X},w_{Y},w_{Z})=(9,6,5)$. Its corank, Milnor number and Tjurina numbers are 33,1919 and $19$ respectively.
Definition 5.2.
We write $\mathrm{QH}(2,3,5;12)_{21}(t)$ for the isolated hypersurface singularity analytically equivalent to $tX^{6}+X^{3}Y^{2}+X^{2}YZ+XZ^{2}+Y^{4}$, where $t\in\mathbb{C}^{\times}$ is a parameter. It is quasi-homogeneous of total degree $12$ with respect to weights $(w_{X},w_{Y},w_{Z})=(2,3,5)$. Its corank, Milnor number and Tjurina numbers are 33,2121 and $21$ respectively.
Definition 5.3.
We write $\mathrm{QH}(3,3,4;12)_{18}(t)$ for the isolated hypersurface singularity analytically equivalent to $Z^{3}+X^{4}+tX^{2}Y^{2}+Y^{4}$. It is quasi-homogeous of total degree $12$ with respect to weights $(w_{X},w_{Y},w_{Z})=(3,3,4)$. Its corank, Milnor number and Tjurina number are $3$, $18$ and $18$ respectively.
Definition 5.4.
We write $\mathrm{QH}(1,1;6)_{25}$ for the isolated hypersurface singularity analytically equivalent to a homogeneous sextic
$\Psi(X,Y)=\prod_{i=1}^{6}L_{i}(X,Y),$
where $L_{1},\dots,L_{6}$ are pairwise distinct linear forms in $(X,Y)$. Equivalently, after a linear change of coordinates, one may write $\Psi(X,Y)=\prod_{i=1}^{6}(X-\lambda_{i}Y)$ with $\lambda_{i}$ pairwise distinct in $\mathbb{P}^{1}$.
It is quasi-homogeneous of total degree $6$ with respect to weights $(w_{X},w_{Y})=(1,1)$. Its corank, Milnor number and Tjurina number are $2$, $25$ and $25$ respectively.
Definition 5.5.
We write $QH(1,1,1,1;3)_{16}([c])$ for the isolated hypersurface singularity analytically equivalent to a homogeneous cubic
$G(X_{1},X_{2},X_{3},X_{4})=\Sigma_{i=0}^{4}c_{i}L_{i}^{3}$
where $L_{i}=L_{i}(X_{1},X_{2},X_{3},X_{4})$ are generic linear forms and $c]=(c_{0}:c_{1}:c_{2}:c_{3}:c_{4})\in\mathbb{P}^{4}$. Its corank, Milnor number and Tjurina number are $4$, $16$ and $16$. It has four-dimensional moduli.
Definition 5.6.
$\mathrm{QH}(1,2,2;6)_{20}(t)$
$F(X,Y,Z)=Y^{3}+Z^{3}+X^{6}+t_{1}X^{2}Y^{2}+t_{2}X^{2}YZ+t_{3}Y^{2}Z+t_{4}X^{4}Y+t_{5}X^{4}Z,$
where $t=(t_{1},t_{2},\cdots,t_{5})$. Its corank, Milnor number and Tjurina number are $3$, 2020 and $20$.
Remark 5.7 (Remark for Definition 5.4).
The germ $\Psi(X,Y)=\prod_{i=1}^{6}L_{i}(X,Y)$ with pairwise distinct linear forms defines an ordinary sextuple point (also called an ordinary $6$-fold point): it has $6$ smooth branches meeting transversely at the origin, with $6$ distinct tangent directions in $\mathbb{P}^{1}$.
We will frequently use the following shorthand labels for isolated singularity types:
[…]
5.1 Case $k=1$
Proposition 5.8.
Let $X_{1}=V(\phi^{\mathrm{nf}}_{1})\subset\mathbb{P}^{6}$. The set-theoretic singular locus is
$\mathrm{Sing}(X_{1})=C\cup\{P\},$
where
$C=\{\,x_{0}=x_{1}=x_{2}=x_{3}=0,\;x_{5}^{\,2}+x_{4}x_{6}=0\,\}$
is a smooth conic in the plane $\{x_{0}=x_{1}=x_{2}=x_{3}=0\}\cong\mathbb{P}^{2}$ (hence, $\deg C=2$), and
$P=(1:0:0:0:0:0:0)$
is an isolated singular point.
The isolated singular point is of type $\mathsf{E}_{19}$.
5.2 Case $k=2$
Proposition 5.9.
Let $X_{2}=V(\varphi^{\mathrm{nf}}_{2})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{2})=L_{01}\cup C,$
with $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})\simeq\mathbb{P}^{1}$, and
$C=V(x_{0},x_{1},x_{2},\;x_{5}^{2}+x_{3}x_{6},\;x_{4}^{2}+x_{3}x_{5})\subset\Pi,\quad\Pi=\{x_{0}=x_{1}=x_{2}=0\}\simeq\mathbb{P}^{3}.$
Here $C$ is a complete intersection $\mathrm{CI}(2,2)$ of degree $4$, with $L_{01}\cap C=\varnothing$.
5.3 Case $k=3$
Proposition 5.10.
Let $X_{3}=V(\varphi^{\mathrm{nf}}_{3})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{3})=L_{01}\cup C,$
with $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})\simeq\mathbb{P}^{1}$, and
$C=V(x_{0},x_{1},x_{2},x_{5}^{2}+x_{3}x_{6}+x_{4}x_{6},x_{3}^{2}+x_{3}x_{4}+x_{4}^{2})\subset\Pi,\quad\Pi=\{x_{0}=x_{1}=x_{2}=0\}\simeq\mathbb{P}^{3}.$
Here $C$ is a complete intersection $\mathrm{CI}(2,2)$ of degree $4$, with $L_{01}\cap C=\varnothing$.
5.4 Case $k=4$
Proposition 5.11.
Let $X_{4}=V(\varphi^{\mathrm{nf}}_{4})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{4})=L_{01}\cup L_{56},$
with $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})\simeq\mathbb{P}^{1}$ and $L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4})\simeq\mathbb{P}^{1}$. In particular, $L_{01}\cap L_{56}=\varnothing$ and no isolated singular points occur.
5.5 Case $k=5$
Proposition 5.12.
Let $X_{5}=V(\varphi^{\mathrm{nf}}_{5})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{5})=P\cup Q$
$P=(1:0:0:0:0:0:0)$, $Q=(0:0:0:0:0:0:1)$ are isolated singular points.
Then $(X_{5},P)$ and $(X_{5},Q)$ are isolated hypersurface singularities of corank $3$, and they are analytically equivalent. The isolated singular points $P,Q$ are of type $\mathsf{E}_{21}$ and have same parameter $t$.
5.6 Case $k=6$
Proposition 5.13.
Let $X_{6}=V(\varphi^{\mathrm{nf}}_{6})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{6})=\{P\}\cup V(x_{3},x_{4},x_{5},x_{6},\;x_{1}^{2}+x_{0}x_{2}),\quad P=(0:0:0:0:0:0:1).$
The isolated singular point $P$ is of type $\mathsf{E}_{19}$.
5.7 Case $k=7$
Proposition 5.14.
Let $X_{7}=V(\varphi^{\mathrm{nf}}_{7})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{7})=C\cup L_{56},\quad L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4}),$
$C=V(x_{4},x_{5},x_{6},\;x_{2}^{2}+x_{1}x_{3},\;x_{1}^{2}+x_{0}x_{3})\subset\Pi=\{x_{4}=x_{5}=x_{6}=0\}.$
$C$ is a $\mathrm{CI}(2,2)$ of degree $4$; $C\cap L_{56}=\varnothing$.
5.8 Case $k=8$
Proposition 5.15.
Let $X_{8}:=V(\phi^{\mathrm{nf}}_{8})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{8})=C\cup\{P\}$
, where
$C:=V\!\bigl(x_{0},x_{1},x_{2},x_{3},\;x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)\subset\Pi:=\{x_{0}=x_{1}=x_{2}=x_{3}=0\}\cong\mathbb{P}^{2}$
is a (smooth) conic of degree $2$ (hence $\mathrm{codim}_{\mathbb{P}^{6}}C=5$), and
$P:=V\!\bigl(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\bigr)=(1:0:0:0:0:0:0)$
is an isolated point (hence $\mathrm{codim}_{\mathbb{P}^{6}}P=6$, $\deg P=1$). In particular, $C\cap\{P\}=\varnothing$. The isolated singular point $P$ is of type $\mathsf{E}_{18}$.
5.9 Case $k=9$
Proposition 5.16.
Let $X_{9}:=V(\phi^{\mathrm{nf}}_{9})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{9})=L_{01}\cup P$
Where $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})$ and $P=(0:0:0:0:0:0:1)$. The isolated point $P$ is of type $\mathcal{O}_{6}$.
5.10 Case $k=10$
Proposition 5.17.
Let $X_{10}:=V(\phi^{\mathrm{nf}}_{10})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{10})=L_{01}\cup C,$
where $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})$ and $C=V(x_{0},x_{1},x_{2},x_{3},x_{4}x_{5}+x_{6}^{2})$ which is a curve of degree $2$. We have $L_{01}\cap C=\varnothing$.
5.11 Case $k=11$
Proposition 5.18.
Let $X_{11}:=V(\phi^{\mathrm{nf}}_{11})\subset\mathbb{P}^{6}$. Then the singular locus has three irreducible components:
$\mathrm{Sing}(X_{11})=P_{1}\cup P_{2}\cup P_{3},$
where $P_{1},P_{2}$ and $P_{3}$ are isolated singular points of type $\mathcal{C}_{cub}$.
5.12 Case $k=12$
Proposition 5.19.
Let $X_{12}:=V(\phi^{\mathrm{nf}}_{12})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{12})=L_{56}\cup C,$
where $L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4})$ is a projective line and $C=V(x_{2}^{2}+x_{3}^{2},x_{1}^{2}+x_{0}x2,x_{4},x_{5},x_{6})$ is a $(2,2)$-complete intersection. We have $L_{56}\cap C=\varnothing$.
5.13 Case $k=13$
Proposition 5.20.
Let $X_{13}:=V(\phi^{\mathrm{nf}}_{13})\subset\mathbb{P}^{6}$. Then the singular locus has two connected components:
$\mathrm{Sing}(X_{13})=C_{1}\cup C_{2},$
where $C_{1}=V(x_{1}^{2}+3x_{2}^{2},x_{3},x_{4},x_{5},x_{6})$ and $C_{2}=V(x_{0},x_{1},x_{2},x_{3},x_{4}^{2}+x_{5}^{2})$. Each $C_{i}$ consists of two lines which intersect at one point. We have $C_{1}\cap C_{2}=\varnothing$.
5.14 Case $k=14$
Proposition 5.21.
Let $X_{14}:=V(\phi^{\mathrm{nf}}_{14})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{14})=L_{01}\cup C,$
where $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})$ and $C$ is a $(2,2)$-complete intersection in $V(x_{0},x_{1},x_{2})\simeq\mathbb{P}^{3}$. We have $L_{01}\cap C=\varnothing$.
5.15 Case $k=15$
Proposition 5.22.
Let $X_{15}:=V(\phi^{\mathrm{nf}}_{15})\subset\mathbb{P}^{6}$. Then the singular locus consists of three irreducible components:
$\mathrm{Sing}(X_{15})=C_{1}\cup C_{2}\cup L,$
where $C_{1}=V(x_{1},x_{3},x_{6},2x_{4}^{2}+x_{0}x_{5},3x_{2}^{2}+x_{0}x_{4})$, $C_{2}=V(x_{2},x_{4},x_{5},x_{3}^{2}+x_{0}x_{6},x_{1}^{2}+x_{0}x_{3})$, and $L=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C_{1}$ and $C_{2}$ are curves of degree $4$ (complete intersections of two quadrics in $\mathbb{P}^{3}$), and $L$ is a line.
5.16 Case $k=16$
Proposition 5.23.
Let $X_{16}:=V(\phi^{\mathrm{nf}}_{16})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{16})=P\cup Q,$
$P=(1:0:0:0:0:0:0)$, $Q=(0:0:0:0:0:0:1)$, where $P,Q$ are isolated singular points of type $\mathsf{E}_{20}$ with the same modulus.
5.17 Case $k=17$
Proposition 5.24.
Let $X_{17}:=V(\phi^{\mathrm{nf}}_{17})\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
$\mathrm{Sing}(X_{17})=C\cup L_{56},$
where $C=V(x_{3},x_{4},x_{5},x_{6},x_{0}^{2}+x_{1}^{2}+3x_{2}^{2})$ and $L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C$ is a smooth conic (degree $2$) and $L$ is a line (degree $1$). We have $C\cap L=\varnothing$.
5.18 Case $k=18$
Proposition 5.25.
Let $X_{18}:=V(\phi^{\mathrm{nf}}_{18})\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
$\mathrm{Sing}(X_{18})=L_{1}\cup L_{2},$
where $L_{1}=V(x_{2},x_{3},x_{4},x_{5},x_{6})$ and $L_{2}=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, both $L_{1}$ and $L_{2}$ are lines (degree $1$). We have $L_{1}\cap L_{2}=\varnothing$.
5.19 Case $k=19$
Proposition 5.26.
Let $X_{19}:=V(\phi^{\mathrm{nf}}_{19})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
$\mathrm{Sing}(X_{19})=P\cup C,$
$P=(0:0:0:0:0:0:1)$, $C=V(x_{0}^{2}+\rho x_{1}^{2}+\sigma x_{0}x_{2}+x_{2}^{2},x_{3},x_{4},x_{5},x_{6})$. Here $P$ is an isolated singular point of type $\mathsf{E}_{18}$.
5.20 Case $k=20$
Proposition 5.27.
Let $X_{20}:=V(\phi^{\mathrm{nf}}_{20})\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
$\mathrm{Sing}(X_{20})=C\cup L,$
where $C=V(x_{4},x_{5},x_{6},4x_{1}^{2}-2x_{2}^{2}+x_{3}^{2},4x_{0}^{2}+6x_{2}^{2}+3x_{3}^{2})$ and $L=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C$ is a curve of degree $4$ (a complete intersection of two quadrics, which is an elliptic curve) and $L$ is a line (degree $1$). We have $C\cap L=\varnothing$.
5.21 Case $k=21$
Proposition 5.28.
Let $X_{21}:=V(\phi^{\mathrm{nf}}_{21})\subset\mathbb{P}^{6}$. Then the singular locus has three irreducible components:
$\mathrm{Sing}(X_{21})=P_{1}\cup P_{2}\cup P_{3},$
where $P_{1},P_{2}$ and $P_{3}$ are isolated singular points of type $\mathcal{C}_{cub}$.
Table 3: Singular loci of the 21 closed-orbit representatives (Section 5)
$k$ Singular locus (notation of §5) Type / degree of top-dimensional part Isolated point(s) / invariants
1 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{19}$
2 $L_{01}\cup C$ ($L_{01}\cong\mathbb{P}^{1}$, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
3 $L_{01}\cup C$ ($L_{01}\cong\mathbb{P}^{1}$, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
4 $L_{01}\cup L_{56}$ (two disjoint lines) two disjoint lines (deg 1+1)
5 $\{P\}\cup\{Q\}$ two points $P,Q$ of type $\mathsf{E}_{21}$ (same modulus)
6 $\{P\}\cup C$ ($C$ conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{19}$
7 $C\cup L_{56}$ ($C$ a CI(2,2)(2,2), $L_{56}$ line) line ++ quartic curve (deg 4)
8 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{18}$
9 $L_{01}\cup\{P\}$ ($L_{01}$ line) line (deg 1) one point $P$ of type $\mathcal{O}_{6}$
10 $L_{01}\cup C$ ($L_{01}$ line, $C$ conic) line ++ conic (deg 2)
11 $\{P_{1}\}\cup\{P_{2}\}\cup\{P_{3}\}$ three points three points $P_{i}$ of type $\mathcal{C}_{cub}$
12 $L_{56}\cup C$ ($L_{56}$ line, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
13 $C_{1}\cup C_{2}$ (disjoint) two pairs of intersecting lines
14 $L_{01}\cup C$ ($L_{01}$ line, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
15 $C_{1}\cup C_{2}\cup L$ (two CI(2,2)(2,2)s + line) two quartics ++ line
16 $\{P\}\cup\{Q\}$ two points $P,Q$ of type $\mathsf{E}_{20}$ (same modulus)
17 $C\cup L$ ($C$ smooth conic, $L$ line) conic (deg 2) ++ line
18 $L_{1}\cup L_{2}$ (two disjoint lines) two disjoint lines
19 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{18}$
20 $C\cup L$ ($C$ elliptic curve, $L$ line) elliptic curve (deg 4) ++ line
21 $\{P_{1}\}\cup\{P_{2}\}\cup\{P_{3}\}$ three points three points $P_{i}$ of type $\mathcal{C}_{cub}$
Table 4: List of isolated singularities
Symbol Type Defining Equation (Normal Form)
$\mathsf{E}_{19}$ $\mathrm{QH}(9,6,5;24)_{19}$ $X^{2}Y+Y^{4}+XZ^{3}=0$
$\mathsf{E}_{21}$ $\mathrm{QH}(2,3,5;12)_{21}(t)$ $tX^{6}+X^{3}Y^{2}+X^{2}YZ+XZ^{2}+Y^{4}=0$
$\mathsf{E}_{18}$ $\mathrm{QH}(3,3,4;12)_{18}(t)$ $Z^{3}+X^{4}+tX^{2}Y^{2}+Y^{4}=0$
$\mathcal{O}_{6}$ $\mathrm{QH}(1,1;6)_{25}$ $\prod_{i=1}^{6}(X-\lambda_{i}Y)=0$
$\mathcal{C}_{cub}$ $\mathrm{QH}(1,1,1,1;3)_{16}([c])$ $\sum_{i=0}^{4}c_{i}L_{i}^{3}=0$
$\mathsf{E}_{20}$ $\mathrm{QH}(1,2,2;6)_{20}(t)$ $Y^{3}+Z^{3}+X^{6}+t_{1}X^{2}Y^{2}+\cdots+t_{5}X^{4}Z=0$
5.22 Minimal exponents and connections to Park’s stability criterion
In Definitions 5.1–5.4 and 5.6–5.7 we introduced six quasi-homogeneous isolated hypersurface singularities that occur as the isolated points on the GIT boundary. The next theorem records a uniform numerical feature of these singularities, and explains why they sit exactly at the critical value in Park’s stability criterion.
Theorem 5.29.
Let $(X,x)$ be one of the six isolated boundary singularities from Definitions 5.1–5.6. Then the local minimal exponent is
$\widetilde{\alpha}_{x}(X)=\frac{7}{3}.$
Proof.
By the holomorphic splitting lemma, the local defining equation of $X$ at an isolated singular point $x$ is analytically equivalent to
$u_{1}^{2}+\cdots+u_{6-c}^{2}\;+\;g(z_{1},\dots,z_{c}),$
where $c=\mathrm{corank}_{x}(X)$ and $g$ is one of the weighted-homogeneous germs from Definitions 5.1–5.4 and 5.6–5.7.
By Thom–Sebastiani for minimal exponents [Par25, Theorem 2.3],
$\widetilde{\alpha}_{x}(X)=\widetilde{\alpha}_{0}\!\left(u_{1}^{2}+\cdots+u_{6-c}^{2}\right)+\widetilde{\alpha}_{0}(g)=\frac{6-c}{2}+\widetilde{\alpha}_{0}(g).$
For a weighted-homogeneous isolated hypersurface singularity of type $\mathrm{QH}(w_{1},\dots,w_{c};D)$, we have $\widetilde{\alpha}_{0}(g)=\frac{w_{1}+\cdots+w_{c}}{D}$. In our list, this yields
$\widetilde{\alpha}_{0}(g)=\begin{cases}\frac{5}{6},&g\in\{\mathrm{QH}(9,6,5;24)_{19},\,\mathrm{QH}(2,3,5;12)_{21}(t),\,\mathrm{QH}(3,3,4;12)_{18}(t),\,\mathrm{QH}(1,2,2;6)_{20}(t)\},\\[2.0pt] \frac{1}{3},&g=\mathrm{QH}(1,1;6)_{25},\\[2.0pt] \frac{4}{3},&g=\mathrm{QH}(1,1,1,1;3)_{16}([c]).\end{cases}$
The corresponding coranks are $c=3,2,4$ respectively (cf. the definitions), hence in each case
$\widetilde{\alpha}_{x}(X)=\frac{6-c}{2}+\widetilde{\alpha}_{0}(g)=\frac{7}{3}.$
Significance (relation to Park). Since $\frac{7}{3}>2$, these isolated singularities are $1$-rational (hence in particular rational and Du Bois) by [Par25, Theorem 2.2]. More importantly, for cubic fivefolds $(n,d)=(6,3)$ we have $\frac{n+1}{d}=\frac{7}{3}$, which is exactly the threshold in Park’s GIT criterion: a degree $d$ hypersurface $X\subset\mathbb{P}^{n}$ is GIT stable (resp. semistable) if $\widetilde{\alpha}(X)>\frac{n+1}{d}$ (resp. $\geq$) [Par25, Theorem A]. Thus, any cubic fivefold containing one of these singular points satisfies $\widetilde{\alpha}(X)\leq\frac{7}{3}$, so it cannot lie in the region $\widetilde{\alpha}(X)>\frac{7}{3}$ guaranteed to be GIT stable by Park. This gives a conceptual explanation for why these singularities occur precisely on the GIT boundary.
Definition 5.30.
We call the above six singularity types extremal cubic fivefold singularities, since they realize the critical value
$\frac{n+1}{d}=\frac{7}{3}\qquad(n=6,\ d=3)$
in Park’s GIT criterion. In particular, each of these types has local minimal exponent $\frac{7}{3}$. Motivated by this, we refer to isolated cubic fivefold singularities with local minimal exponent $\frac{7}{3}$ as $\frac{7}{3}$-singularities; the six types above occur on the GIT boundary.
Remark 5.31 (Unification via minimal exponent).
Park observes that, for cubic hypersurfaces $X\subset\mathbb{P}^{n}$ with $n\leq 5$, the classical explicit GIT analyses can be reformulated in terms of the same numerical threshold $\frac{n+1}{3}$ for the (global) minimal exponent; see [Par25, Remark 7.6]. Theorem 5.29 shows that for cubic fivefolds $(n=6)$ the isolated boundary singularities also satisfy equality $\widetilde{\alpha}_{x}(X)=\frac{7}{3}=\frac{n+1}{3}$, fitting the same pattern.
6 Adjacency relations among strictly semistable components
In this section, we record the adjacency relations among the closed strata $\{\Phi_{k}\}_{k=1}^{21}\subset\mathbb{P}(W)^{\mathrm{ss}}$, where $W=\mathrm{Sym}^{3}\mathbb{C}^{7}$. We adopt the following convention: two closed strata $\Phi_{i}$ and $\Phi_{j}$ are adjacent if there exists a codimension-one wall in the Hilbert–Mumford weight space such that the maximally destabilizing one-parameter subgroups (1-PS’s) for general points of $\Phi_{i}$ and $\Phi_{j}$ coincide on that wall, and the corresponding degenerations specialize to the same unstable limit in $\mathbb{P}(W)$ (see Definition 6.2).
This notion of adjacency captures wall-crossing in Kirwan’s stratification and, from the combinatorial viewpoint, corresponds to codimension-one faces of the cones $I(r)_{\geq 0}$ determined by 1-PS weight vectors $r$.
Our strategy is as follows.
  1. (1)
    For each pair of reduced weight vectors $r_{i},r_{j}$ listed in Section 2, we study the slices
    $I(r)_{\geq 0}\cap I(r_{i})_{=0},\qquad I(r)_{\geq 0}\cap I(r_{j})_{=0},$
    and determine whether they are inclusion-maximal subsets inside the corresponding hyperplanes. The same algorithm as in Section 2 applies verbatim.
  2. (2)
    If these slices are maximal precisely for $r=r_{j}$ (resp. $r=r_{i}$), then $\Phi_{i}$ and $\Phi_{j}$ admit a common wall degeneration.
  3. (3)
    Using the explicit normal forms from Section 4, we verify adjacency by checking that suitable representatives $\widetilde{\varphi}_{i}\in\Phi_{i}$ and $\widetilde{\varphi}_{j}\in\Phi_{j}$ satisfy
    $\lim_{t\to 0}\lambda_{j}(t)\cdot\widetilde{\varphi}_{i}\;=\;\lim_{t\to 0}\lambda_{i}(t)\cdot\widetilde{\varphi}_{j},$
    where $\lambda_{k}$ denotes the 1-PS corresponding to $r_{k}$.
  4. (4)
    For all remaining pairs, we rule out adjacency by support considerations.
The main result of this section is the following classification.
Theorem 6.1.
Among the closed strata $\Phi_{k}$ in the strictly semistable locus, the only adjacent pairs are the following eight pairs:
$\{\Phi_{1},\Phi_{7}\},\ \{\Phi_{2},\Phi_{6}\},\ \{\Phi_{3},\Phi_{12}\},\ \{\Phi_{8},\Phi_{19}\},\ \{\Phi_{9},\Phi_{15}\},\ \{\Phi_{10},\Phi_{17}\},\ \{\Phi_{11},\Phi_{21}\},\ \{\Phi_{14},\Phi_{20}\}.$
All other pairs of closed strata are not adjacent. Equivalently, the adjacency graph of the strictly semistable boundary consists precisely of these eight edges (see Propositions 6.3–6.10).
For the following definition, see [Kir84, DH98, Tha96].
Definition 6.2 (Adjacency via wall-crossing).
Let $\Phi_{i},\Phi_{j}$ be closed strata in the strictly semistable locus. We say that $\Phi_{i}$ and $\Phi_{j}$ are adjacent if there exists a codimension-one wall in the Hilbert–Mumford weight space such that for general $f\in\Phi_{i}$ and $g\in\Phi_{j}$ one has $\mu(f,\lambda_{i})=\mu(f,\lambda_{j})$ on the wall, and both 1-PS degenerations specialize to the same unstable limit in $\mathbb{P}(W)$.
The following propositions verify Theorem 6.1 by the same combinatorial procedure as in Section 2. More precisely, for each $k$ we work inside the hyperplane $I(r_{k})_{=0}\subset I$ and determine inclusion-maximal subsets of the form $I(r)_{\geq 0}\cap I(r_{k})_{=0}$, where $r\in\mathbb{Z}^{7}_{(0)}$ ranges over 1-PS’s of the maximal torus $T$.
Proposition 6.3.
The closed strata $\Phi_{1}$ and $\Phi_{7}$ are both zero-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{1})_{=0}$ is inclusion-maximal in $I(r_{1})_{=0}$ if and only if $r=r_{7}$, and $I(r)_{\geq 0}\cap I(r_{7})_{=0}$ is inclusion-maximal in $I(r_{7})_{=0}$ if and only if $r=r_{1}$. Moreover, for suitable specializations $\widetilde{\varphi}_{1},\widetilde{\varphi}_{7}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{7}(t)\cdot\widetilde{\varphi}_{1}$ $=\lim_{t\to 0}\lambda_{1}(t)\cdot\widetilde{\varphi}_{7}=x_{0}x_{5}^{2}+x_{1}^{2}x_{6}.$
Proposition 6.4.
The closed strata $\Phi_{2}$ and $\Phi_{6}$ are both zero-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{2})_{=0}$ is inclusion-maximal in $I(r_{2})_{=0}$ if and only if $r=r_{6}$, and $I(r)_{\geq 0}\cap I(r_{6})_{=0}$ is inclusion-maximal in $I(r_{6})_{=0}$ if and only if $r=r_{2}$. Moreover, for suitable specializations $\widetilde{\varphi}_{2},\widetilde{\varphi}_{6}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{6}(t)\cdot\widetilde{\varphi}_{2}$ $=\lim_{t\to 0}\lambda_{2}(t)\cdot\widetilde{\varphi}_{6}=x_{1}x_{3}x_{5}+x_{0}x_{3}x_{6}.$
Proposition 6.5.
The closed strata $\Phi_{3}$ and $\Phi_{12}$ are both zero-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{3})_{=0}$ is inclusion-maximal in $I(r_{3})_{=0}$ if and only if $r=r_{12}$, and $I(r)_{\geq 0}\cap I(r_{12})_{=0}$ is inclusion-maximal in $I(r_{12})_{=0}$ if and only if $r=r_{3}$. Moreover, for suitable specializations $\widetilde{\varphi}_{3},\widetilde{\varphi}_{12}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{12}(t)\cdot\widetilde{\varphi}_{3}$ $=\lim_{t\to 0}\lambda_{3}(t)\cdot\widetilde{\varphi}_{12}=x_{1}x_{4}^{2}+x_{2}^{2}x_{5}+x_{0}x_{3}x_{6}.$
Proposition 6.6.
The closed strata $\Phi_{8}$ and $\Phi_{19}$ are both two-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{8})_{=0}$ is inclusion-maximal in $I(r_{8})_{=0}$ if and only if $r=r_{19}$, and $I(r)_{\geq 0}\cap I(r_{19})_{=0}$ is inclusion-maximal in $I(r_{19})_{=0}$ if and only if $r=r_{8}$. Moreover, for suitable specializations $\widetilde{\varphi}_{8},\widetilde{\varphi}_{19}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{19}(t)\cdot\widetilde{\varphi}_{8}$ $=\lim_{t\to 0}\lambda_{8}(t)\cdot\widetilde{\varphi}_{19}$
$=x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{4}x_{5}+x_{0}x_{5}^{2}+x_{1}^{2}x_{6}+x_{1}x_{2}x_{6}+x_{2}^{2}x_{6}.$
Proposition 6.7.
The closed strata $\Phi_{9}$ and $\Phi_{15}$ are both two-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{9})_{=0}$ is inclusion-maximal in $I(r_{9})_{=0}$ if and only if $r=r_{15}$, and $I(r)_{\geq 0}\cap I(r_{15})_{=0}$ is inclusion-maximal in $I(r_{15})_{=0}$ if and only if $r=r_{9}$. Moreover, for suitable specializations $\widetilde{\varphi}_{9},\widetilde{\varphi}_{15}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{15}(t)\cdot\widetilde{\varphi}_{9}$ $=\lim_{t\to 0}\lambda_{9}(t)\cdot\widetilde{\varphi}_{15}=x_{3}^{3}+x_{0}x_{4}x_{5}+x_{1}x_{2}x_{6}+x_{0}x_{3}x_{6}.$
Proposition 6.8.
The closed strata $\Phi_{10}$ and $\Phi_{17}$ are both two-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{10})_{=0}$ is inclusion-maximal in $I(r_{10})_{=0}$ if and only if $r=r_{17}$, and $I(r)_{\geq 0}\cap I(r_{17})_{=0}$ is inclusion-maximal in $I(r_{17})_{=0}$ if and only if $r=r_{10}$. Moreover, for suitable specializations $\widetilde{\varphi}_{10},\widetilde{\varphi}_{17}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{17}(t)\cdot\widetilde{\varphi}_{10}$ $=\lim_{t\to 0}\lambda_{10}(t)\cdot\widetilde{\varphi}_{17}=x_{3}^{3}+x_{1}x_{3}x_{5}+x_{0}x_{4}x_{5}+x_{1}x_{2}x_{6}.$
Proposition 6.9.
The closed strata $\Phi_{11}$ and $\Phi_{21}$ are both four-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{11})_{=0}$ is inclusion-maximal in $I(r_{11})_{=0}$ if and only if $r=r_{21}$, and $I(r)_{\geq 0}\cap I(r_{21})_{=0}$ is inclusion-maximal in $I(r_{21})_{=0}$ if and only if $r=r_{11}$. Moreover, for suitable specializations $\widetilde{\varphi}_{11},\widetilde{\varphi}_{21}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{21}(t)\cdot\widetilde{\varphi}_{11}$ $=\lim_{t\to 0}\lambda_{11}(t)\cdot\widetilde{\varphi}_{21}$
$=x_{2}^{3}+x_{2}^{2}x_{3}+x_{2}x_{3}^{2}+x_{3}^{3}+x_{2}^{2}x_{4}+x_{2}x_{3}x_{4}+x_{3}^{2}x_{4}+x_{2}x_{4}^{2}+x_{3}x_{4}^{2}+x_{4}^{3}.$
Proposition 6.10.
The closed strata $\Phi_{14}$ and $\Phi_{20}$ are both one-dimensional and adjacent. For $r\in\mathbb{Z}^{7}_{(0)}$, the slice $I(r)_{\geq 0}\cap I(r_{14})_{=0}$ is inclusion-maximal in $I(r_{14})_{=0}$ if and only if $r=r_{20}$, and $I(r)_{\geq 0}\cap I(r_{20})_{=0}$ is inclusion-maximal in $I(r_{20})_{=0}$ if and only if $r=r_{14}$. Moreover, for suitable specializations $\widetilde{\varphi}_{14},\widetilde{\varphi}_{20}$ (with all nonzero coefficients normalized to $1$), one has
$\lim_{t\to 0}\lambda_{20}(t)\cdot\widetilde{\varphi}_{14}$ $=\lim_{t\to 0}\lambda_{14}(t)\cdot\widetilde{\varphi}_{20}$
$=x_{0}x_{3}x_{5}+x_{1}x_{3}x_{5}+x_{0}x_{3}x_{6}+x_{1}x_{3}x_{6}.$
Collecting Propositions 6.36.10, we obtain Theorem 6.1.